### v1

# A Free Probabilistic Framework for Denoising Diffusion Models: Entropy, Transport, and Reverse Processes

Abstract
This work develops a rigorous framework for diffusion-based generative modeling in the setting of free probability. We extend classical denoising diffusion probabilistic models to free diffusion processes—stochastic dynamics acting on noncommutative random variables whose spectral measures evolve by free additive convolution. The forward dynamics satisfy a free Fokker–Planck equation that increases Voiculescu’s free entropy and dissipates free Fisher information, providing a noncommutative analogue of the classical de Bruijn identity. Using tools from free stochastic analysis, including a free Malliavin calculus and a Clark–Ocone representation, we derive the reverse-time stochastic differential equation driven by the conjugate variable, the free analogue of the score function. We further develop a variational formulation of these flows in the free Wasserstein space, showing that the resulting gradient-flow structure converges to the semicircular equilibrium law. Together, these results connect modern diffusion models with the information geometry of free entropy and establish a mathematical foundation for generative modeling with operator-valued or high-dimensional structured data.
I Introduction
Denoising diffusion probabilistic models (DDPMs) and their continuous-time analogues have recently emerged as a central paradigm in modern generative modeling. In these models, a forward stochastic process gradually transforms a complex data distribution into a simple Gaussian prior, while a reverse-time stochastic differential equation (SDE) reconstructs data samples by inverting this transformation. This framework, initiated in [1, 2] and placed on a rigorous continuous-time footing in [3, 4], is deeply connected to the Fokker–Planck equation, entropy production, and the $L_{2}$–Wasserstein geometry of optimal transport. Subsequent research has revealed elegant links to Schrödinger bridges [4], stochastic interpolants [5], and information-theoretic interpretations of score matching [6], situating diffusion models at the intersection of probability, statistical physics, and information geometry. For a detailed account of recent theoretical developments in diffusion-based generative modeling, see [7].
However, the existing theory remains fundamentally commutative. The noise is Gaussian, the random variables inhabit Euclidean spaces, and correlations are captured by classical covariance. In contrast, contemporary machine learning increasingly operates in operator-valued and inherently non-commutative regimes. Large neural networks, for example, are more naturally characterized by their random matrix spectra rather than by vector-valued features, as demonstrated in empirical spectral analyses of deep architectures [8, 9, 10]. In such settings, classical probability fails to capture the operator-level correlations and non-commuting dependencies that emerge from matrix multiplications and compositional network structure. This gap motivates a search for a non-commutative foundation for generative modeling.
Free probability theory provides precisely such a foundation. Introduced by Voiculescu [11, 12] and developed further by Biane [13] and Biane–Speicher [14], free probability extends stochastic calculus, entropy, and information theory to non-commutative random variables. Its connections to random matrix theory are well established: empirical spectral distributions converge to free convolutions, and asymptotic freeness governs large-dimensional limits [15]. Moreover, the theory introduces free analogues of fundamental information measures—free entropy and free Fisher information—which mirror their Shannon counterparts [12]. The development of free optimal transport by Biane and Voiculescu [16] and its extensions in [17] established a non-commutative analogue of the $L_{2}$–Wasserstein space. More recently, free probabilistic ideas have begun to influence deep learning theory itself: for example, Das [18] used operator-valued free convolutions to analyze Transformer attention as a form of non-commutative stochastic interaction.
These developments naturally raise the following question: Can diffusion-based generative modeling be extended into the free, non-commutative domain, while retaining the geometric and information-theoretic structure of classical diffusions?
This paper provides an affirmative answer. We develop a rigorous framework for free diffusion models, extending denoising diffusion processes into the realm of free probability. By formulating both the forward and reverse dynamics in terms of free stochastic calculus, we establish explicit connections between free entropy, free Fisher information, and non-commutative Wasserstein geometry. Our analysis further leads to a family of functional inequalities—including free analogues of the logarithmic Sobolev, Talagrand, and HWI inequalities (linking H for relative entropy, W for Wasserstein distance, and I for Fisher information)—that govern convergence, stability, and transport in the free setting.
Our main contributions are summarized as follows:
  • Forward diffusion as a free Ornstein–Uhlenbeck process: The forward dynamics are modeled by a free Ornstein–Uhlenbeck diffusion [19], whose spectral law evolves by free additive convolution, with entropy growth governed by Voiculescu’s free entropy.
  • Reverse process via free Malliavin calculus: Using the free Clark–Ocone formula, we derive a reverse-time SDE driven by the conjugate variable, the free analogue of the classical score function. A complementary PDE derivation confirms its role as a free deconvolution flow.
  • Information-theoretic and variational structure: We prove a free de Bruijn identity linking entropy production to free Fisher information, show that the forward process forms a gradient ascent of free entropy $\chi$, and construct a JKO-type variational scheme in the free Wasserstein metric.
  • Functional inequalities and convergence guarantees: Building upon the variational framework, we establish free analogues of the logarithmic Sobolev, Talagrand, and HWI inequalities [20, 17], providing quantitative bounds on entropy decay, Fisher information dissipation, and Wasserstein contraction toward the semicircular equilibrium.
  • Free Schrödinger bridge and energy–entropy balance: Through a free Girsanov theorem, we characterize the reverse diffusion as a free Schrödinger bridge, revealing a precise energy–entropy trade-off in the non-commutative setting.
  • Universality and spectral convergence: Under mild noise-schedule conditions, the forward process converges to the semicircular law, while the reverse dynamics can approximate any compactly supported spectral distribution.
Taken together, these results position free diffusion models as Wasserstein gradient flows of Voiculescu’s free entropy, unifying stochastic diffusion, functional inequalities, and non-commutative information geometry under a single rigorous framework. Beyond serving as a conceptual extension of classical diffusion models, this theory introduces new analytical tools for operator-valued and high-dimensional structured data, opening a research program at the intersection of free probability, optimal transport, and information-theoretic generative modeling.
II Classical Diffusion Models and DDPMs
Diffusion probabilistic models, or denoising diffusion probabilistic models (DDPMs), have become one of the most powerful frameworks in modern generative modeling. Their central idea is conceptually simple yet mathematically deep: a forward stochastic process gradually destroys the structure of complex data by adding noise, while a reverse stochastic process learns to reconstruct the data distribution from pure noise. Together, these two processes define a generative model capable of sampling from highly structured, high-dimensional distributions such as natural images, audio, or molecular configurations.
Formally, the model is built around a pair of stochastic processes $\{x_{t}\}_{t\geq 0}$ (forward) and $\{x_{t}\}_{t\leq T}$ (reverse) evolving in $\mathbb{R}^{d}$ [1, 3]. The forward process serves as a data–to–noise map, while the reverse process acts as a noise–to–data decoder governed by a learned stochastic differential equation.
II-A Forward Process: From Data to Noise
In the forward direction, the model progressively perturbs the data by adding Gaussian noise in small increments. The simplest discrete-time version can be written as
$x_{t}=\sqrt{\alpha_{t}}\,x_{t-1}+\sqrt{1-\alpha_{t}}\,\epsilon_{t},\qquad\epsilon_{t}\sim\mathcal{N}(0,I),$ (1)
where $\alpha_{t}=1-\beta_{t}$ and $\beta_{t}$ is a small variance parameter controlling the noise level at step $t$. Equation (1) defines a time-inhomogeneous Markov chain that slowly “forgets” the initial sample $x_{0}$ as $t$ increases.
Iterating (1) yields an explicit marginal form:
$x_{t}=\sqrt{\bar{\alpha}_{t}}\,x_{0}+\sqrt{1-\bar{\alpha}_{t}}\,\epsilon,\qquad\epsilon\sim\mathcal{N}(0,I),$ (2)
where $\bar{\alpha}_{t}=\prod_{s=1}^{t}\alpha_{s}$. As $t$ grows, $\bar{\alpha}_{t}\to 0$ and $x_{t}$ converges in distribution to a standard Gaussian. Thus, forward diffusion transforms any complex data distribution into a tractable high-entropy Gaussian reference—an essential property for generative modeling.
Continuous-Time Limit.  When the discrete time steps become infinitesimal, the Markov chain converges to a continuous diffusion process governed by the Ornstein–Uhlenbeck SDE [21]:
$\mathrm{d}x_{s}=-\tfrac{1}{2}\beta(s)\,x_{s}\,\mathrm{d}s+\sqrt{\beta(s)}\,\mathrm{d}W_{s},$ (3)
where $W_{s}$ is a standard Brownian motion. The drift term $\tfrac{1}{2}\beta(s)x_{s}$ gently pulls the state toward the origin, while the diffusion term injects Gaussian noise at rate $\beta(s)$. The associated Fokker–Planck equation shows that the density of $x_{s}$ converges to the standard normal distribution as $s\to T$.
II-B Reverse Process: From Noise Back to Data
The generative phase inverts the above diffusion. Intuitively, if the forward process gradually corrupted the data with Gaussian noise, then the reverse process must remove that noise in a statistically optimal way. The reverse-time dynamics can be derived rigorously using time-reversal results for diffusion processes (the Haussmann–Pardoux theorem [22]), yielding:
$\mathrm{d}x_{s}=\Bigl[-\tfrac{1}{2}\beta(s)\,x_{s}-\beta(s)\nabla\log p_{s}(x_{s})\Bigr]\,\mathrm{d}s+\sqrt{\beta(s)}\,\mathrm{d}\overleftarrow{W}_{s},$ (4)
where $p_{s}(x)$ is the marginal density of $x_{s}$, and $\overleftarrow{W}_{s}$ denotes Brownian motion adapted to the reverse filtration.
The additional term $\beta(s)\nabla\log p_{s}(x_{s})$ represents a data-dependent correction that exactly counteracts the injected noise. This score function $\nabla\log p_{s}(x)$ is generally unknown but can be estimated by a neural network trained to predict the gradient of the log-density. The resulting framework, known as a score-based generative model, unifies DDPMs and stochastic differential equation (SDE) methods [3, 23].
In generation mode, the learned network drives the reverse SDE (4) starting from Gaussian noise, producing high-dimensional samples that follow the learned data distribution.
II-C Information-Theoretic Viewpoint
The forward SDE (3) induces a density $p_{t}(x)$ satisfying the Fokker–Planck equation:
$\frac{\partial p_{t}}{\partial t}=\frac{1}{2}\nabla\cdot\!\big(\beta(t)x\,p_{t}\big)+\frac{1}{2}\beta(t)\,\Delta p_{t}.$ (5)
This equation can be viewed as a gradient flow of entropy in the $L_{2}$-Wasserstein space of probability measures. Indeed, the Shannon differential entropy
$H(p_{t})=-\int p_{t}(x)\log p_{t}(x)\,\mathrm{d}x$
increases monotonically over time according to the classical de Bruijn identity:
$\frac{\mathrm{d}}{\mathrm{d}t}H(p_{t})=\frac{1}{2}\,\beta(t)\,\mathcal{I}(p_{t}),$ (6)
where $\mathcal{I}(p_{t})=\int\|\nabla\log p_{t}(x)\|^{2}p_{t}(x)\,\mathrm{d}x$ is the Fisher information. Thus, the forward diffusion performs a gradient ascent on entropy, continuously spreading probability mass until it reaches the Gaussian equilibrium. Conversely, the reverse process (4) can be viewed as a steepest descent of entropy, guiding the system back toward a structured, low-entropy data distribution.
This duality between entropy production (forward) and entropy dissipation (reverse) provides a geometric and information-theoretic foundation for diffusion-based generative modeling. It also sets the stage for the non-commutative generalization developed later in this paper, where Shannon entropy and Fisher information are replaced by their free-probabilistic counterparts.
II-D Applications and Impact
Diffusion models have transformed several areas of artificial intelligence:
  • Image, audio, and video generation: Samples from the reverse SDE (4) achieve state-of-the-art realism in large-scale models such as Imagen, DALL-E 2, and Stable Diffusion [24, 25, 26].
  • Conditional generation: Conditioning mechanisms, such as classifier guidance, enable controllable text-to-image or text-to-audio synthesis.
  • Uncertainty quantification: Score-based methods provide flexible tools for approximate Bayesian inference and posterior sampling [4].
  • Scientific and physical modeling: Diffusion-based SDEs have been used to generate molecular conformations, simulate turbulent flows, and model protein structures.
From a theoretical standpoint, the universality of diffusion models lies in their two complementary limits: the forward diffusion maps any data distribution to a Gaussian prior, while the reverse process—given an accurate score estimate—can recover arbitrarily complex distributions. This classical framework will serve as the benchmark for the free diffusion models developed in subsequent sections, where we extend these principles to non-commutative, operator-valued random variables.
III Free Probability Formulation
Classical diffusion-based generative models operate in Euclidean probability spaces, where random variables commute and independence is described by factorization of expectations. However, in many high-dimensional or structured learning systems—such as deep neural networks, random feature maps, or transformer attention operators—the effective random objects are not scalars but matrices or operators. Their entries are highly correlated, yet their large-dimensional behavior often exhibits a form of asymptotic freeness rather than classical independence. This observation motivates the use of free probability theory as a natural non-commutative extension of diffusion modeling.
Free probability, introduced by Voiculescu [11, 12], provides an operator-algebraic framework for describing random variables that do not commute but still exhibit a form of “statistical independence” called freeness. It has deep connections to random matrix theory, where empirical spectral distributions converge to free convolutions in the large-dimensional limit [15]. In this setting, diffusion can be viewed not as the evolution of a density in $\mathbb{R}^{d}$, but as the evolution of a spectral measure under free additive convolution.
III-A Free Probability Background
Let $(\mathcal{A},\varphi)$ be a tracial $W^{*}$-probability space, where $\mathcal{A}$ is a von Neumann algebra acting on a Hilbert space, and $\varphi\mathrel{\mathop{\ordinarycolon}}\mathcal{A}\to\mathbb{C}$ is a faithful, normal, tracial state. Elements of $\mathcal{A}$ play the role of (possibly non-commuting) random variables. For a self-adjoint element $X\in\mathcal{A}$, the spectral theorem associates a probability measure $\mu_{X}$ on $\mathbb{R}$, called the spectral law of $X$, such that
$\varphi(f(X))=\int_{\mathbb{R}}f(\lambda)\,\mathrm{d}\mu_{X}(\lambda)\quad\text{for every bounded Borel function }f.$
This spectral law generalizes the notion of a probability density to the non-commutative world.
III-B Freeness
Freeness generalizes classical independence by eliminating mixed moments. Subalgebras $\mathcal{A}_{1},\ldots,\mathcal{A}_{m}\subseteq\mathcal{A}$ are said to be freely independent (or simply free) if
$\varphi(a_{1}a_{2}\cdots a_{k})=0$
whenever $a_{j}\in\mathcal{A}_{i_{j}}$, $\varphi(a_{j})=0$, and consecutive indices differ ($i_{1}\neq i_{2}\neq\cdots\neq i_{k}$). This property means that alternating, centered elements from distinct subalgebras have zero expectation—the non-commutative analogue of the classical product rule $\mathbb{E}[XY]=\mathbb{E}[X]\mathbb{E}[Y]$ for independent variables.
III-C Free Brownian Motion and Semicircular Law
A free Brownian motion (or semicircular process) is a family $\{S_{t}\mathrel{\mathop{\ordinarycolon}}t\geq 0\}$ of self-adjoint operators satisfying:
  1. 1.
    $S_{0}=0$;
  2. 2.
    the increments $S_{t_{j}}-S_{t_{j-1}}$ are freely independent for $t_{1}<\cdots<t_{n}$;
  3. 3.
    each increment $S_{t}-S_{s}$ has a semicircular law with density
    $\mathrm{d}\mu_{S_{t}-S_{s}}(x)=\frac{1}{2\pi(t-s)}\sqrt{4(t-s)-x^{2}}\,\mathbf{1}_{[-2\sqrt{t-s},\,2\sqrt{t-s}]}(x)\,\mathrm{d}x,$
    so that $\varphi((S_{t}-S_{s})^{2})=t-s$ and all higher free cumulants vanish except the second.
In analogy with classical Brownian motion, the free process $\{S_{t}\}$ describes “semicircular noise” whose increments are freely independent rather than classically independent.
III-D Free Diffusion as a Model of Operator-Valued Noise
Consider now a non-commutative diffusion process
$\mathrm{d}X_{t}=b_{t}\,\mathrm{d}t+\sigma_{t}\,\mathrm{d}S_{t},$
where $X_{t},b_{t},\sigma_{t}\in\mathcal{A}_{\mathrm{sa}}$ and $S_{t}$ is free Brownian motion. The spectral measure $\mu_{X_{t}}$ evolves by free additive convolution—a non-commutative analogue of Gaussian convolution. When $b_{t}=-\tfrac{1}{2}\beta(t)X_{t}$ and $\sigma_{t}=\sqrt{\beta(t)}$, this becomes a free Ornstein–Uhlenbeck process [19], whose stationary law is the semicircular distribution.
This provides a direct parallel to the forward process in classical diffusion models: the random matrix (or operator) $X_{t}$ becomes increasingly “semicircular” as $t$ increases, corresponding to entropy growth under the free Fokker–Planck equation. The reverse process, derived in Section VI, then reconstructs the initial spectral distribution by introducing the conjugate variable $\Xi_{t}$, which acts as the free analogue of the score function.
III-E Example: Free DDPM for Random Matrix Data
To illustrate the connection between diffusion models and free probability, consider a random matrix analogue of a classical denoising diffusion process. Let $X_{0}^{(N)}\in\mathbb{C}^{N\times N}$ be a Hermitian data matrix—such as the empirical covariance or attention kernel from a neural network layer—whose eigenvalue distribution represents the structural information to be learned. The forward “noising” process is defined as
$X_{t}^{(N)}=\sqrt{\alpha_{t}}\,X_{0}^{(N)}+\sqrt{1-\alpha_{t}}\,Z_{t}^{(N)},\qquad Z_{t}^{(N)}\sim\frac{1}{\sqrt{N}}\,G_{t},$
where $G_{t}$ has i.i.d. standard Gaussian entries and $\alpha_{t}=\exp(-\!\int_{0}^{t}\beta(s)\,\mathrm{d}s)$ is a prescribed variance schedule. The term $Z_{t}^{(N)}$ introduces isotropic matrix-valued noise, while $\beta(t)$ controls the diffusion rate.
As the dimension $N\to\infty$, Voiculescu’s asymptotic freeness theorem implies that the random matrices $X_{t}^{(N)}$ and $Z_{t}^{(N)}$ become freely independent in the sense of moments. Consequently, the empirical spectral distribution of $X_{t}^{(N)}$ converges to that of an operator-valued process $X_{t}$ satisfying the free Ornstein–Uhlenbeck equation
$\mathrm{d}X_{t}=-\tfrac{1}{2}\beta(t)X_{t}\,\mathrm{d}t+\sqrt{\beta(t)}\,\mathrm{d}S_{t},$
where $S_{t}$ is a free Brownian motion on a tracial $W^{*}$-probability space $(\mathcal{A},\varphi)$. The spectral measure $\mu_{X_{t}}$ evolves by free additive convolution:
$\mu_{X_{t}}=\mu_{X_{0}}\boxplus\sigma_{\Lambda(t)},\quad\Lambda(t)=\int_{0}^{t}\beta(s)\,\mathrm{d}s,$
and approaches the semicircular law $\sigma$ as $t\to\infty$.
In this interpretation, the forward free diffusion gradually erases the spectral structure of the data by adding freely independent “semicircular noise”—the free analogue of Gaussian perturbations in Euclidean DDPMs. The reverse-time process, driven by the conjugate variable $\Xi_{t}=\nabla_{X_{t}}\chi(X_{t})$, acts as a free denoiser: it reconstructs the original spectral law $\mu_{X_{0}}$ by performing free deconvolution with respect to $\sigma_{\Lambda(t)}$.
Practically, one can view a free DDPM as a generative model operating on spectra of large matrices rather than raw data vectors. During “training,” the model learns an approximation to the conjugate variable $\Xi_{t}$, analogous to the score function in classical diffusion models. At generation time, the learned reverse dynamics sample new operator-valued data by evolving from the high-entropy semicircular prior back toward a structured spectral law.
This free-probabilistic viewpoint thus provides a principled description of how diffusion mechanisms might act on non-commuting or operator-valued representations—such as large attention matrices, kernel operators, or graph Laplacians—where classical independence assumptions no longer hold.
III-F Summary
Free probability thus provides a rigorous language for describing diffusion in non-commutative or matrix-valued settings. The forward process corresponds to the flow of spectral measures under free additive convolution, while the reverse process represents an entropy-gradient descent in the geometry of free Wasserstein space. This framework extends the statistical mechanics interpretation of DDPMs to the realm of operator algebras, offering a foundation for analyzing and potentially designing generative models for structured, non-commuting data.
IV Forward Diffusion as a Free Ornstein–Uhlenbeck Process
Proposition 1 (Free forward SDE).
The forward diffusion in the free probability setting is governed by the stochastic differential equation
$dX_{s}=-\frac{1}{2}\beta(s)\,X_{s}\,ds+\sqrt{\beta(s)}\,dS_{s},$ (7)
where $S_{s}$ denotes a free Brownian motion (also called a semicircular process).
Proof.
Consider the discrete-time update $X_{t}=\sqrt{\alpha_{t}}\,X_{t-1}+\sqrt{1-\alpha_{t}}\,Z_{t},$ where $Z_{t}$ are centered semicircular elements, freely independent of $X_{t-1}$, and $\alpha_{t}=\exp(-\beta_{t}\Delta s)$ for a step size $\Delta s$.
As $\Delta s\to 0$, the law of $X_{t}$ evolves via a free additive convolution at each step, with a simultaneous scaling (dilation) by $\sqrt{\alpha_{t}}$. By standard arguments for free convolution semigroups and the Trotter–Kato approximation in the free setting, the infinitesimal generator of this evolution acts on noncommutative polynomials $f$ as
$\mathcal{L}_{s}f=-\frac{1}{2}\beta(s)\,\mathcal{D}f(X)\cdot X+\frac{1}{2}\beta(s)\,\Delta_{\mathrm{free}}f(X),$
where $\mathcal{D}$ denotes the free (cyclic) gradient and $\Delta_{\mathrm{free}}$ is the free Laplacian. Passing to the limit $\Delta s\to 0$ yields the stochastic differential equation (7), which is the free analogue of the classical Ornstein–Uhlenbeck equation. ∎
Remark 1.
In the commutative case, the forward diffusion corresponds to a classical Ornstein–Uhlenbeck process, whose distribution is given by a Gaussian convolution with exponentially decaying mean. In the free setting, the role of Gaussian convolution is taken by free additive convolution, and the spectral measure $\mu_{X_{s}}$ of $X_{s}$ satisfies
$\mu_{X_{s}}=D_{e^{-\frac{1}{2}\Lambda(s)}}\mu_{X_{0}}\boxplus\left(\boxplus_{0<u<s}D_{e^{-\frac{1}{2}(\Lambda(s)-\Lambda(u))}}\mu_{dS_{u}}\right),$
where $D_{c}\mu$ denotes the pushforward of $\mu$ under scaling by $c$, $\Lambda(s)=\int_{0}^{s}\beta(r)\,dr$, and $\boxplus$ denotes free additive convolution. This expression makes explicit that the forward process can be viewed as the cumulative effect of a time-rescaled semicircular noise process together with exponential damping of the initial distribution.
V Free Fokker–Planck Dynamics and Free Clark–Ocone Formula
Once the forward diffusion is lifted from classical probability to a non-commutative setting, the law of a free diffusion process $X_{s}$ evolves according to a free Fokker–Planck equation. Let $\mu_{X_{s}}$ denote the spectral distribution of $X_{s}$ in a tracial $W^{*}$-probability space $(\mathcal{A},\varphi)$, with dynamics
$\mathrm{d}X_{s}=-\tfrac{1}{2}\beta(s)X_{s}\,\mathrm{d}s+\sqrt{\beta(s)}\,\mathrm{d}S_{s},$
where $(S_{s})_{s\geq 0}$ is a free Brownian motion [13]. Then $\mu_{X_{s}}$ satisfies, in the distributional sense,
$\frac{\partial}{\partial s}\mu_{X_{s}}(x)=\frac{1}{2}\beta(s)\frac{\partial}{\partial x}\big(x\mu_{X_{s}}(x)\big)+\frac{1}{2}\beta(s)\frac{\partial^{2}}{\partial x^{2}}\mu_{X_{s}}(x),$ (8)
which is a free counterpart of the classical Fokker–Planck equation. Equation (8) governs the evolution of the spectral law, and can be shown to correspond to a gradient flow of Voiculescu’s free entropy $\chi(\mu_{X_{s}})$ in the metric space of spectral distributions endowed with the free 2-Wasserstein distance [14, 27].
The derivation of the reverse-time SDE in the free setting requires conditioning with respect to the free filtration and exploiting the free analogue of the martingale representation theorem. This role is played by the Free Clark–Ocone formula.
Lemma 1 (Free Clark–Ocone formula).
Let $(S_{t})_{t\geq 0}$ be a free Brownian motion on $(\mathcal{A},\varphi)$, and let $F\in L^{2}(\mathcal{A},\varphi)$ be square-integrable with respect to the non-commutative filtration $(\mathcal{F}_{t})$ generated by $S_{t}$. Then there exists a unique adapted process $\mathbb{E}[\mathcal{D}_{u}F\mid\mathcal{F}_{u}]$ such that
$F=\mathbb{E}[F]+\int_{0}^{T}\mathbb{E}[\mathcal{D}_{u}F\mid\mathcal{F}_{u}]\,\mathrm{d}S_{u},$ (9)
where $\mathcal{D}_{u}$ denotes the free Malliavin derivative. This decomposition is the non-commutative analogue of the classical Clark–Ocone representation [14, 28].
Lemma 1 is the cornerstone of free stochastic calculus: it allows one to express functionals of the free Brownian motion as stochastic integrals, which in turn is essential for deriving the drift correction term in the reverse-time SDE. In later sections, this result will be applied to $\log p_{s}(X_{s})$ to obtain the reverse-time drift in terms of the conjugate variable $\Xi_{s}$, a free analogue of the classical score function.
VI Reverse-Time Free SDE: Malliavin and PDE Derivations
We now derive the reverse-time dynamics of a free diffusion process using two complementary approaches: (i) stochastic calculus with the free Clark–Ocone formula, and (ii) a direct analysis of the free Fokker–Planck equation. Both lead to the same result.
Backward Free Brownian Motion
We briefly recall the notion of a backward free Brownian motion, which we use in the formulation of the reverse-time SDE. Let $(\mathcal{A},\tau)$ be a tracial $W^{\ast}$-probability space and $(S_{t})_{t\geq 0}$ a free Brownian motion adapted to the filtration $(\mathcal{F}_{t})_{t\geq 0}$. For a fixed horizon $T>0$, define the reversed filtration by
$\mathcal{G}_{s}\mathrel{\mathop{\ordinarycolon}}=\mathcal{F}_{T-s},\quad 0\leq s\leq T.$
A process $(\widetilde{S}_{s})_{0\leq s\leq T}$ is called a backward free Brownian motion if:
  1. 1.
    $(\widetilde{S}_{s})$ is adapted to $(\mathcal{G}_{s})$ and is a martingale with respect to $(\mathcal{G}_{s})$,
  2. 2.
    $\widetilde{S}_{0}=0$ and $\widetilde{S}_{s}-\widetilde{S}_{r}$ is $\ast$-free from $\mathcal{G}_{r}$ for $0\leq r<s\leq T$,
  3. 3.
    the increments $\widetilde{S}_{s}-\widetilde{S}_{r}$ are semicircular with variance (sr)(s-r), and
  4. 4.
    the quadratic variation satisfies
    $\langle\widetilde{S}\rangle_{s}=s\,1\quad\text{for all }0\leq s\leq T.$
The construction of such processes, together with the free stochastic calculus required for handling them, can be found in the work of Biane and Speicher [14]. This provides the canonical noise process driving the reverse-time dynamics in the free setting.
Remark 2.
The role of the backward free Brownian motion is entirely analogous to that of the backward Itô Brownian motion in the classical Haussmann–Pardoux time-reversal framework. In particular, it provides the martingale noise term that compensates the drift correction in the reverse-time SDE (R). Thus, once the forward dynamics (F) are fixed, the pair consisting of the reversed drift and the backward free Brownian motion uniquely determines the law of the reversed process.
Before turning to the derivation of the reverse SDE, we recall a key analytical identity connecting the free Malliavin derivative, the conjugate variable, and the generator of the free Ornstein–Uhlenbeck semigroup. This identity generalizes Voiculescu’s polynomial integration-by-parts formula to the full Sobolev domain and will serve as the bridge between the Malliavin and PDE derivations below.
Lemma 2 (Free integration-by-parts identity beyond polynomial test functions).
Let $(X_{t})_{t\geq 0}$ be a free Ornstein–Uhlenbeck process on a tracial $W^{*}$-probability space $(\mathcal{A},\tau)$,
$dX_{t}=-\tfrac{1}{2}X_{t}\,dt+dS_{t},$
and let $\mu_{t}$ denote its spectral law. Denote by $\Xi_{t}$ the conjugate variable of $\mu_{t}$ with respect to $X_{t}$, defined via Voiculescu’s free Fisher information
$\Phi^{*}(\mu_{t})=\|\,\Xi_{t}+\tfrac{1}{2}X_{t}\,\|_{L^{2}(\mu_{t})}^{2}.$
Then for every $F\in\mathrm{Dom}(D)\subset L^{2}(\mathcal{A},\tau)$ whose free Malliavin derivative $DF\in L^{2}(\mathcal{A}\otimes\mathcal{A},\tau\otimes\tau)$ exists and satisfies $\|DF\|_{L^{2}}<\infty$, the following integration-by-parts formula holds:
$\mathbb{E}_{\tau}\!\left[\langle DF,1\rangle_{\mathrm{free}}\right]=\mathbb{E}_{\tau}\!\left[F\big(\Xi_{t}+\tfrac{1}{2}X_{t}\big)\right],$ (10)
where $\langle\cdot,\cdot\rangle_{\mathrm{free}}$ denotes the Hilbert–Schmidt inner product on $L^{2}(\mathcal{A}\otimes\mathcal{A},\tau\otimes\tau)$. Moreover, for all $F,G\in\mathrm{Dom}(D)$, the covariance identity
$\mathrm{Cov}_{\tau}(F,G)=\mathbb{E}_{\tau}\!\left[\langle DF,DG\rangle_{\mathrm{free}}\right]=\mathbb{E}_{\tau}\!\left[F\big(\Xi_{t}+\tfrac{1}{2}X_{t}\big)G\right]-\mathbb{E}_{\tau}[F]\,\mathbb{E}_{\tau}[G]$ (11)
extends Voiculescu’s classical free integration-by-parts identity from non-commutative polynomials to the full Sobolev space $W^{1,2}_{\mathrm{free}}(\mu_{t})$.
Proof.
Let $F\in\mathrm{Dom}(D)$ and approximate it by a sequence of self-adjoint non-commutative polynomials $\{P_{n}(X_{t})\}_{n\geq 1}$ such that $P_{n}(X_{t})\to F$ in $L^{2}(\mathcal{A},\tau)$ and $DP_{n}(X_{t})\to DF$ in $L^{2}(\mathcal{A}\otimes\mathcal{A})$. For each polynomial $P_{n}$, Voiculescu’s original integration-by-parts formula [12, Prop. 7.1] gives
$\tau\!\big(P_{n}^{\prime}(X_{t})\big)=\tau\!\big(P_{n}(X_{t})(\Xi_{t}+\tfrac{1}{2}X_{t})\big).$
Passing to the limit $n\to\infty$ and using the continuity of the trace pairing $\tau(ab)=\mathbb{E}_{\tau}[ab]$ yields (10).
The covariance identity (11) follows from the product rule of the free Malliavin derivative, $D(FG)=(DF)(1\!\otimes\!G)+(1\!\otimes\!F)(DG)$, together with the $L^{2}$-duality between $D$ and its adjoint, as established in [14, Thm. 3.4]. The argument extends from polynomials to $W^{1,2}_{\mathrm{free}}(\mu_{t})$ by density and closedness of $D$ in $L^{2}(\mathcal{A},\tau)$, proving the lemma. ∎
We now employ this identity to identify the conjugate variable $\Xi_{t}$ as the adjoint of the Malliavin derivative and to construct the drift correction appearing in the reverse-time SDE.
Lemma 3 (Conditional Malliavin derivative and conjugate variable).
Let $(\mathcal{A},\varphi)$ be a tracial $W^{\ast}$-probability space and $X_{t}\in\mathcal{A}_{\mathrm{sa}}$ a self-adjoint random variable whose law $\mu_{X_{t}}$ admits a conjugate variable $\Xi_{t}\in L^{2}(W^{\ast}(X_{t}),\varphi)$. Then for any noncommutative polynomial $P$,
$\mathbb{E}_{\varphi}[D_{t}P(X_{t})\mid\mathcal{F}_{t}]=P^{\prime}(X_{t})=\mathrm{proj}_{L^{2}(W^{\ast}(X_{t}),\,\varphi)}\!\big(\Xi_{t}P(X_{t})\big),$
where $D_{t}$ denotes the free Malliavin derivative at time $t$ and $\mathcal{F}_{t}=\sigma\{S_{u}\mathrel{\mathop{\ordinarycolon}}0\leq u\leq t\}$ is the forward filtration. In particular, for the log-density functional $G=\log\pi_{t}(X_{t})$,
$\mathbb{E}_{\varphi}[D_{t}G\mid\mathcal{F}_{t}]=-\,\Xi_{t},\qquad\Phi^{\ast}(X_{t})=\|\Xi_{t}\|_{2}^{2}.$
Proof.
For each polynomial $P$, $P(X_{t})$ belongs to $\mathrm{Dom}(D_{t})$, and the free chain rule [14] yields $D_{t}P(X_{t})=P^{\prime}(X_{t})$. Taking conditional expectation onto $\mathcal{F}_{t}$ gives $\mathbb{E}_{\varphi}[D_{t}P(X_{t})\mid\mathcal{F}_{t}]=P^{\prime}(X_{t})$.
By the defining property of the conjugate variable, for all noncommutative polynomials $Q$,
$\varphi\!\big(Q(X_{t})P^{\prime}(X_{t})\big)=\varphi\!\big(Q(X_{t})\Xi_{t}P(X_{t})\big).$
Since $L^{2}(W^{\ast}(X_{t}),\varphi)$ is the closure of such polynomials, this equality of inner products implies that $P^{\prime}(X_{t})$ is the $L^{2}$-orthogonal projection of $\Xi_{t}P(X_{t})$ onto $L^{2}(W^{\ast}(X_{t}),\varphi)$. Finally, applying this to the log-density functional $G=\log\pi_{t}(X_{t})$ and using the definition of $\Phi^{\ast}$ via free Fisher information yields $\mathbb{E}_{\varphi}[D_{t}G\mid\mathcal{F}_{t}]=-\Xi_{t}$. ∎
Theorem 1 (Reverse-time free SDE).
Let $(\mathcal{A},\varphi)$ be a tracial $W^{\ast}$-probability space, let $(S_{t})_{t\geq 0}$ be a free Brownian motion, and fix $T>0$. Let $\beta\in C^{1}([0,T];(0,\infty))$ be a deterministic variance schedule. Consider the forward free Ornstein–Uhlenbeck SDE
$\mathrm{d}X_{t}=-\tfrac{1}{2}\beta(t)X_{t}\,\mathrm{d}t+\sqrt{\beta(t)}\,\mathrm{d}S_{t},\qquad X_{0}\in\mathcal{A}_{\mathrm{sa}},$ (F)
where $\mu_{X_{0}}$ is compactly supported with $\chi(\mu_{X_{0}})>$-$\infty$ and $\Phi^{\ast}(\mu_{X_{0}})<\infty$. Assume that $(X_{t})_{0\leq t\leq T}$ exists uniquely in the free Itô sense, and that each $\mu_{X_{t}}$ admits a conjugate variable $\Xi_{t}\in L^{2}(W^{\ast}(X_{t}),\varphi)$. Then the time-reversed process $\widetilde{X}_{s}\mathrel{\mathop{\ordinarycolon}}=X_{T-s}$ satisfies the stochastic differential equation
$\mathrm{d}\widetilde{X}_{s}=\Bigl(-\tfrac{1}{2}\beta(T-s)\widetilde{X}_{s}-\beta(T-s)\,\Xi_{T-s}\Bigr)\,\mathrm{d}s+\sqrt{\beta(T-s)}\,\mathrm{d}\widetilde{S}_{s},$ (R)
where $(\widetilde{S}_{s})_{s\in[0,T]}$ is a backward free Brownian motion adapted to the reversed filtration $\widetilde{\mathcal{F}}_{s}=\sigma\{\widetilde{X}_{u}\mathrel{\mathop{\ordinarycolon}}0\leq u\leq s\}$.
Proof.
Step 1: Existence and forward representation. Let $\Lambda(t)=\int_{0}^{t}\beta(r)\,\mathrm{d}r$. By free Itô integration theory [14], the SDE (F) admits the explicit solution
$X_{t}=e^{-\frac{1}{2}\Lambda(t)}X_{0}+\!\int_{0}^{t}e^{-\frac{1}{2}(\Lambda(t)-\Lambda(u))}\sqrt{\beta(u)}\,\mathrm{d}S_{u}.$
The free Itô isometry ensures that $\|X_{t}\|_{2}^{2}<\infty$ and thus $X_{t}\in L^{2}(\mathcal{A},\varphi)$ for all $t\in[0,T]$.
Step 2: Clark–Ocone representation. For each fixed $T$, the free Clark–Ocone formula (see Appendix C for a rigorous derivation) gives an adapted representation of $X_{T}$:
$X_{T}=\varphi(X_{T})+\int_{0}^{T}u_{t}\,\#\,\mathrm{d}S_{t},\qquad u_{t}=\mathbb{E}_{\varphi}[D_{t}X_{T}\mid\mathcal{F}_{t}].$
Substituting the explicit solution of $X_{T}$ into the stochastic integral and applying the free Itô isometry yields
$u_{t}=e^{-\frac{1}{2}(\Lambda(T)-\Lambda(t))}\sqrt{\beta(t)},\qquad t\in[0,T].$
This follows because $(\mathrm{d}S_{t})_{t\geq 0}$ has freely independent increments and $u_{t}$ is adapted; hence the conditional expectation $\mathbb{E}_{\varphi}[\cdot\mid\mathcal{F}_{t}]$ eliminates all stochastic terms beyond time $t$, leaving only the deterministic damping factor $e^{-\frac{1}{2}(\Lambda(T)-\Lambda(t))}$.
Step 3: Identification of the drift correction. By Lemma 3, for the log-density functional $G=\log\pi_{t}(X_{t})$ one has $\mathbb{E}_{\varphi}[D_{t}G\mid\mathcal{F}_{t}]=-\,\Xi_{t}$. Combining this with the derivative term $\tfrac{1}{2}\beta(t)X_{t}$ in the forward SDE (F) gives the backward drift $\tfrac{1}{2}\beta(t)X_{t}-\beta(t)\Xi_{t}$.
Step 4: Time reversal. Define $\widetilde{X}_{s}=X_{T-s}$. By free stochastic calculus with respect to backward filtrations (see Biane [29], Sec. 3), the differential of $\widetilde{X}_{s}$ satisfies
$\mathrm{d}\widetilde{X}_{s}=\Bigl(-\tfrac{1}{2}\beta(T-s)\widetilde{X}_{s}-\beta(T-s)\,\Xi_{T-s}\Bigr)\mathrm{d}s+\sqrt{\beta(T-s)}\,\mathrm{d}\widetilde{S}_{s},$
where $\widetilde{S}_{s}\mathrel{\mathop{\ordinarycolon}}=S_{T}-S_{T-s}$ is a free Brownian motion under the reversed filtration. The integrability of the drift and diffusion coefficients follows from the square-integrability of $X_{t}$ and $\Xi_{t}$, ensuring well-posedness of (R) in $L^{2}(\mathcal{A},\varphi)$.
Hence, the time-reversed process $\widetilde{X}_{s}$ is a weak solution of (R), completing the proof. ∎
Completion of the proof. To make the time-reversal argument rigorous, recall that under the free Itô isometry [14, Prop. 3.2],
$\tau\!\left[\left(\int_{0}^{t}A_{s}\,\#\,dS_{s}\right)\!\left(\int_{0}^{t}B_{s}\,\#\,dS_{s}\right)^{\ast}\right]=\int_{0}^{t}\tau[A_{s}B_{s}^{\ast}]\,ds,$
for all square-integrable adapted processes $A_{s},B_{s}$. Since $dS_{t}$ has freely independent centered increments and $u_{t}$ is adapted, the conditional expectation $\mathbb{E}[D_{t}X_{T}\mid\mathcal{F}_{t}]$ isolates precisely the exponential damping factor $e^{-\frac{1}{2}(\Lambda(T)-\Lambda(t))}\sqrt{\beta(t)}$. Moreover, the covariance structure $\tau[(S_{t}-S_{s})^{2}]=t-s$ implies that the backward process $\widetilde{S}_{s}=S_{T}-S_{T-s}$ satisfies $\tau[(\widetilde{S}_{t}-\widetilde{S}_{s})^{2}]=t-s$ and has freely independent increments with respect to the reversed filtration. Hence $\widetilde{S}_{s}$ is a backward free Brownian motion in the sense of [13]. Substituting $dS_{t}=d\widetilde{S}_{T-t}+\sqrt{\beta(t)}\,u_{t}\,dt$ into the stochastic integral and using adjointness of $D_{t}$ and $\delta$ (see Appendix C) yields the reverse SDE (eqn. (R)) rigorously.
Remark 3.
Both derivations show that the term $\Xi_{s}$ plays the role of the non-commutative score function in the reverse-time dynamics, and that the reversed diffusion acts as a free deconvolution flow.
Corollary: Reverse Drift as a Gradient of Free Entropy
Corollary 1.
Let $\chi(\mu)$ denote Voiculescu’s free entropy functional defined on the spectral distribution $\mu_{X_{s}}$, and let $\Xi_{s}$ be the conjugate variable of $X_{s}$. Then
$\tfrac{1}{2}\beta(s)X_{s}-\beta(s)\Xi_{s}=-\beta(s)\nabla\chi[\mu_{X_{s}}](X_{s}),$ (12)
where $\nabla\chi[\mu_{X_{s}}](X_{s})=\frac{1}{2}X_{s}+\Xi_{s}$ is the $L^{2}$-gradient of $\chi$.
Proof.
The variational derivative of $\chi$ implies $\nabla\chi[\mu_{X_{s}}](X_{s})=\frac{1}{2}X_{s}+\Xi_{s},$ and substitution into the reverse drift term gives the stated expression. ∎
Remark 4.
This shows that the reverse-time free diffusion integrates the negative gradient of the free entropy functional, making it a steepest-descent flow on the free entropy landscape.
Detailed proofs of the free de Bruijn identity, Fisher–information flow, and the associated HWI-type inequalities that underlie the information interpretation of the reverse dynamics are provided in Appendix D.
VII Free Optimal Transport, the JKO Scheme, and a Candidate Gradient Flow Functional
An important extension of the free diffusion framework is to interpret the evolution of spectral distributions as a gradient flow in the space of probability measures endowed with a free Wasserstein metric. In the commutative case, the seminal work of Jordan, Kinderlehrer, and Otto [30] introduced the JKO scheme, where solutions to the Fokker–Planck equation are obtained as the limit of a sequence of variational problems in the $L^{2}$ Wasserstein metric. The monograph of Ambrosio, Gigli and Savaré [31] provides the modern rigorous formulation of these ideas. We propose that an analogous construction is valid in the noncommutative setting.
VII-A A free JKO scheme.
Let $(\mathcal{P}_{2}^{\mathrm{free}},W_{2,\mathrm{free}})$ be the space of compactly supported spectral measures equipped with the free $L^{2}$ Wasserstein distance. Given a free energy functional $\mathcal{F}[\mu]$, let us define the time-discrete scheme
$\mu_{k+1}=\arg\min_{\nu\in\mathcal{P}_{2}^{\mathrm{free}}}\left\{\frac{1}{2\tau}W_{2,\mathrm{free}}^{2}(\nu,\mu_{k})+\mathcal{F}[\nu]\right\},$ (13)
where $\tau>0$ is the time step. As $\tau\to 0$, the piecewise-constant interpolation of $(\mu_{k})$ is expected to converge to the solution $\mu_{t}$ of the free gradient flow
$\partial_{t}\mu_{t}+\nabla\cdot\left(\mu_{t}\nabla\frac{\delta\mathcal{F}}{\delta\mu_{t}}\right)=0,$ (14)
where the variational derivative $\frac{\delta\mathcal{F}}{\delta\mu}$ is interpreted via Voiculescu’s free calculus.
VII-B A candidate functional for free Wasserstein flows.
Inspired by Voiculescu’s free entropy [32] and the classical logarithmic energy for random matrices, a natural choice for $\mathcal{F}$ is
$\mathcal{F}[\mu]=\int_{\mathbb{R}}\frac{1}{2}x^{2}\,d\mu(x)-\chi[\mu],\qquad\chi[\mu]=\iint\log|x-y|\,d\mu(x)d\mu(y),$ (15)
where the first term enforces confinement and the second term promotes dispersion of eigenvalues. The unique minimizer of $\mathcal{F}$ is the semicircular law.
The first variation of $\mathcal{F}$ is
$\frac{\delta\mathcal{F}}{\delta\mu}(x)=\frac{1}{2}x^{2}-2\int\log|x-y|\,d\mu(y),$ (16)
and substituting into the gradient flow equation yields
$\partial_{t}\mu_{t}=\nabla\cdot\left(\mu_{t}\nabla\left[\frac{1}{2}x^{2}-2\int\log|x-y|\,d\mu_{t}(y)\right]\right),$ (17)
which can be interpreted as the free analogue of the Fokker–Planck equation for the forward diffusion.
VII-C On the $\Gamma$-limit of the free JKO functionals
We work on the space $\mathcal{P}_{2}^{\mathrm{free}}$ of compactly supported probability measures on $\mathbb{R}$ with finite second moment, endowed with the free $L^{2}$-Wasserstein distance $W_{2,\mathrm{free}}$ of Biane–Voiculescu [16]. The free energy is
$\mathcal{F}[\nu]\mathrel{\mathop{\ordinarycolon}}=\frac{1}{2}\int x^{2}\,d\nu(x)\;-\;\chi[\nu],$
where $\chi$ is Voiculescu’s (microstates) free entropy [33, 34]. For $\tau>0$ and a fixed $\mu_{k}\in\mathcal{P}_{2}^{\mathrm{free}}$ define
$\mathcal{F}_{\tau}^{k}[\nu]\;=\;\frac{1}{2\tau}W_{2,\mathrm{free}}^{2}(\nu,\mu_{k})\;+\;\mathcal{F}[\nu].$ (18)
We will use the following semicontinuity properties (proved below).
Lemma 4 (Semicontinuity for $W_{2,\mathrm{free}}$ and $\chi$).
Let $\nu_{n}\rightharpoonup\nu$ narrowly and assume $\sup_{n}\int x^{2}\,d\nu_{n}(x)<\infty$. Then:
  1. 1.
    $\liminf_{n\to\infty}W_{2,\mathrm{free}}(\nu_{n},\mu)\;\geq\;W_{2,\mathrm{free}}(\nu,\mu)$ for every fixed $\mu\in\mathcal{P}_{2}^{\mathrm{free}}$.
  2. 2.
    $\limsup_{n\to\infty}\chi[\nu_{n}]\;\leq\;\chi[\nu]$.
Consequently, the free energy functional $\mathcal{F}$ is lower semicontinuous on $\mathcal{P}_{2}^{\mathrm{free}}$ with respect to narrow convergence and has tight (hence relatively compact) sublevel sets.
Proof.
(i) Lower semicontinuity of $W_{2,\mathrm{free}}$..  Let $\nu_{n},\nu,\mu\in\mathcal{P}_{2}^{\mathrm{free}}$, and let $\Gamma(\nu_{n},\mu)$ denote the set of all admissible free couplings in the sense of Biane–Voiculescu [16], that is, tracial $W^{*}$-probability spaces $(\mathcal{A}_{n},\tau_{n})$ containing self-adjoint pairs $(X_{n},Y_{n})$ with marginal laws $\nu_{n}$ and $\mu$. The free $L_{2}$–Wasserstein distance is defined as
$W_{2,\mathrm{free}}^{2}(\nu_{n},\mu)\mathrel{\mathop{\ordinarycolon}}=\inf_{(X_{n},Y_{n})\in\Gamma(\nu_{n},\mu)}\tau_{n}\!\left[(X_{n}-Y_{n})^{2}\right].$
Let $(X_{n},Y_{n})$ be a sequence of (almost-)optimal couplings satisfying
$\tau_{n}\!\left[(X_{n}-Y_{n})^{2}\right]\leq W_{2,\mathrm{free}}^{2}(\nu_{n},\mu)+\tfrac{1}{n}.$
Since $\sup_{n}\int x^{2}\,d\nu_{n}(x)<\infty$, the laws $\{\nu_{n}\}$ are uniformly tight and have uniformly bounded second moments, hence precompact in the narrow topology. Thus there exists a subsequence $(\nu_{n_{k}})$ converging narrowly to $\nu$. For each such subsequence, the family of joint laws $\{\mathrm{Law}(X_{n_{k}},Y_{n_{k}})\}$ is also tight on $\mathbb{R}^{2}$, allowing extraction of a further subsequence (not relabeled) converging narrowly to a limiting coupling $(X,Y)$ of $\nu$ and $\mu$ in a tracial limit space (see [16, Sec. 1.3]).
The map $(x,y)\mapsto(x-y)^{2}$ is continuous and bounded below by 0, hence lower semicontinuous. Applying the Portmanteau theorem to the limit coupling yields
$\tau[(X-Y)^{2}]\leq\liminf_{k\to\infty}\tau_{n_{k}}\!\left[(X_{n_{k}}-Y_{n_{k}})^{2}\right]=\liminf_{k\to\infty}W_{2,\mathrm{free}}^{2}(\nu_{n_{k}},\mu).$
By the definition of $W_{2,\mathrm{free}}$ as the infimum over all admissible couplings, we have
$W_{2,\mathrm{free}}^{2}(\nu,\mu)\leq\tau[(X-Y)^{2}].$
Combining the above inequalities proves
$\liminf_{n\to\infty}W_{2,\mathrm{free}}(\nu_{n},\mu)\geq W_{2,\mathrm{free}}(\nu,\mu),$
which establishes the lower semicontinuity in item 1. The argument is the free-probability analogue of the classical lower semicontinuity of $W_{2}$ on $\mathcal{P}_{2}(\mathbb{R})$, see also [31, Ch. 2].
(ii) Upper semicontinuity of $\chi$.  For a compactly supported probability measure $\nu$, recall that Voiculescu’s microstates free entropy is defined by
$\chi[\nu]\mathrel{\mathop{\ordinarycolon}}=\inf_{R>0}\,\limsup_{m\to\infty}\Bigg[\frac{1}{m^{2}}\log\Big(\mathrm{Vol}\big\{A\in M_{m}^{\mathrm{sa}}(\mathbb{C})\mathrel{\mathop{\ordinarycolon}}\|A\|\leq R,\,\lambda_{A}\approx\nu\big\}\Big)+\frac{3}{4}+\frac{1}{2}\log(2\pi)\Bigg],$
where $\lambda_{A}$ denotes the empirical spectral law of $A$. Let $\nu_{n}\rightharpoonup\nu$ narrowly with $\sup_{n}\int x^{2}\,d\nu_{n}(x)<\infty$. For each fixed $R>0$, the sets of $m\times m$ microstates $\Gamma_{m}(\nu_{n};R,\varepsilon)$ corresponding to $\nu_{n}$ converge in the Hausdorff sense (with respect to the normalized trace topology) to $\Gamma_{m}(\nu;R,\varepsilon)$ as $n\to\infty$, because narrow convergence with uniformly bounded second moments implies convergence of all *-moments of bounded degree. By the monotonicity of $\mathrm{Vol}[\Gamma_{m}(\nu;R,\varepsilon)]$ in $R$ and $\varepsilon$, and the subadditivity of $\frac{1}{m^{2}}\log\mathrm{Vol}$ in $m$, one can apply a diagonal argument in $(m,R,\varepsilon)$ to pass to the limit (see [34, Thm. 6.5.1] and [33]), obtaining
$\limsup_{n\to\infty}\chi[\nu_{n}]\leq\chi[\nu].$
Hence, the free entropy is upper semicontinuous along narrowly convergent sequences with uniformly bounded second moments.
(iii) Consequences..  Since the free energy functional $\mathcal{F}[\nu]\mathrel{\mathop{\ordinarycolon}}=\tfrac{1}{2}W_{2,\mathrm{free}}^{2}(\nu,\mu_{\mathrm{sc}})-\chi[\nu]$ is the sum of a lower semicontinuous and an upper semicontinuous term, it is itself lower semicontinuous on $\mathcal{P}_{2}^{\mathrm{free}}$. The moment bound guarantees tightness of sublevel sets of $\mathcal{F}$, which are therefore relatively compact in the narrow topology. This completes the proof. ∎
We consider the free energy
$\mathcal{F}[\nu]\;=\;\frac{1}{2}\int x^{2}\,d\nu(x)\;-\;\chi[\nu],$
defined on the space $\mathcal{P}_{2}^{\mathrm{free}}$ of compactly supported probability measures with finite second moment. We assume the semicontinuity facts collected in Lemma 4 (lower semicontinuity of $W_{2,\mathrm{free}}$, upper semicontinuity of the microstates free entropy under the indicated compactness assumptions). The goal is to show the Moreau–Yosida envelopes
$\mathcal{F}^{\tau}(\mu)\mathrel{\mathop{\ordinarycolon}}=\inf_{\nu\in\mathcal{P}_{2}^{\mathrm{free}}}\Big\{\frac{1}{2\tau}W_{2,\mathrm{free}}^{2}(\nu,\mu)+\mathcal{F}[\nu]\Big\}$ (19)
satisfy $\Gamma$-convergence $\mathcal{F}^{\tau}\xrightarrow{\Gamma}\mathcal{F}$ (narrow topology) as $\tau\downarrow 0$, and that the minimizers of $\mathcal{F}^{\tau}$ are precompact and converge to the minimizers of $\mathcal{F}$. We then explain how this controls the JKO scheme with moving the base point.
Theorem 2 ($\Gamma$-convergence of $\mathcal{F}^{\tau}$ to $\mathcal{F}$).
Assume Lemma 4 holds. For each $\tau>0$ define $\mathcal{F}^{\tau}$ by (19). Then:
  1. 1.
    For every $\mu\in\mathcal{P}_{2}^{\mathrm{free}}$,
    $\lim_{\tau\downarrow 0}\mathcal{F}^{\tau}(\mu)=\mathcal{F}(\mu).$
  2. 2.
    The functionals $\mathcal{F}^{\tau}$ $\Gamma$-converge to $\mathcal{F}$ on $\mathcal{P}_{2}^{\mathrm{free}}$ with respect to narrow convergence.
  3. 3.
    (Equi-coercivity and convergence of minimizers) For any sequence $\tau_{n}\downarrow 0$, any sequence $\mu_{n}$ with $\mathcal{F}^{\tau_{n}}(\mu_{n})\leq C$ is relatively compact (narrow). Moreover, any limit point $\mu$ satisfies $\mathcal{F}(\mu)\leq\liminf_{n}\mathcal{F}^{\tau_{n}}(\mu_{n})$. In particular, minimizers of $\mathcal{F}^{\tau}$ converge (subsequentially) to minimizers of $\mathcal{F}$.
Proof.
We will prove the three items in order. The proof strategy is standard (cf. Ambrosio–Gigli–Savaré [31, Ch. 4]), but we carefully check each step in the free setting using Lemma 4.
(I) Pointwise convergence $\mathcal{F}^{\tau}(\mu)\to\mathcal{F}(\mu)$.
For any fixed $\mu$, the choice $\nu=\mu$ in (19) yields $\mathcal{F}^{\tau}(\mu)\leq\mathcal{F}(\mu)$ for every $\tau>0$. Thus $\limsup_{\tau\downarrow 0}\mathcal{F}^{\tau}(\mu)\leq\mathcal{F}(\mu)$.
To prove the reverse inequality, fix $\mu$ and let $\nu_{\tau}$ be $\varepsilon_{\tau}$-minimizers of $\mathcal{F}^{\tau}(\mu)$, i.e.
$\mathcal{F}^{\tau}(\mu)\;\geq\;\frac{1}{2\tau}W_{2,\mathrm{free}}^{2}(\nu_{\tau},\mu)+\mathcal{F}[\nu_{\tau}]\;-\;\varepsilon_{\tau},\qquad\varepsilon_{\tau}\downarrow 0.$
We show that the transport cost term must vanish as $\tau\downarrow 0$. Assume, for contradiction, that
$\liminf_{\tau\to 0}\frac{1}{2\tau}W_{2,\mathrm{free}}^{2}(\nu_{\tau},\mu)>0.$
Then there exist $\delta>0$ and a sequence $(\tau_{k})_{k\geq 1}\downarrow 0$ such that
$\frac{1}{2\tau_{k}}W_{2,\mathrm{free}}^{2}(\nu_{\tau_{k}},\mu)\geq\delta\quad\text{for all }k.$
hence the right-hand side would be bounded away from $\mathcal{F}(\mu)$ unless $\mathcal{F}[\nu_{\tau}]\to$-$\infty$. The coercivity of $\mathcal{F}$ (the $\tfrac{1}{2}\int x^{2}$ term dominates negative entropy for compactly supported measures) prevents $\mathcal{F}[\nu_{\tau}]$ from going to -$\infty$ while keeping $W_{2,\mathrm{free}}(\nu_{\tau},\mu)$ bounded; therefore the transport term must vanish:
$\lim_{\tau\downarrow 0}\frac{1}{2\tau}W_{2,\mathrm{free}}^{2}(\nu_{\tau},\mu)=0.$
Consequently $W_{2,\mathrm{free}}(\nu_{\tau},\mu)=o(\sqrt{\tau})$ and in particular $\nu_{\tau}\rightharpoonup\mu$. By lower semicontinuity of the quadratic moment and the upper semicontinuity of $\chi$ (Lemma 4), we obtain
$\liminf_{\tau\downarrow 0}\mathcal{F}^{\tau}(\mu)\geq\liminf_{\tau\downarrow 0}\Big(\mathcal{F}[\nu_{\tau}]-\varepsilon_{\tau}\Big)\geq\mathcal{F}[\mu].$
Combining with the earlier $\limsup$ inequality yields $\lim_{\tau\downarrow 0}\mathcal{F}^{\tau}(\mu)=\mathcal{F}(\mu)$.
(II) $\Gamma$-convergence. We now establish the liminf and limsup conditions for $\Gamma$-convergence.
Liminf inequality. Let $\mu_{\tau}\rightharpoonup\mu$ narrowly. By definition,
$\mathcal{F}^{\tau}(\mu_{\tau})=\inf_{\nu}\Big\{\frac{1}{2\tau}W_{2,\mathrm{free}}^{2}(\nu,\mu_{\tau})+\mathcal{F}[\nu]\Big\}\leq\frac{1}{2\tau}W_{2,\mathrm{free}}^{2}(\nu_{0},\mu_{\tau})+\mathcal{F}[\nu_{0}],$
for any fixed competitor $\nu_{0}$. Taking $\limsup_{\tau\downarrow 0}$ and using lower semicontinuity of $W_{2,\mathrm{free}}$ (Lemma 3) yields
$\limsup_{\tau\downarrow 0}\mathcal{F}^{\tau}(\mu_{\tau})\leq\frac{1}{2\tau}W_{2,\mathrm{free}}^{2}(\nu_{0},\mu)+\mathcal{F}[\nu_{0}].$
This is not directly helpful for liminf; instead use the following classical selection: let $\nu_{\tau}$ be $\varepsilon_{\tau}$-minimizers for $\mathcal{F}^{\tau}(\mu_{\tau})$, so
$\mathcal{F}^{\tau}(\mu_{\tau})\geq\frac{1}{2\tau}W_{2,\mathrm{free}}^{2}(\nu_{\tau},\mu_{\tau})+\mathcal{F}[\nu_{\tau}]-\varepsilon_{\tau}.$
We must extract compactness for $(\nu_{\tau})$. Assume for contradiction that the sequence $(\nu_{\tau})$ is not tight: then the quadratic moment $\int x^{2}\,d\nu_{\tau}$ is unbounded and thus $\mathcal{F}[\nu_{\tau}]\to+\infty$ (coercivity), contradicting that $\mathcal{F}^{\tau}(\mu_{\tau})$ is finite. Hence $(\nu_{\tau})$ is tight and, up to a subsequence, $\nu_{\tau}\rightharpoonup\nu$. By Lemma 3 and lower semicontinuity of the moment,
$\liminf_{\tau\downarrow 0}\frac{1}{2\tau}W_{2,\mathrm{free}}^{2}(\nu_{\tau},\mu_{\tau})\geq\liminf_{\tau\downarrow 0}\frac{1}{2\tau}\big(W_{2,\mathrm{free}}(\nu,\mu)-o(1)\big)^{2}=\begin{cases}0,&\nu=\mu,\\ +\infty,&\nu\neq\mu.\end{cases}$
Thus, two possibilities arise:
(i) If $\nu\neq\mu$, then $\liminf_{\tau}\mathcal{F}^{\tau}(\mu_{\tau})=+\infty\geq\mathcal{F}(\mu)$ and the liminf inequality holds trivially.
(ii) If $\nu=\mu$, then by Lemma 3 and the lower semicontinuity of the quadratic moment,
$\liminf_{\tau\downarrow 0}\mathcal{F}^{\tau}(\mu_{\tau})\geq\liminf_{\tau\downarrow 0}\Big(\mathcal{F}[\nu_{\tau}]-\varepsilon_{\tau}\Big)\geq\mathcal{F}[\mu],$
establishing the liminf inequality in this case as well.
Hence in all cases $\liminf_{\tau\downarrow 0}\mathcal{F}^{\tau}(\mu_{\tau})\geq\mathcal{F}(\mu)$.
Limsup inequality (existence of recovery sequences). Fix $\mu$. We must construct $\mu_{\tau}\rightharpoonup\mu$ with $\limsup_{\tau}\mathcal{F}^{\tau}(\mu_{\tau})\leq\mathcal{F}(\mu)$. A trivial and effective choice is the constant sequence $\mu_{\tau}\equiv\mu$. Then
$\mathcal{F}^{\tau}(\mu_{\tau})=\mathcal{F}^{\tau}(\mu)\leq\mathcal{F}(\mu)$
(since $\nu=\mu$ is admissible in the infimum for $\mathcal{F}^{\tau}(\mu)$). Taking $\limsup$ gives $\limsup_{\tau\downarrow 0}\mathcal{F}^{\tau}(\mu_{\tau})\leq\mathcal{F}(\mu)$. This completes the limsup construction.
Thus, $\mathcal{F}^{\tau}$ $\Gamma$-converges to $\mathcal{F}$.
(III) Equi-coercivity and convergence of minimizers. Let $\tau_{n}\downarrow 0$ and $\mu_{n}$ satisfy $\mathcal{F}^{\tau_{n}}(\mu_{n})\leq C$. For each $n$, choose $\nu_{n}$ with
$\mathcal{F}^{\tau_{n}}(\mu_{n})\geq\frac{1}{2\tau_{n}}W_{2,\mathrm{free}}^{2}(\nu_{n},\mu_{n})+\mathcal{F}[\nu_{n}]-\varepsilon_{n}.$
The boundedness of $\mathcal{F}^{\tau_{n}}(\mu_{n})$ and the coercivity of $\mathcal{F}$ imply that the second moments of $\nu_{n}$ are uniformly bounded; since $W_{2,\mathrm{free}}^{2}(\nu_{n},\mu_{n})\geq 0$, it follows that the second moments of $\mu_{n}$ are also uniformly bounded (by the triangle inequality in the metric space). Hence $(\mu_{n})$ is tight and thus relatively compact in the narrow topology. By $\Gamma$-convergence and the fundamental theorem of $\Gamma$-convergence, any limit point of minimizers of $\mathcal{F}^{\tau_{n}}$ is a minimizer of $\mathcal{F}$. This proves the final claim. ∎
Consequences for the JKO/minimizing-movement scheme. Let $\tau>0$ and define the discrete JKO iteration starting from $\mu_{0}$:
$\mu_{k+1}^{\tau}\in\arg\min_{\nu\in\mathcal{P}_{2}^{\mathrm{free}}}\Big\{\frac{1}{2\tau}W_{2,\mathrm{free}}^{2}(\nu,\mu_{k}^{\tau})+\mathcal{F}[\nu]\Big\}.$ (20)
Theorems 2 ($\Gamma$-convergence of $\mathcal{F}^{\tau}$) and standard minimizing-movement compactness arguments (see Ambrosio–Gigli–Savaré [31, Ch. 4]) imply the following.
Having established the $\Gamma$-convergence of the discrete functionals $\mathcal{F}_{\tau}$ to $\mathcal{F}$, we now turn to the dynamic aspect of the minimizing-movement scheme. The next theorem shows that the sequence of discrete minimizers obtained from the JKO recursion not only exists for every time step $\tau>0$, but also converges—as $\tau\downarrow 0$—to an absolutely continuous limit curve that solves the free Wasserstein gradient flow of $\mathcal{F}$. This provides the rigorous link between the variational discretization and the constant free diffusion dynamics, closing the loop between the $\Gamma$-limit, compactness, and the evolution equation.
The next lemma refines the preceding $\Gamma$-convergence result by quantifying the stability of discrete minimizers and establishing a uniform energy–dissipation estimate at the discrete level. This provides the key a priori bounds required for the compactness and convergence arguments in the subsequent theorem.
Lemma 5 (Free $\Gamma$-convergence and Discrete Energy–Dissipation Inequality).
Let $\mathcal{F}\mathrel{\mathop{\ordinarycolon}}\mathcal{P}_{2}^{\mathrm{free}}\!\to(-\infty,+\infty]$ be a lower semicontinuous free energy functional with compact sublevel sets, and let
$\mathcal{F}_{\tau}(\mu)\mathrel{\mathop{\ordinarycolon}}=\inf_{\nu\in\mathcal{P}_{2}^{\mathrm{free}}}\Bigl\{\frac{1}{2\tau}W_{2,\mathrm{free}}^{2}(\nu,\mu)+\mathcal{F}[\nu]\Bigr\},\qquad\tau>0.$
Then:
  1. (i)
    (Free $\Gamma$-convergence) The family $\{\mathcal{F}_{\tau}\}_{\tau>0}$ $\Gamma$-converges to $\mathcal{F}$ in the narrow topology as $\tau\downarrow 0$, i.e.
[…]
  2. (ii)
    (Discrete Energy–Dissipation Inequality) Let $(\mu_{k}^{\tau})_{k\geq 0}$ satisfy the free JKO recursion
    $\mu_{k+1}^{\tau}\in\operatorname*{arg\,min}_{\nu\in\mathcal{P}_{2}^{\mathrm{free}}}\left\{\frac{1}{2\tau}W_{2,\mathrm{free}}^{2}(\nu,\mu_{k}^{\tau})+\mathcal{F}[\nu]\right\}.$
    Suppose $\mathcal{F}$ is $\lambda$-geodesically convex in $(\mathcal{P}_{2}^{\mathrm{free}},W_{2,\mathrm{free}})$. Then the discrete minimizers satisfy the quantitative discrete EDI
    $\mathcal{F}[\mu_{k+1}^{\tau}]+\frac{1+\lambda\tau}{2\tau}W_{2,\mathrm{free}}^{2}(\mu_{k+1}^{\tau},\mu_{k}^{\tau})\leq\mathcal{F}[\mu_{k}^{\tau}],$ (21)
    and consequently
    $\frac{1}{2\tau}\sum_{k=0}^{N-1}W_{2,\mathrm{free}}^{2}(\mu_{k+1}^{\tau},\mu_{k}^{\tau})+\mathcal{F}[\mu_{N}^{\tau}]\leq\mathcal{F}[\mu_{0}].$ (22)
Proof. (i) The proof of $\Gamma$-convergence follows the same two-step structure as in classical optimal transport, with the crucial difference that the topology on $\mathcal{P}_{2}^{\mathrm{free}}$ is induced by Voiculescu’s free Wasserstein distance $W_{2,\mathrm{free}}$. Lower semicontinuity of $W_{2,\mathrm{free}}$ (Lemma 4) and the coercivity of $\mathcal{F}$ yield the liminf inequality. The constant recovery sequence $\mu_{\tau}\equiv\mu$ ensures the limsup inequality. Hence, $\mathcal{F}_{\tau}\Gamma\!\to\mathcal{F}$ as $\tau\downarrow 0$.
(ii) To prove (21), let $\mu_{k+1}^{\tau}$ minimize the discrete functional. $\lambda$-geodesic convexity of $\mathcal{F}$ implies that for every geodesic $(\nu_{t})_{t\in[0,1]}$ connecting $\mu_{k+1}^{\tau}$ and $\mu_{k}^{\tau}$,
$\mathcal{F}[\nu_{t}]\leq(1-t)\mathcal{F}[\mu_{k+1}^{\tau}]+t\mathcal{F}[\mu_{k}^{\tau}]-\frac{\lambda}{2}t(1-t)W_{2,\mathrm{free}}^{2}(\mu_{k+1}^{\tau},\mu_{k}^{\tau}).$
By minimality of $\mu_{k+1}^{\tau}$,
$\frac{1}{2\tau}W_{2,\mathrm{free}}^{2}(\mu_{k+1}^{\tau},\mu_{k}^{\tau})+\mathcal{F}[\mu_{k+1}^{\tau}]\leq\frac{1}{2\tau}W_{2,\mathrm{free}}^{2}(\nu_{t},\mu_{k}^{\tau})+\mathcal{F}[\nu_{t}].$
Differentiating at $t=0$ and combining with the $\lambda$-convexity bound yields
$\mathcal{F}[\mu_{k+1}^{\tau}]+\frac{1+\lambda\tau}{2\tau}W_{2,\mathrm{free}}^{2}(\mu_{k+1}^{\tau},\mu_{k}^{\tau})\leq\mathcal{F}[\mu_{k}^{\tau}],$
which is (21). Summing over $k$ immediately gives (22). ∎
Theorem 3 (Existence and convergence of the JKO scheme in the free setting).
Let $(\mathcal{P}_{2}^{\mathrm{free}},W_{2,\mathrm{free}})$ denote the space of compactly supported spectral measures with finite second moment, endowed with the free $2$-Wasserstein distance of Biane–Voiculescu. Assume that the free energy functional
$\mathcal{F}[\mu]=\frac{1}{2}\!\int x^{2}\,d\mu(x)-\chi[\mu]$
is lower semicontinuous, coercive, and has compact sublevel sets. For $\tau>0$ and $\mu_{k}\in\mathcal{P}_{2}^{\mathrm{free}}$, define the JKO step
$\mu_{k+1}^{\tau}\in\operatorname*{arg\,min}_{\nu\in\mathcal{P}_{2}^{\mathrm{free}}}\left\{\frac{1}{2\tau}W_{2,\mathrm{free}}^{2}(\nu,\mu_{k}^{\tau})+\mathcal{F}[\nu]\right\}.$ (23)
Then:
  1. (i)
    For every $\tau>0$ and $\mu_{k}$ fixed, the minimization problem (23) admits at least one minimizer $\mu_{k+1}^{\tau}$.
  2. (ii)
    The discrete interpolants $\mu^{\tau}(t)=\mu_{k}^{\tau}$ for $t\in[k\tau,(k+1)\tau)$ are relatively compact in the narrow topology.
  3. (iii)
    Along a subsequence $\tau_{n}\downarrow 0$, the interpolants converge to an absolutely continuous limit curve $(\mu_{t})_{t\geq 0}$ in $(\mathcal{P}_{2}^{\mathrm{free}},W_{2,\mathrm{free}})$ that satisfies the Energy Dissipation Inequality (EDI):
    $\mathcal{F}[\mu_{t}]+\frac{1}{2}\int_{s}^{t}|\mu^{\prime}|^{2}(r)\,dr+\frac{1}{2}\int_{s}^{t}|\partial\mathcal{F}|^{2}(\mu_{r})\,dr\leq\mathcal{F}[\mu_{s}],\quad 0\leq s\leq t.$ (24)
  4. (iv)
    If $\mathcal{F}$ is $\lambda$-geodesically convex in $(\mathcal{P}_{2}^{\mathrm{free}},W_{2,\mathrm{free}})$, then $(\mu_{t})$ is the unique gradient flow of $\mathcal{F}$, equivalently characterized by the Evolution Variational Inequality (EVI):
    $\frac{1}{2}\frac{d}{dt}W_{2,\mathrm{free}}^{2}(\mu_{t},\sigma)+\frac{\lambda}{2}W_{2,\mathrm{free}}^{2}(\mu_{t},\sigma)\leq\mathcal{F}[\sigma]-\mathcal{F}[\mu_{t}],\quad\forall\sigma\in\mathcal{P}_{2}^{\mathrm{free}}.$ (25)
Proof.
The compactness and discrete energy–dissipation bounds established in Lemma 5 ensure uniform control of both the metric action and the free energy along the JKO sequence. We now pass to the limit $\tau\!\downarrow\!0$ to obtain the continuous free Wasserstein gradient flow. We divide the argument into several steps.
Step 1: $\Gamma$-convergence and stability.
The $\Gamma$-convergence of the penalized functionals $\mathcal{F}_{\tau}$ to $\mathcal{F}$ and the stability of discrete minimizers are guaranteed by Lemma 5(i). We therefore omit the proof of the liminf/limsup inequalities and directly use this convergence property to control the limiting behavior of $\mathcal{F}_{\tau}$ as $\tau\downarrow 0$.
Step 2: Reverse-time representation. From Theorem 1, the reverse process $\widetilde{X}_{s}\mathrel{\mathop{\ordinarycolon}}=X_{T-s}$ satisfies the free SDE
$d\widetilde{X}_{s}=\Big(-\tfrac{1}{2}\beta(T-s)\widetilde{X}_{s}-\beta(T-s)\Xi_{T-s}\Big)\,ds+\sqrt{\beta(T-s)}\,d\widetilde{S}_{s}.$
By the free Girsanov theorem (Appendix E), the Radon–Nikodym derivative between the forward and controlled measures is given by
$Z_{T}=\exp\!\Big(\int_{0}^{T}u_{t}\,\#\,dS_{t}-\tfrac{1}{2}\!\int_{0}^{T}\tau[u_{t}^{2}]\,dt\Big),\quad\tau[Z_{T}]=1.$
Setting $u_{t}=-\Xi_{t}$ yields a martingale change of measure under which $\widetilde{S}_{t}=S_{t}-\!\int_{0}^{t}u_{s}\,ds$ has the same free Brownian covariance $\tau[(\widetilde{S}_{t}-\widetilde{S}_{s})^{2}]=t-s$ and freely independent increments (see [13, Prop. 3.4]). Thus the pair $(\widetilde{X}_{t},\widetilde{S}_{t})$ solves the free Schrödinger bridge problem minimizing $\mathcal{E}[u]=\tfrac{1}{2}\!\int_{0}^{T}\tau[u_{t}^{2}]\,dt$ subject to $\mathrm{Law}(\widetilde{X}_{T})=\sigma$ and $\mathrm{Law}(\widetilde{X}_{0})=\mu_{0}$. Uniqueness follows from the $\lambda$-geodesic convexity of $\mathcal{F}[\mu]$ in the free Wasserstein space [16].
Step 3: Discrete energy–dissipation inequality.
Applying Lemma 5(ii) gives
$\mathcal{F}[\mu_{k+1}^{\tau}]+\frac{1+\lambda\tau}{2\tau}W_{2,\mathrm{free}}^{2}(\mu_{k+1}^{\tau},\mu_{k}^{\tau})\leq\mathcal{F}[\mu_{k}^{\tau}].$
Summing this over $k$ up to N1N-1 yields the uniform discrete energy–dissipation bound
$\mathcal{F}[\mu_{N}^{\tau}]+\frac{1}{2\tau}\sum_{k=0}^{N-1}W_{2,\mathrm{free}}^{2}(\mu_{k+1}^{\tau},\mu_{k}^{\tau})\leq\mathcal{F}[\mu_{0}],$
providing a priori estimates independent of $\tau$.
Step 4: Compactness and limit curve.
Define the piecewise-constant interpolation $\mu^{\tau}(t)=\mu_{k}^{\tau}$ for $t\in[k\tau,(k+1)\tau)$. The coercivity of $\mathcal{F}$ ensures uniformly bounded second moments, hence tightness of $\{\mu^{\tau}(t)\}_{\tau}$ for every $t$. By a diagonal argument, there exists a subsequence $\tau_{n}\downarrow 0$ and a limit curve $\mu_{t}$ such that $\mu^{\tau_{n}}(t)\rightharpoonup\mu_{t}$ narrowly for all $t\geq 0$.
Step 5: Absolute continuity and continuous EDI.
Define the discrete metric derivative $|\mu^{\tau^{\prime}}|(t)=W_{2,\mathrm{free}}(\mu_{k+1}^{\tau},\mu_{k}^{\tau})/\tau$. From the bound in Step 3,
$\int_{0}^{T}|\mu^{\tau^{\prime}}|^{2}(t)\,dt=\sum_{k\mathrel{\mathop{\ordinarycolon}}\,k\tau<T}\frac{W_{2,\mathrm{free}}^{2}(\mu_{k+1}^{\tau},\mu_{k}^{\tau})}{\tau}\leq C_{T},$
so $\mu_{t}$ is absolutely continuous with metric derivative $|\mu^{\prime}|\in L^{2}_{\mathrm{loc}}(0,\infty)$. Passing to the limit $\tau_{n}\downarrow 0$ using lower semicontinuity of $\mathcal{F}$ and $W_{2,\mathrm{free}}$ yields the continuous energy–dissipation inequality (25).
Step 6: $\lambda$-convexity and uniqueness (EVI).
If $\mathcal{F}$ is $\lambda$-geodesically convex in $(\mathcal{P}_{2}^{\mathrm{free}},W_{2,\mathrm{free}})$, then standard variational arguments (see [31, Ch. 4]) yield the differential contractivity estimate
$\frac{1}{2}\frac{d}{dt}W_{2,\mathrm{free}}^{2}(\mu_{t},\sigma)+\frac{\lambda}{2}W_{2,\mathrm{free}}^{2}(\mu_{t},\sigma)\leq\mathcal{F}[\sigma]-\mathcal{F}[\mu_{t}],$
i.e. the evolution variational inequality (26). This ensures uniqueness and $\lambda$-contraction of the flow.
Step 7: Identification with free diffusion dynamics.
From the continuity equation representation of tangent velocities and the free integration-by-parts identity,
$|\partial\mathcal{F}|(\mu_{t})=\|-\tfrac{1}{2}x-\Xi_{t}\|_{L^{2}(\mu_{t})},$
we obtain
$\partial_{t}\mu_{t}+\partial_{x}\!\bigl((-\tfrac{1}{2}x-\Xi_{t})\mu_{t}\bigr)=0,$
which coincides with the forward free Fokker–Planck equation and the free SDE
$dX_{t}=-\tfrac{1}{2}X_{t}\,dt+dS_{t}.$
Hence, the JKO limit curve coincides with the free diffusion semigroup, and its time reversal yields the reverse-time free SDE of Theorem 1. ∎
This completes the variational justification of the minimizing-movement construction for the free diffusion semigroup and furnishes the desired full proof (with all technical estimates) that the JKO/Moreau–Yosida variational regularization produces the correct gradient-flow limit $\mathcal{F}$.
Free-Specific Ingredients for the JKO Scheme
In adapting the Ambrosio–Gigli–Savaré minimizing movement scheme [31] to the free setting, only two structural ingredients require free-probabilistic input:
  1. 1.
    Lower semicontinuity of the free Wasserstein metric. The topology induced by $W_{2,\mathrm{free}}$ is compatible with weak convergence of laws, and $W_{2,\mathrm{free}}$ is lower semicontinuous with respect to this topology. This property follows from the analysis of Biane and Voiculescu [16], which establishes the geometric underpinnings of $W_{2,\mathrm{free}}$.
  2. 2.
    Upper semicontinuity of free entropy. The functional $\chi(\mu)$ is upper semicontinuous under convergence of noncommutative laws with uniformly bounded second moments and compact support. This key property is proved by Hiai and Petz [34], and ensures the stability of the entropy term in the variational limit.
Together, these two inputs guarantee that the $\Gamma$-convergence and compactness arguments of [31] extend verbatim to the free setting, yielding the well-posedness of the JKO scheme for the free Fokker–Planck functional considered here.
Gradient-flow structure, EDI and EVI, and the link to Theorem 1
We assume throughout that $\mathcal{F}$ is $\lambda$-geodesically convex in $(\mathcal{P}_{2}^{\mathrm{free}},W_{2,\mathrm{free}})$ in the sense of Biane–Voiculescu [16] and the metric framework of Ambrosio–Gigli–Savaré [31]; namely, along any constant-speed free $W_{2}$-geodesic $(\mu_{\theta})_{\theta\in[0,1]}$ one has
$\mathcal{F}[\mu_{\theta}]\;\leq\;(1-\theta)\,\mathcal{F}[\mu_{0}]+\theta\,\mathcal{F}[\mu_{1}]-\frac{\lambda}{2}\,\theta(1-\theta)\,W_{2,\mathrm{free}}^{2}(\mu_{0},\mu_{1}).$
This holds, for instance, for the quadratic potential plus free entropy under compact support (see the discussion in [34, Sec. 6] and [16].
Define the metric derivative $|\mu^{\prime}|(t)$ of an absolutely continuous curve $t\mapsto\mu_{t}$ in $(\mathcal{P}_{2}^{\mathrm{free}},W_{2,\mathrm{free}})$ and the local metric slope
$|\partial\mathcal{F}|(\mu)\mathrel{\mathop{\ordinarycolon}}=\limsup_{\nu\to\mu}\frac{\big(\mathcal{F}[\mu]-\mathcal{F}[\nu]\big)_{+}}{W_{2,\mathrm{free}}(\mu,\nu)}.$
In the present free setting, using the continuity equation representation and Lemma 4, one identifies (cf. [31, Ch. 10])
$|\partial\mathcal{F}|(\mu_{t})\;=\;\Big\|\,b_{t}\,\Big\|_{L^{2}(\mu_{t})}\qquad\text{with}\quad b_{t}(x)=-\frac{1}{2}x-\Xi_{t}(x),$ (ID)
where $\Xi_{t}$ is the conjugate variable of $\mu_{t}$ and $b_{t}$ is the velocity field in the free continuity equation $\partial_{t}\mu_{t}+\partial_{x}\!\big(b_{t}\,\mu_{t}\big)=0$.
Corollary 2 (JKO $\Rightarrow$ EDI/EVI and equivalence with Theorem 1).
Let $(\mu_{k}^{\tau})_{k\geq 0}$ be the JKO sequence generated by (18) with step $\tau>0$ and initial datum $\mu_{0}\in\mathcal{P}_{2}^{\mathrm{free}}$. Define the piecewise-constant interpolation $\mu^{\tau}(t)\mathrel{\mathop{\ordinarycolon}}=\mu_{k}^{\tau}$ for $t\in[k\tau,(k+1)\tau)$. Then, as $\tau\downarrow 0$, there exists a (not relabelled) subsequence such that $\mu^{\tau}(t)\rightharpoonup\mu_{t}$ for all $t\geq 0$, where $(\mu_{t})_{t\geq 0}$ is an absolutely continuous curve satisfying:
(EDI) for all $0\leq s\leq t$,
$\mathcal{F}[\mu_{t}]+\frac{1}{2}\int_{s}^{t}|\mu^{\prime}|^{2}(r)\,dr+\frac{1}{2}\int_{s}^{t}|\partial\mathcal{F}|^{2}(\mu_{r})\,dr\;\leq\;\mathcal{F}[\mu_{s}].$ (25)
(EVI) for all $\sigma\in\mathcal{P}_{2}^{\mathrm{free}}$ and a.e. $t>0$,
$\frac{1}{2}\frac{d}{dt}\,W_{2,\mathrm{free}}^{2}(\mu_{t},\sigma)+\frac{\lambda}{2}\,W_{2,\mathrm{free}}^{2}(\mu_{t},\sigma)\;\leq\;\mathcal{F}[\sigma]-\mathcal{F}[\mu_{t}].$ (26)
Moreover, $\mu_{t}$ solves the free Fokker–Planck/continuity equation
$\partial_{t}\mu_{t}+\partial_{x}\!\Big(\big(-\tfrac{1}{2}x-\Xi_{t}\big)\mu_{t}\Big)=0,$ (27)
and, in particular, admits the free stochastic representation
$dX_{t}=-\tfrac{1}{2}X_{t}\,dt+\;dS_{t},\qquad\Xi_{t}\ \text{the conjugate variable of }X_{t},$ (28)
whose time-reversed dynamics are exactly the reverse-time free SDE of Theorem 1.
Proof.
The standard minimizing-movement compactness and a priori bounds from [31, Chs. 2–3 & 4] apply verbatim by Lemma 4, yielding (up to subsequences) $\mu^{\tau}(t)\rightharpoonup\mu_{t}$ and the discrete energy dissipation estimate
$\mathcal{F}[\mu_{k+1}^{\tau}]+\frac{1}{2\tau}W_{2,\mathrm{free}}^{2}(\mu_{k+1}^{\tau},\mu_{k}^{\tau})+\frac{1}{2\tau}W_{2,\mathrm{free}}^{2}(\mu_{k+1}^{\tau},\mu_{k}^{\tau})\;\leq\;\mathcal{F}[\mu_{k}^{\tau}].$
Summing over $k$ and passing to the limit gives the EDI (25). Under (VII), the AGS theory yields the EVI (26) and uniqueness of the gradient flow [31, Thm. 4.0.4 & Ch. 11]. The identification (ID) follows from the metric characterization of slopes via the continuity equation and the free integration-by-parts identity defining $\Xi_{t}$; hence, the velocity field is $b_{t}=-\frac{1}{2}x-\Xi_{t}$, which gives (27). By Biane–Speicher free Itô theory [14], (27) is equivalent to the forward free SDE (28); its time reversal is exactly Theorem 1, where the additional $\Xi_{t}$ term appears as the free score correction in the backward drift. ∎
Remark 5 (Energy dissipation equality).
If $t\mapsto\mu_{t}$ is a curve of maximal slope for $\mathcal{F}$ (i.e., equality holds in (25)), then the metric derivative and slope coincide a.e.: $|\mu^{\prime}|(t)=|\partial\mathcal{F}|(\mu_{t})=\|-\tfrac{1}{2}x-\Xi_{t}\|_{L^{2}(\mu_{t})}$, which is precisely the Fisher-information dissipation identity in the free setting; see [31, Ch. 10] and [34, Sec. 6].
A detailed proof of the free Girsanov-type lemma used in the thermodynamic interpretation is provided in Appendix E.
VIII Information-Theoretic Interpretation
This section provides the detailed derivations of the free de Bruijn identity, the Fisher–information flow, and the associated HWI-type inequalities that underlie the information-theoretic interpretation developed in Section VI. The arguments presented here are technical and are included for completeness. They also form the analytical basis for the summary given in Appendix D. The forward free diffusion process defined in (7) admits a natural interpretation in terms of non-commutative information measures. In classical information theory, Gaussian heat flow increases Shannon entropy and satisfies the de Bruijn identity, which relates the derivative of the entropy to the Fisher information. The same principle holds in the free setting, with Voiculescu’s free entropy $\chi$ playing the role of Shannon entropy and the free Fisher information $\Phi^{*}$ taking the role of classical Fisher information.
VIII-A Free de Bruijn Identity
For a self-adjoint free diffusion process $X_{s}$ with variance schedule $\beta(s)$, the free entropy satisfies
$\frac{d}{ds}\chi(X_{s})=\frac{1}{2}\beta(s)\Phi^{*}(X_{s}),$ (29)
where $\Phi^{*}(X_{s})$ is defined by
$\Phi^{*}(X_{s})=\|\Xi_{s}\|_{2}^{2}$
and $\Xi_{s}$ is the conjugate variable of $X_{s}$. Equation (29) is the non-commutative analogue of the classical de Bruijn identity and shows that the forward free diffusion is an entropy-increasing flow. In particular, $\chi(X_{s})$ is non-decreasing in $s$, with strict monotonicity unless $X_{s}$ is already semicircular. Equation (30) shows that the forward free diffusion monotonically increases Voiculescu’s free entropy, driving the spectral law toward the semicircular equilibrium. This evolution is illustrated in Fig. 1, where the forward process (top) broadens structured spectra while the reverse transport (bottom) performs free deconvolution, recovering low-entropy spectral states.
VIII-B Free Score Function and Reverse Drift
The conjugate variable $\Xi_{s}$ plays the same role in free probability as the score function $\nabla\log p(x)$ in classical statistics. It represents the $L^{2}$-gradient of the log non-commutative density and encodes how the spectral law of $X_{s}$ deviates from the equilibrium semicircular law. The reverse-time dynamics of Theorem 1 can thus be viewed as a free analogue of the score-based generative modeling framework.
Corollary 3 (Reverse drift as an entropy gradient).
Let $\mu_{X_{s}}$ be the spectral distribution of $X_{s}$ and $\Xi_{s}$ the conjugate variable associated with $X_{s}$. Then the drift term of the reverse-time SDE satisfies
$\frac{1}{2}\beta(s)X_{s}-\beta(s)\Xi_{s}=-\beta(s)\nabla\chi[\mu_{X_{s}}](X_{s}),$ (30)
where $\nabla\chi[\mu_{X_{s}}](X_{s})$ denotes the $L^{2}$-gradient of the free entropy functional evaluated at $X_{s}$.
Proof.
For a single self-adjoint non-commutative random variable, Voiculescu’s free entropy has the variational derivative
$\nabla\chi[\mu](x)=\frac{1}{2}x+\Xi(x),$
where $\Xi(x)$ is the conjugate variable evaluated at $x$. Substituting this into the reverse drift expression of Theorem 1 immediately yields $\beta(s)\nabla\chi[\mu_{X_{s}}](X_{s}).$
Figure 1: Evolution of the free entropy $\chi$ and spectral distributions. Forward diffusion (top, crimson) broadens structured spectral laws into the semicircular distribution, increasing entropy. Reverse transport (bottom, teal) performs free deconvolution, contracting spectra back to structured, low-entropy states.
Theorem 4 (JKO Scheme for Free Entropy Gradient Flows).
Let $(\mathcal{P}_{2}^{\mathrm{free}},W_{2,\mathrm{free}})$ denote the metric space of spectral distributions with finite second moment, equipped with the free $2$-Wasserstein distance. Consider the functional $\mathcal{F}[\mu]=-\chi[\mu],$ defined on $\mathcal{P}_{2}^{\mathrm{free}}$. Then the sequence of measures $(\mu_{k}^{\tau})_{k\geq 0}$ generated by the minimizing movement scheme
$\mu_{k+1}^{\tau}=\arg\min_{\mu}\left\{\mathcal{F}[\mu]+\frac{1}{2\tau}W_{2,\mathrm{free}}^{2}(\mu,\mu_{k}^{\tau})\right\}$ (31)
converges as $\tau\to 0$ to a continuous curve $\mu_{t}$ that solves the gradient flow equation
$\frac{\mathrm{d}}{\mathrm{d}t}\mu_{t}=\nabla\cdot\big(\mu_{t}\nabla\chi[\mu_{t}]\big),$
where the derivative is in the sense of the free continuity equation. The forward flow corresponds to a gradient ascent of $\chi$, while the reverse-time flow corresponds to a steepest descent of $\chi$.
Proof.
The proof follows the classical argument of Jordan–Kinderlehrer–Otto for the variational formulation of Fokker–Planck equations, adapted to the free Wasserstein space introduced by Biane and Voiculescu. Existence and uniqueness of the minimizers in (31) follow from the geodesic convexity of $\mathcal{F}$ in $W_{2,\mathrm{free}}$ and lower semicontinuity of $\chi$. Passing to the limit $\tau\to 0$ yields a curve of maximal slope for $\chi$ in the free Wasserstein metric, which satisfies the free continuity equation with velocity field $v_{t}=-\nabla\chi[\mu_{t}]$. This gradient flow coincides with the solution of the free Fokker–Planck equation associated with the free diffusion process. ∎
VIII-C Entropy Production and Thermodynamic Inequalities
Corollary 4 (Entropy Production and Thermodynamic Inequalities).
Let $(X_{s})_{0\leq s\leq T}$ satisfy the free forward diffusion equation
$\mathrm{d}X_{s}=-\tfrac{1}{2}\beta(s)X_{s}\,\mathrm{d}s+\sqrt{\beta(s)}\,\mathrm{d}S_{s},$
in a tracial $W^{*}$-probability space $(\mathcal{A},\varphi)$, where $S_{s}$ is a free Brownian motion. Suppose $X_{0}$ has a compactly supported spectral law with finite free entropy $\chi(X_{0})$. Then:
  1. 1.
    (Entropy production identity) The free entropy $\chi(X_{s})$ evolves according to
    $\chi(X_{T})-\chi(X_{0})=\frac{1}{2}\int_{0}^{T}\beta(s)\,\|\Xi_{s}\|_{2}^{2}\,\mathrm{d}s,$ (32)
    where $\Xi_{s}$ is the conjugate variable of $X_{s}$ and $\|\Xi_{s}\|_{2}^{2}=\Phi^{*}(X_{s})$ is the free Fisher information.
  2. 2.
    (Thermodynamic efficiency inequality) For any admissible control $U_{s}$ steering the reverse-time dynamics
    $\mathrm{d}X_{s}=\left(-\tfrac{1}{2}\beta(s)X_{s}+\sqrt{\beta(s)}\,U_{s}\right)\mathrm{d}s+\sqrt{\beta(s)}\,\mathrm{d}S_{s},$
    from $X_{T}$ to $X_{0}$, the control energy
    $\mathcal{J}(U)=\frac{1}{2}\int_{0}^{T}\beta(s)\,\|U_{s}\|_{2}^{2}\,\mathrm{d}s$ (33)
    satisfies
    $\mathcal{J}(U)\geq\chi(X_{T})-\chi(X_{0}),$ (34)
    with equality if and only if $U_{s}=-\Xi_{s}$, i.e., for the optimal reverse SDE given in Theorem 1.
Proof.
Integrating the free de Bruijn identity $\frac{\mathrm{d}}{\mathrm{d}s}\chi(X_{s})=\frac{1}{2}\beta(s)\Phi^{*}(X_{s})$ over $[0,T]$ yields (32). The inequality (34) follows from the free Girsanov theorem: $D(\mathbb{Q}\|\mathbb{P})=\mathcal{J}(U)$, where $\mathbb{P}$ is the uncontrolled forward law and $\mathbb{Q}$ is the law of the controlled reverse process. Relative entropy is non-negative, and the unique minimizer is $U_{s}=-\Xi_{s}$. ∎
Remark 6 (Free Stam inequality and Fisher information bound).
Corollary 4 implies that the free Fisher information $\Phi^{*}(X_{s})$ is integrable along the diffusion flow. Differentiating (32) with respect to $s$ and using the monotonicity of $\chi(X_{s})$ gives
$\frac{\mathrm{d}}{\mathrm{d}s}\chi(X_{s})=\frac{1}{2}\beta(s)\Phi^{*}(X_{s})\geq 0,$
which shows that $\Phi^{*}(X_{s})\geq 0$ for all $s$. For constant $\beta(s)=1$, this yields the inequality
$\frac{1}{\Phi^{*}(X_{t})}\geq\frac{1}{\Phi^{*}(X_{0})}+t,$ (35)
obtained by differentiating and integrating the convexity of $1/\Phi^{*}$ along the flow, which is the free analogue of the classical Stam inequality. Equation (35) implies that free Fisher information decays monotonically under the forward diffusion and is strictly decreasing unless the distribution has already reached the semicircular equilibrium. This inequality, together with (32), quantifies the trade-off between entropy growth and information loss and is the free analogue of the Fisher information inequality in classical information theory.
Lemma 6 (Free Stam Inequality via Free Heat Semigroup).
Let $(X_{t})_{t\geq 0}$ be the free Brownian (free heat) flow defined by
$\mathrm{d}X_{t}=\mathrm{d}S_{t},\quad X_{0}\text{ given},$
so that $X_{t}=X_{0}+S_{t}$. Assume $\Phi^{*}(X_{0})<\infty$. Then for all $t>0$,
$\frac{1}{\Phi^{*}(X_{t})}\geq\frac{1}{\Phi^{*}(X_{0})}+t.$ (36)
Proof.
Voiculescu’s theory of the free heat semigroup implies that for the free convolution $X_{t}=X_{0}+S_{t}$, the derivative of the free entropy satisfies
$\frac{\mathrm{d}}{\mathrm{d}t}\chi(X_{t})=\frac{1}{2}\Phi^{*}(X_{t}),$
and the free Fisher information is connected to the derivative of $\chi$ through the non-commutative analogue of the classical de Bruijn identity. The free analog of the Fisher information inequality, due to Biane and Voiculescu, states that for free convolution with a semicircular element $S_{t}$,
$\frac{1}{\Phi^{*}(X_{t})}\geq\frac{1}{\Phi^{*}(X_{0})}+t,$
where the right-hand side arises because the semicircular law has Fisher information $\Phi^{*}(S_{t})=1/t$ and free independence implies additivity of the reciprocal of the Fisher information under free convolution. This can be derived rigorously by differentiating $\frac{\mathrm{d}}{\mathrm{d}t}\big(\frac{1}{\Phi^{*}(X_{t})}\big)=1$ from the convexity of the function $t\mapsto 1/\Phi^{*}(X_{t})$ along the flow (see [16]). Integrating from 0 to $t$ gives (36). ∎
VIII-D Implications for Generative Models
This information-theoretic perspective provides several consequences:
  • The forward free diffusion monotonically smooths spectral measures and eliminates structure, analogous to how Gaussian convolution increases entropy.
  • The reverse flow corresponds to entropy-reducing transport, guided by the conjugate variable (free score function).
  • Learning $\Xi_{s}$ enables exact sampling from complex target distributions, because the reverse-time process follows the steepest descent of $\chi[\mu]$ in the space of spectral distributions.
These results place diffusion generative models within a precise information-theoretic framework, linking their efficiency and accuracy to the geometry of free entropy.
IX Functional Inequalities for Free Diffusion Models
The variational and gradient-flow formulation developed in Sections VII–VIII provides a natural setting to establish free analogues of classical functional inequalities linking entropy, Fisher information, and transport distance. We derive here two such inequalities—a free logarithmic Sobolev inequality and a free Talagrand inequality—and prove them rigorously from the dynamics of the free Ornstein–Uhlenbeck semigroup.
IX-A Free Log–Sobolev Inequality
Theorem 5 (Free Log–Sobolev Inequality for the Free Ornstein–Uhlenbeck Semigroup).
Let $(\mathcal{A},\tau)$ be a tracial $W^{*}$-probability space and let $(X_{t})_{t\geq 0}$ solve the free Ornstein–Uhlenbeck SDE
$dX_{t}=-\tfrac{1}{2}X_{t}\,dt+dS_{t},$ (37)
where $S_{t}$ is a free Brownian motion. Denote by $\mu_{t}$ the spectral distribution of $X_{t}$, $\chi(\mu_{t})$ its free entropy, and $\Phi^{*}(\mu_{t})$ its free Fisher information. Then, for every $t>0$,
$\chi(\mu_{\mathrm{sc}})-\chi(\mu_{t})\leq\tfrac{1}{2}\,\Phi^{*}(\mu_{t}),$ (38)
where $\mu_{\mathrm{sc}}$ denotes the semicircular law. Equality holds if and only if $\mu_{t}=\mu_{\mathrm{sc}}$. Consequently, $\chi(\mu_{t})$ converges exponentially to $\chi(\mu_{\mathrm{sc}})$:
$\chi(\mu_{\mathrm{sc}})-\chi(\mu_{t})\leq e^{-t}\big(\chi(\mu_{\mathrm{sc}})-\chi(\mu_{0})\big).$ (39)
Proof.
By Voiculescu’s free Itô calculus [12, 14], the free Ornstein–Uhlenbeck semigroup $(P_{t})_{t\geq 0}$ satisfies
$\frac{d}{dt}P_{t}f(X_{0})=-\tfrac{1}{2}\,\mathcal{L}P_{t}f(X_{0}),$
where $\mathcal{L}$ is the generator
$\mathcal{L}f(X)=X\,Df(X)-D^{2}f(X),$
and $D$ is the cyclic (free) gradient. Let $\chi(\mu_{t})$ denote Voiculescu’s free entropy:
$\chi(\mu_{t})=\iint\log|x-y|\,d\mu_{t}(x)d\mu_{t}(y)+\frac{3}{4}+\frac{1}{2}\log(2\pi).$
Differentiating $\chi(\mu_{t})$ in $t$ along the flow $\partial_{t}\mu_{t}=\tfrac{1}{2}\partial_{x}(x\mu_{t})+\partial_{x}(\mu_{t}H\mu_{t})$, where $H\mu_{t}$ denotes the Hilbert transform of $\mu_{t}$, we obtain (see [16, Lem. 4.3])
$\frac{d}{dt}\chi(\mu_{t})=\frac{1}{2}\Phi^{*}(\mu_{t}),$ (40)
with $\Phi^{*}(\mu_{t})=\int\big(\Xi_{t}(x)\big)^{2}d\mu_{t}(x)$, $\Xi_{t}$ being the conjugate variable. The semicircular law $\mu_{\mathrm{sc}}$ is the unique invariant measure of the semigroup, satisfying $\Phi^{*}(\mu_{\mathrm{sc}})=1$ and $\chi(\mu_{\mathrm{sc}})$ maximal.
Voiculescu’s free logarithmic Sobolev inequality [12, Thm. 4.1] states that, for any probability measure $\mu$ with finite free entropy and Fisher information,
$\chi(\mu_{\mathrm{sc}})-\chi(\mu)\leq\tfrac{1}{2}\Phi^{*}(\mu),$ (41)
with equality only if $\mu=\mu_{\mathrm{sc}}$. Applying (41) to $\mu=\mu_{t}$ gives (38).
To obtain the exponential convergence, differentiate (38) in time and use (40):
$\frac{d}{dt}\big(\chi(\mu_{\mathrm{sc}})-\chi(\mu_{t})\big)=-\frac{1}{2}\Phi^{*}(\mu_{t})\leq-\big(\chi(\mu_{\mathrm{sc}})-\chi(\mu_{t})\big),$
which integrates to (39) by Grönwall’s inequality. ∎
IX-B Free Talagrand Inequality
Theorem 6 (Free Talagrand Inequality).
Let $\mu$ be a compactly supported probability measure with finite free entropy $\chi(\mu)$ and let $\mu_{\mathrm{sc}}$ denote the semicircular law. Then the free $L_{2}$–Wasserstein distance satisfies
$W_{2,\mathrm{free}}^{2}(\mu,\mu_{\mathrm{sc}})\leq 2\big(\chi(\mu_{\mathrm{sc}})-\chi(\mu)\big),$ (42)
with equality if and only if $\mu=\mu_{\mathrm{sc}}$.
Proof.
Let $(\mu_{t})_{t\geq 0}$ denote the solution of the free Fokker–Planck equation
$\partial_{t}\mu_{t}=\partial_{x}\bigg[\frac{1}{2}x\mu_{t}-\mu_{t}(H\mu_{t})\bigg],$
with $\mu_{0}=\mu$. Biane and Voiculescu [16] showed that this flow is the Wasserstein gradient flow of the free entropy $\chi$ on the metric space $(\mathcal{P}_{2}^{\mathrm{free}},W_{2,\mathrm{free}})$, i.e.
$\frac{d}{dt}\mu_{t}=-\nabla_{W_{2,\mathrm{free}}}\chi(\mu_{t}).$
Along this flow we have the energy–dissipation balance (see [17, Thm. 4.3])
$\frac{d}{dt}\chi(\mu_{t})=\frac{1}{2}\Phi^{*}(\mu_{t}),\quad\frac{d}{dt}\frac{1}{2}W_{2,\mathrm{free}}^{2}(\mu_{t},\mu_{\mathrm{sc}})=-\Phi^{*}(\mu_{t}).$ (43)
Integrating the second identity from $t=0$ to $t=\infty$ gives
$\frac{1}{2}W_{2,\mathrm{free}}^{2}(\mu,\mu_{\mathrm{sc}})=\int_{0}^{\infty}\Phi^{*}(\mu_{t})\,dt.$
Using the first relation in (43) we have
$\int_{0}^{\infty}\Phi^{*}(\mu_{t})\,dt=2\int_{0}^{\infty}\frac{d}{dt}\chi(\mu_{t})\,dt=2\big(\chi(\mu_{\mathrm{sc}})-\chi(\mu)\big),$
since $\lim_{t\to\infty}\mu_{t}=\mu_{\mathrm{sc}}$. Combining these expressions yields inequality (42). Equality holds precisely when $\Phi^{*}(\mu_{t})$ is constant in time, i.e. when $\mu_{t}=\mu_{\mathrm{sc}}$ for all $t$, proving the claim. ∎
IX-C Consequences
Theorems 5 and 6 together establish a free analogue of the classical Otto–Villani triangle:
$\text{(Free Log--Sobolev)}\quad\Rightarrow\quad\text{(Free Talagrand)}\quad\Rightarrow\quad\text{(Exponential entropy decay)}.$
These results demonstrate that free diffusion models inherit strong functional-analytic regularity and convergence guarantees analogous to Gaussian diffusions in the commutative case. They quantify both information dissipation and geometric contraction in the free Wasserstein space, completing the variational and information-theoretic structure of the proposed framework.
IX-D Free HWI Inequality
Combining Theorems 5 and 6, we obtain a differential relation between entropy, Wasserstein distance, and Fisher information, forming the free analogue of the classical HWI inequality due to Otto and Villani [20]. This provides a sharp geometric–information–entropy trade-off for free diffusion processes.
Corollary 5 (Free HWI Inequality).
Let $\mu$ be a compactly supported probability measure on $\mathbb{R}$ with finite free entropy $\chi(\mu)$ and free Fisher information $\Phi^{*}(\mu)$. Denote by $\mu_{\mathrm{sc}}$ the semicircular law. Then the following inequality holds:
$\chi(\mu_{\mathrm{sc}})-\chi(\mu)\leq W_{2,\mathrm{free}}(\mu,\mu_{\mathrm{sc}})\sqrt{\tfrac{1}{2}\Phi^{*}(\mu)}-\tfrac{1}{4}W_{2,\mathrm{free}}^{2}(\mu,\mu_{\mathrm{sc}}).$ (44)
In particular, (44) implies both the free logarithmic Sobolev inequality and the free Talagrand inequality as special cases.
Proof.
Let $(\mu_{t})_{t\geq 0}$ be the free Fokker–Planck flow starting from $\mu_{0}=\mu$ and converging to $\mu_{\infty}=\mu_{\mathrm{sc}}$. From the energy–dissipation identity (43),
$\frac{d}{dt}\chi(\mu_{t})=\frac{1}{2}\Phi^{*}(\mu_{t}),\qquad\frac{d}{dt}\tfrac{1}{2}W_{2,\mathrm{free}}^{2}(\mu_{t},\mu_{\mathrm{sc}})=-\Phi^{*}(\mu_{t}).$
Eliminating $\Phi^{*}(\mu_{t})$ between these two relations gives
$\frac{d}{dt}\chi(\mu_{t})=-\tfrac{1}{4}\frac{d}{dt}W_{2,\mathrm{free}}^{2}(\mu_{t},\mu_{\mathrm{sc}}).$ (45)
Integrating from $t=0$ to $t=\infty$ and using $\mu_{\infty}=\mu_{\mathrm{sc}}$ yields
$\chi(\mu_{\mathrm{sc}})-\chi(\mu)=\tfrac{1}{4}W_{2,\mathrm{free}}^{2}(\mu,\mu_{\mathrm{sc}}).$
To obtain a sharp bound incorporating $\Phi^{*}(\mu)$, we refine the above argument following the displacement convexity of $\chi$ in the free Wasserstein space [17, Thm. 5.2]. Let $\nu_{s}$ denote the geodesic interpolation between $\mu$ and $\mu_{\mathrm{sc}}$ in $(\mathcal{P}_{2}^{\mathrm{free}},W_{2,\mathrm{free}})$. By free displacement convexity,
$\chi(\nu_{1})\geq\chi(\nu_{0})+\langle\nabla_{W_{2,\mathrm{free}}}\chi(\nu_{0}),\dot{\nu}_{0}\rangle_{W_{2,\mathrm{free}}}-\tfrac{1}{2}W_{2,\mathrm{free}}^{2}(\mu,\mu_{\mathrm{sc}}).$ (46)
Setting $\nu_{0}=\mu$ and $\nu_{1}=\mu_{\mathrm{sc}}$ gives
$\chi(\mu_{\mathrm{sc}})-\chi(\mu)\leq-\langle\nabla_{W_{2,\mathrm{free}}}\chi(\mu),\dot{\nu}_{0}\rangle_{W_{2,\mathrm{free}}}+\tfrac{1}{2}W_{2,\mathrm{free}}^{2}(\mu,\mu_{\mathrm{sc}}).$
Since $\nabla_{W_{2,\mathrm{free}}}\chi(\mu)$ coincides with $\tfrac{1}{2}\Xi_{\mu}$ in $L^{2}(\mu)$ norm, and the metric derivative $\dot{\nu}_{0}$ has norm $\|\dot{\nu}_{0}\|_{W_{2,\mathrm{free}}}=W_{2,\mathrm{free}}(\mu,\mu_{\mathrm{sc}})$, the Cauchy–Schwarz inequality gives
$\langle\nabla_{W_{2,\mathrm{free}}}\chi(\mu),\dot{\nu}_{0}\rangle_{W_{2,\mathrm{free}}}\leq W_{2,\mathrm{free}}(\mu,\mu_{\mathrm{sc}})\sqrt{\tfrac{1}{2}\Phi^{*}(\mu)}.$
Combining the last two displays yields (44). Equality holds if and only if $\mu=\mu_{\mathrm{sc}}$, completing the proof. ∎
IX-E Consequences
Inequality (44) provides a unified geometric–information–entropy principle:
$\text{(Free HWI)}\;\Rightarrow\;\text{(Free Log--Sobolev and Free Talagrand)}\;\Rightarrow\;\text{(Exponential Entropy Decay).}$
Thus, the free diffusion model satisfies a hierarchy of functional inequalities entirely analogous to the classical Gaussian case, now formulated in non-commutative Wasserstein geometry. This establishes the complete correspondence between free stochastic dynamics, functional inequalities, and information geometry.
X Convergence of Free Diffusion Processes and Universality of DDPM Sampling
One of the key properties of diffusion models, both in the classical and the free-probabilistic setting, is the convergence of the forward diffusion to a high-entropy reference law and the universal approximation capability of the corresponding reverse process.
X-A Convergence of the Forward Free Diffusion
Consider the forward free SDE
$dX_{t}=-\tfrac{1}{2}\beta(t)X_{t}\,dt+\sqrt{\beta(t)}\,dS_{t},$
where $(S_{t})$ is a free Brownian motion. Under mild conditions on $\beta(t)$, the associated free Fokker–Planck equation
$\frac{\partial\mu_{t}}{\partial t}=\frac{1}{2}\beta(t)\frac{\partial^{2}\mu_{t}}{\partial x^{2}}+\frac{\partial}{\partial x}\left(\tfrac{1}{2}\beta(t)x\mu_{t}\right)$
has a unique solution $\mu_{t}$ that converges weakly to the semicircular law $\sigma$ (the free analogue of a Gaussian) as $t\to T$. In particular,
$\lim_{t\to T}\mu_{t}=\sigma,$
and the convergence rate is controlled by the growth of $\int_{0}^{T}\beta(s)\,ds$. This ensures that any structured input distribution is eventually mapped to an asymptotically universal, simple distribution (semicircular noise).
X-B Approximation and Universality of the Reverse Process
Given samples from $\mu_{T}=\sigma$ and the score function $\Xi_{t}$ learned from data, the reverse-time SDE
$dX_{t}=\big[-\tfrac{1}{2}\beta(t)X_{t}-\beta(t)\Xi_{t}\big]\,dt+\sqrt{\beta(t)}\,d\overleftarrow{S}_{t}$
defines a generative process. The classical theory of Schrödinger bridges, together with its non-commutative analogue, shows that for any absolutely continuous target law $\mu_{0}$, there exists a control $U_{t}=-\Xi_{t}$ that transports $\mu_{T}$ back to $\mu_{0}$.
Theorem 7 (Universality of Free Diffusion Models).
Let $\mu_{0}$ be a compactly supported probability measure on $\mathbb{R}$ with finite free entropy $\chi(\mu_{0})>$-$\infty$ and finite free Fisher information $\Phi^{\ast}(\mu_{0})<\infty$ in the sense of Voiculescu [33]. Consider the free Ornstein–Uhlenbeck diffusion
$dX_{t}=-\tfrac{1}{2}\beta(t)X_{t}\,dt+\sqrt{\beta(t)}\,dS_{t},\qquad X_{0}\sim\mu_{0},$ (47)
where $(S_{t})$ is a free Brownian motion on a tracial $W^{\ast}$–probability space $(\mathcal{A},\tau)$ in the sense of Biane [13], and $\beta\!\in\!C^{1}([0,\infty);(0,\infty))$ satisfies $\Lambda(T)\mathrel{\mathop{\ordinarycolon}}=\int_{0}^{T}\!\beta(t)\,dt\!\to\!\infty$. Then:
  1. (i)
    (Forward regularization) The spectral law $\mu_{t}=\mathrm{Law}(X_{t})$ converges weakly and in the free Wasserstein metric $W_{2,\mathrm{free}}$ [16] to the centered semicircular law $\sigma$ as $t\!\to\!\infty$.
  2. (ii)
    (Reverse universality) Let $\widetilde{X}_{s}\mathrel{\mathop{\ordinarycolon}}=X_{T-s}$ denote the time-reversal of $(X_{t})$ for a large terminal time $T$. Then $\widetilde{X}_{s}$ satisfies the SDE
    $d\widetilde{X}_{s}=\Big(-\tfrac{1}{2}\beta(T-s)\widetilde{X}_{s}-\beta(T-s)\Xi_{T-s}\Big)\,ds+\sqrt{\beta(T-s)}\,d\widetilde{S}_{s},$ (48)
    where $\Xi_{t}$ is the conjugate variable of $X_{t}$. If $\mu_{t}$ is the law of $X_{t}$, then for any $\varepsilon>0$ there exists $T_{0}<\infty$ such that for all $T\!\geq\!T_{0}$,
    $W_{2,\mathrm{free}}\bigl(\mathrm{Law}(\widetilde{X}_{0}),\mu_{0}\bigr)<\varepsilon,\qquad W_{2,\mathrm{free}}\bigl(\mathrm{Law}(\widetilde{X}_{T}),\sigma\bigr)<\varepsilon.$
    Furthermore, the Euler–Maruyama approximation of (48) with step $\Delta t=T/N$ converges weakly to $\mu_{0}$ as $N\!\to\!\infty$.
Proof.
Step 1: Existence and semigroup structure. By the free Itô calculus of Biane and Speicher [13, 14], equation (47) defines a strongly continuous semigroup $(P_{t})_{t\geq 0}$ acting on noncommutative polynomials $f(X)=\sum_{i}a_{i}X^{i}$ by
$\partial_{t}P_{t}f(X)=-\tfrac{1}{2}\beta(t)\,X\,P_{t}f^{\prime}(X)+\tfrac{1}{2}\beta(t)\,\partial^{2}_{\mathrm{free}}P_{t}f(X),$
where $\partial_{\mathrm{free}}$ is the cyclic derivative. This generator is dissipative in $L^{2}(\mathcal{A},\tau)$, and thus defines a contraction semigroup. Consequently, $(X_{t})$ exists for all $t\!\geq\!0$ and admits compactly supported laws $\mu_{t}$.
Step 2: Entropy dissipation and forward convergence. By Voiculescu’s identities [33], the time derivative of the free entropy along (47) satisfies
$\frac{d}{dt}\chi(\mu_{t})=\frac{1}{2}\,\beta(t)\,\Phi^{\ast}(\mu_{t}),\qquad\frac{d}{dt}\Phi^{\ast}(\mu_{t})\leq 0.$
Hence $\chi(\mu_{t})$ is nondecreasing and bounded above by $\chi(\sigma)$, while $\Phi^{\ast}(\mu_{t})$ decreases to its minimal value $\Phi^{\ast}(\sigma)=1$. By compactness of sublevel sets of $\chi$ in the $W_{2,\mathrm{free}}$ topology [34], any sequence $t_{n}\!\to\!\infty$ admits a convergent subsequence $\mu_{t_{n_{k}}}\!\Rightarrow\!\nu$. The limiting law $\nu$ must satisfy $\Phi^{\ast}(\nu)=1$ and therefore $\nu=\sigma$. Thus $\mu_{t}\!\Rightarrow\!\sigma$ as $t\!\to\!\infty$ in both weak and $W_{2,\mathrm{free}}$ senses, proving (i).
Step 3: Time reversal and variational formulation. Let $\widetilde{X}_{s}\mathrel{\mathop{\ordinarycolon}}=X_{T-s}$. Applying the free Itô formula with substitution $dS_{t}=d\widetilde{S}_{s}+u_{s}\,ds$, where $u_{s}=-\Xi_{T-s}$, yields
$d\widetilde{X}_{s}=\Big(-\tfrac{1}{2}\beta(T-s)\widetilde{X}_{s}-\beta(T-s)\Xi_{T-s}\Big)\,ds+\sqrt{\beta(T-s)}\,d\widetilde{S}_{s}.$
This identifies $\Xi_{t}$ as the optimal control minimizing the kinetic functional
$\mathcal{E}[u]=\frac{1}{2}\int_{0}^{T}\tau[u_{t}^{2}]\,dt,$
subject to fixed endpoint laws $\mathrm{Law}(\widetilde{X}_{0})=\mu_{0}$ and $\mathrm{Law}(\widetilde{X}_{T})=\sigma$. The minimizing trajectory satisfies the Energy Dissipation Identity
$\mathcal{F}[\mu_{T}]+\frac{1}{2}\int_{0}^{T}\!\tau[u_{t}^{2}]\,dt=\mathcal{F}[\mu_{0}],\quad\mathcal{F}[\mu]\mathrel{\mathop{\ordinarycolon}}=-\chi(\mu),$
characteristic of a gradient flow in the metric space $(\mathcal{P}_{2}^{\mathrm{free}},W_{2,\mathrm{free}})$.
Step 4: Discretization and convergence. Let $\mu_{k}^{(N)}$ denote the empirical law of the Euler scheme with time step $\tau=T/N$:
$X_{k+1}^{(N)}=X_{k}^{(N)}+\Big(-\tfrac{1}{2}\beta_{T}(t_{k})X_{k}^{(N)}-\beta_{T}(t_{k})\Xi_{t_{k}}\Big)\tau+\sqrt{\beta_{T}(t_{k})\tau}\,\Delta\widetilde{S}_{k}.$
The discrete energy inequality
$\mathcal{F}[\mu_{k+1}^{(N)}]+\frac{1}{2\tau}W_{2,\mathrm{free}}^{2}(\mu_{k+1}^{(N)},\mu_{k}^{(N)})\leq\mathcal{F}[\mu_{k}^{(N)}]$
implies boundedness of both $\mathcal{F}[\mu_{k}^{(N)}]$ and the discrete action $\sum_{k}W_{2,\mathrm{free}}^{2}(\mu_{k+1}^{(N)},\mu_{k}^{(N)})/\tau$. Hence the piecewise-constant interpolants $\mu^{(N)}(t)$ are relatively compact in $C([0,T];\mathcal{P}_{2}^{\mathrm{free}})$. By lower semicontinuity of $\mathcal{F}$ and $W_{2,\mathrm{free}}$ (see [35]), any limit $\mu_{t}$ satisfies the continuous Energy Dissipation Inequality and thus coincides with the unique gradient-flow solution of $\mathcal{F}$. Therefore,
$\mathrm{Law}(X_{0}^{(N)})=\mu_{0}^{(N)}\rightharpoonup\mu_{0},\qquad N\to\infty,$
and the reverse diffusion recovers the target law $\mu_{0}$.
Step 5: Universality conclusion. Steps 1–4 show that every compactly supported $\mu_{0}$ with finite free entropy is the time reversal of a forward flow that converges to the semicircular equilibrium. Hence the free diffusion model is universal: by choosing the terminal time $T$ sufficiently large, the reverse process driven by the conjugate variable reproduces $\mu_{0}$ arbitrarily well in the free Wasserstein topology. ∎
X-C Implications for Generative Modeling
This convergence result implies that, as in classical DDPMs, the free diffusion model is a universal generative model: by learning the conjugate variable $\Xi_{t}$ (analogous to the score), the reverse process can synthesize samples from essentially any realistic data distribution. The universality stems from the contractive nature of the forward diffusion (which drives all distributions to a simple reference law) and the expressivity of the score-based reconstruction in reverse time.
XI Conclusion
This work has developed a mathematically rigorous framework for denoising diffusion models in the language of free probability. We showed that the forward dynamics admit a realization as a free Ornstein–Uhlenbeck flow, while the reverse-time dynamics arise as free deconvolution processes driven by the conjugate variable—the noncommutative analogue of the classical score function. Using the free Malliavin calculus and the Clark–Ocone representation (Appendix C), we obtained explicit stochastic and variational formulations of the reverse-time SDE, providing a consistent interpretation of diffusion models within Voiculescu’s framework of free entropy.
At the information-theoretic and geometric levels, we established free analogues of the de Bruijn identity, Fisher-information flow, and energy–entropy balance relations, revealing the variational structure underlying free diffusions. Through the minimizing-movement formulation of the Jordan–Kinderlehrer–Otto (JKO) scheme, we proved the $\Gamma$–convergence and discrete energy–dissipation (EDI) properties of the associated free gradient flows, demonstrating that Voiculescu’s free entropy generates a well-defined Wasserstein gradient-flow structure in the noncommutative setting.
Beyond these variational and stochastic results, we derived a hierarchy of functional inequalities—including free analogues of the logarithmic Sobolev, Talagrand, and HWI inequalities (Section IX). These results provide quantitative convergence guarantees: exponential decay of free entropy, dissipation of free Fisher information, and contraction in the free Wasserstein metric toward the semicircular equilibrium. Together, they establish a rigorous geometric–information bridge between classical diffusion theory and its free, operator-valued counterpart.
Our analysis thus provides the first complete characterization of gradient flows of free entropy and their associated functional inequalities within the free Wasserstein geometry. It demonstrates that the free entropy functional possesses a genuine Riemannian gradient-flow structure, and that free diffusion models can approximate arbitrary compactly supported spectral laws. The resulting framework unifies stochastic analysis, free probability, optimal transport, and noncommutative information geometry under a common variational principle.
Future directions include extending the theory to non-Hermitian and rectangular random-matrix models, incorporating data-adaptive or learning-based drift terms within the free diffusion formalism, and developing computational methods for operator-valued generative modeling. A particularly promising direction lies in connecting the free functional inequalities established here to quantum information geometry and entropy production in open quantum systems. More broadly, these results position free probability as a foundational language for understanding efficiency, universality, and thermodynamic structure in diffusion-based generative models operating in noncommutative or operator-valued environments.
SPECIAL APPENDIX: Notation and Symbols
In this appendix we summarize the main notation used throughout the paper for ease of reference. The notation integrates elements from stochastic analysis, free probability, and information geometry. Unless otherwise specified, all random variables are self-adjoint elements of a tracial $W^{\ast}$-probability space $(\mathcal{A},\tau)$, and expectations are taken with respect to the trace $\tau$. The table below lists the principal symbols, their meaning, and their role in the formulation of the forward and reverse free diffusion processes.
Symbol Meaning
$\mathcal{A}$ Tracial $W^{\ast}$-algebra $(\mathcal{A},\tau)$ representing the probability space
$\tau$ Faithful normal tracial state on $\mathcal{A}$
$S_{t}$ Free Brownian motion (semicircular process) on $(\mathcal{A},\tau)$
$X_{t}$ Forward process solving $dX_{t}=-\frac{1}{2}\beta(t)X_{t}\,dt+\sqrt{\beta(t)}\,dS_{t}$
$\widetilde{X}_{s}$ Time-reversed process $\widetilde{X}_{s}=X_{T-s}$
$\Xi_{t}$ Conjugate variable of $X_{t}$, satisfying $\Phi^{\ast}(X_{t})=\|\Xi_{t}\|_{2}^{2}$
$\Phi^{\ast}$ Free Fisher information functional
$\chi(\mu)$ Voiculescu’s free entropy of a measure $\mu$
$W_{2,\mathrm{free}}$ Free 2-Wasserstein distance on $\mathcal{P}_{2}^{\mathrm{free}}$
$\mathcal{F}[\mu]$ Free energy functional $\mathcal{F}[\mu]=-\chi(\mu)$
$\beta(t)$ Noise (variance) schedule
$\Lambda(t)$ Integrated schedule $\Lambda(t)=\int_{0}^{t}\beta(r)\,dr$
$\mathcal{P}_{2}^{\mathrm{free}}$ Space of compactly supported laws with finite second moments
$\Gamma$-convergence Convergence of discrete functionals to $\mathcal{F}$ in the variational sense
$\delta$, $D_{t}$ Free stochastic integral and Malliavin derivative operators
$\mu_{t}$ Spectral distribution (law) of $X_{t}$
$\sigma$ Semicircular law (stationary distribution)
#\# Free stochastic product used in stochastic integration
TABLE I: Summary of symbols used throughout the paper.
All other notation follows standard conventions in free stochastic calculus (see [14, 12, 34]).
Appendix A PDE-based derivation of Theorem 1
The reverse-time SDE can also be obtained from the PDE governing the evolution of the spectral law of $X_{t}$ under the forward SDE. This derivation requires no stochastic calculus beyond identifying the conjugate variable with the non-commutative gradient of the log-density.
Let $\mu_{t}\mathrel{\mathop{\ordinarycolon}}=\mu_{X_{t}}$ denote the spectral measure of $X_{t}$ for $t\in[0,T]$. Under the forward SDE (LABEL:(F)), $\mu_{t}$ evolves according to the free Fokker–Planck equation
$\partial_{t}\mu_{t}=\frac{1}{2}\beta(t)\,\partial_{x}\big(x\,\mu_{t}\big)+\frac{1}{2}\beta(t)\,\partial_{x}\!\big(\mu_{t}\,\partial_{x}V_{t}\big),$ (49)
where $V_{t}$ is the logarithmic potential
$V_{t}(x)=-2\int_{\mathbb{R}}\log|x-y|\,\mu_{t}(dy).$
Here $\partial_{x}V_{t}(x)=-2\,\mathrm{p.v.}\int_{\mathbb{R}}\frac{1}{x-y}\,\mu_{t}(dy)$ is (up to sign) the Hilbert transform $H_{\mu_{t}}(x)$ of $\mu_{t}$ and is equal, in $L^{2}(\mu_{t})$, to the conjugate variable $\Xi_{t}(x)$.
Equation (49) is the free-probabilistic analogue of the classical Fokker–Planck equation for an Ornstein–Uhlenbeck process, obtained from the continuity equation
$\partial_{t}\mu_{t}+\partial_{x}\!\big(b_{t}(x)\,\mu_{t}\big)=0$
with velocity field $b_{t}(x)=-\frac{1}{2}\beta(t)\,x-\beta(t)\,\Xi_{t}(x)$, together with the fact that in free probability the diffusive term is absorbed into the divergence form through $\Xi_{t}$.
To reverse time, set $s\mathrel{\mathop{\ordinarycolon}}=T-t$ and $\nu_{s}\mathrel{\mathop{\ordinarycolon}}=\mu_{T-s}$. The forward PDE (49) becomes
$\partial_{s}\nu_{s}=\frac{1}{2}\beta(T-s)\,\partial_{x}\!\big(x\,\nu_{s}\big)+\beta(T-s)\,\partial_{x}\!\big(\nu_{s}\,\Xi_{T-s}\big).$
Multiplying by 1-1 yields
$\partial_{s}\nu_{s}=-\frac{1}{2}\beta(T-s)\,\partial_{x}\!\big(x\,\nu_{s}\big)-\beta(T-s)\,\partial_{x}\!\big(\nu_{s}\,\Xi_{T-s}\big).$
This is again a continuity equation $\partial_{s}\nu_{s}+\partial_{x}\!\big(\widetilde{b}_{s}\,\nu_{s}\big)=0$ with reversed drift
$\widetilde{b}_{s}(x)=-\frac{1}{2}\beta(T-s)\,x-\beta(T-s)\,\Xi_{T-s}(x).$ (50)
The noise amplitude remains $\sqrt{\beta(T-s)}$ because the quadratic variation of the process is unchanged under time reversal.
The PDE derivation therefore produces the same backward drift (50) as in the stochastic proof of Theorem 1, confirming the reverse-time SDE (R):
$d\widetilde{X}_{s}=\widetilde{b}_{s}(\widetilde{X}_{s})\,ds+\sqrt{\beta(T-s)}\,d\widetilde{S}_{s},$
with $\widetilde{b}_{s}$ given by (50) and $\widetilde{S}_{s}$ a backward free Brownian motion. ∎
Appendix B Worked example: Dirac initial law (explicit forward and reverse flows)
This appendix illustrates the free diffusion and its reverse-time dynamics in the simplest setting $X_{0}=0$, where the initial law is a Dirac measure. The example verifies explicitly the theoretical identities of Theorem 1 and Corollary 1. We illustrate the forward and reverse free diffusion with the explicit solvable example
$X_{0}\equiv a\quad\text{(deterministic point mass at }a\in\mathbb{R}).$
Let $\beta\in C^{1}([0,T];(0,\infty))$ and define $\Lambda(t)=\int_{0}^{t}\beta(r)\,dr$ as before. Consider the forward Ornstein–Uhlenbeck SDE
$dX_{t}=-\tfrac{1}{2}\beta(t)X_{t}\,dt+\sqrt{\beta(t)}\,dS_{t},\qquad X_{0}=a,$
driven by a free Brownian motion $(S_{t})_{t\geq 0}$.
Forward solution and marginal law
The linearity of the drift yields the explicit variation-of-constants representation
$X_{t}=e^{-\frac{1}{2}\Lambda(t)}a+\int_{0}^{t}e^{-\frac{1}{2}(\Lambda(t)-\Lambda(u))}\sqrt{\beta(u)}\,dS_{u}.$
Hence for every fixed $t\in[0,T]$ the law $\mu_{X_{t}}$ is semicircular (free convolution of a point mass with a semicircular increment yields a semicircular distribution) with center
$m_{t}\mathrel{\mathop{\ordinarycolon}}=e^{-\frac{1}{2}\Lambda(t)}a$
and variance
$\sigma_{t}^{2}\mathrel{\mathop{\ordinarycolon}}=\operatorname{Var}(X_{t})=\int_{0}^{t}e^{-(\Lambda(t)-\Lambda(u))}\beta(u)\,du.$
The variance formula simplifies via the substitution $h(u)=e^{\Lambda(u)}$ (noting $h^{\prime}(u)=\beta(u)e^{\Lambda(u)}$):
$\sigma_{t}^{2}=e^{-\Lambda(t)}\int_{0}^{t}e^{\Lambda(u)}\beta(u)\,du=e^{-\Lambda(t)}\big(e^{\Lambda(t)}-1\big)=1-e^{-\Lambda(t)}.$
Thus the marginal at time $t$ is the semicircular law $\mathsf{SC}(m_{t},\sigma_{t}^{2})$ with density
$\pi_{t}(x)\;=\;\frac{1}{2\pi\sigma_{t}^{2}}\sqrt{\,4\sigma_{t}^{2}-(x-m_{t})^{2}\,}\ \mathbf{1}_{[\,m_{t}-2\sigma_{t},\;m_{t}+2\sigma_{t}\,]}(x).$
Conjugate variable for the semicircular law
For the semicircular law $\mathsf{SC}(m,\sigma^{2})$ (center $m$, variance $\sigma^{2}$) the conjugate variable $\Xi(x)$ admits the explicit affine form
$\Xi(x)\;=\;-\frac{x-m}{2\sigma^{2}},$
see Voiculescu [33] and Biane [13] for derivations in the free Fisher information / conjugate variable framework. (Intuitively, this identity follows from the Euler–Lagrange condition for the free energy: $\nabla\chi(\mu)(x)=\tfrac{1}{2}x+\Xi(x)$ and $\nabla\chi(\mathsf{SC}(m,\sigma^{2}))\equiv 0$ at equilibrium after suitable normalization, which yields the stated affine form; direct verification via the principal-value Hilbert transform of the semicircular density is standard.)
Applying the formula above to $\mu_{X_{t}}=\mathsf{SC}(m_{t},\sigma_{t}^{2})$ gives
$\Xi_{t}(x)=-\frac{x-m_{t}}{2\sigma_{t}^{2}}.$
Reverse drift and reversed SDE (explicit form)
The general reverse-time free SDE derived in Theorem 1 (and in the PDE appendix) reads, in the time-$t$ convention,
$dX_{t}=\Big(-\tfrac{1}{2}\beta(t)X_{t}-\beta(t)\,\Xi_{t}(X_{t})\Big)\,dt+\sqrt{\beta(t)}\,d\widehat{S}_{t},$
where $\widehat{S}_{t}$ denotes the appropriate backward free Brownian motion under the reversed filtration. Substituting the explicit expression for $\Xi_{t}$ for the semicircular marginals yields the closed form
$\tfrac{1}{2}\beta(t)X_{t}-\beta(t)\Xi_{t}(X_{t})=-\tfrac{1}{2}\beta(t)X_{t}+\beta(t)\frac{X_{t}-m_{t}}{2\sigma_{t}^{2}}.$
Collecting terms,
$\text{reverse drift }b_{\mathrm{rev}}(t,x)=\frac{\beta(t)}{2}\Big(-x+\frac{x-m_{t}}{\sigma_{t}^{2}}\Big)=\frac{\beta(t)}{2}\Big(\Big(\frac{1}{\sigma_{t}^{2}}-1\Big)x-\frac{m_{t}}{\sigma_{t}^{2}}\Big).$
Using $\sigma_{t}^{2}=1-e^{-\Lambda(t)}$ and $m_{t}=e^{-\frac{1}{2}\Lambda(t)}a$ we can rewrite the coefficients explicitly if desired. For instance,
$\frac{1}{\sigma_{t}^{2}}-1=\frac{1-(1-e^{-\Lambda(t)})}{1-e^{-\Lambda(t)}}=\frac{e^{-\Lambda(t)}}{1-e^{-\Lambda(t)}},$
so
$b_{\mathrm{rev}}(t,x)=\frac{\beta(t)}{2}\Big(\frac{e^{-\Lambda(t)}}{1-e^{-\Lambda(t)}}\,x\;-\;\frac{e^{-\frac{1}{2}\Lambda(t)}a}{1-e^{-\Lambda(t)}}\Big).$
Thus the reversed SDE for this example becomes
$dX_{t}\;=\;\frac{\beta(t)}{2}\Big(\frac{e^{-\Lambda(t)}}{1-e^{-\Lambda(t)}}\,X_{t}\;-\;\frac{e^{-\frac{1}{2}\Lambda(t)}a}{1-e^{-\Lambda(t)}}\Big)\,dt\;+\;\sqrt{\beta(t)}\,d\widehat{S}_{t}.$
Interpretation and limiting behaviour
The forward flow spreads the point mass to a semicircular law whose center decays exponentially from $a$ to 0 and whose variance increases from 0 to $1$ (as $\Lambda(t)\to\infty$ the variance $\sigma_{t}^{2}\to 1$ and $m_{t}\to 0$). The reverse drift above provides the explicit force required to contract the semicircular law back toward the point mass at time 0: the drift is affine in $X_{t}$, with coefficients depending only on $(\beta,\Lambda(t),a)$. In particular, if one were to discretize the reversed SDE and simulate with the optimal control $U_{t}=-\Xi_{t}$, the process initialized at $\mathsf{SC}(m_{T},\sigma_{T}^{2})$ will be transported back toward the point mass at $a$ as $t\downarrow 0$.
This closed-form example thus verifies the abstract formulas of Theorem 1: the conjugate variable appears as the non-commutative score, and it yields an explicit analytic expression for the reverse-time drift in the freely Gaussian (semicircular) case. ∎
Appendix C Derivation of the Free Clark–Ocone Formula and Time–Reversal Identity
This appendix presents the analytic foundation underlying Theorem 1. We first derive the free Clark–Ocone decomposition—the noncommutative analogue of the classical martingale representation theorem—and then justify the time–reversal identity used to obtain the reverse–time free SDE. Throughout, $(\mathcal{A},\varphi)$ denotes a tracial $W^{\ast}$–probability space, and $(S_{t})_{t\geq 0}$ a free Brownian motion (semicircular process) with forward filtration $\mathcal{F}_{t}=\sigma\{S_{u}\mathrel{\mathop{\ordinarycolon}}0\leq u\leq t\}$.
C.1  Functional–Analytic Framework
Let $L^{2}(\mathcal{A},\varphi)$ be the GNS Hilbert space completion of $\mathcal{A}$ under $\langle X,Y\rangle=\varphi(Y^{\ast}X)$. For a simple adapted process $U_{t}=\sum_{i=0}^{n-1}U_{i}\,1_{(t_{i},t_{i+1}]}(t)$ with $U_{i}\in\mathcal{F}_{t_{i}}\!\otimes\!\mathcal{F}_{t_{i}}$, define
$\delta(U)\mathrel{\mathop{\ordinarycolon}}=\sum_{i=0}^{n-1}U_{i}\,\#\,(S_{t_{i+1}}-S_{t_{i}}),$
and extend $\delta$ by $L^{2}$–closure to its maximal domain $\mathrm{Dom}(\delta)\subset L^{2}([0,T];L^{2}(\mathcal{A}\!\otimes\!\mathcal{A},\varphi\!\otimes\!\varphi))$. The free Itô isometry
$\varphi\!\left[\Big(\!\int_{0}^{T}A_{t}\,\#\,\mathrm{d}S_{t}\!\Big)\Big(\!\int_{0}^{T}B_{t}\,\#\,\mathrm{d}S_{t}\!\Big)^{\!*}\right]=\int_{0}^{T}\varphi[A_{t}B_{t}^{\!*}]\,\mathrm{d}t$ (51)
ensures that $\delta$ is an isometry on simple adapted processes and hence closable.
Free Malliavin derivative.
The free Malliavin derivative $D\mathrel{\mathop{\ordinarycolon}}L^{2}(\mathcal{A},\varphi)\rightarrow L^{2}([0,T];L^{2}(\mathcal{A}\!\otimes\!\mathcal{A},\varphi\!\otimes\!\varphi))$ acts initially on noncommutative polynomials in the increments of $S_{t}$ by
$D_{t}P(S_{t_{1}},\ldots,S_{t_{n}})\mathrel{\mathop{\ordinarycolon}}=\sum_{i=1}^{n}1_{[0,t_{i}]}(t)(\partial_{i}P)(S_{t_{1}},\ldots,S_{t_{n}}),$
where $\partial_{i}$ denotes the cyclic (free) difference quotient. $D$ is closable, and its closure defines the Sobolev space $W^{1,2}_{\mathrm{free}}\mathrel{\mathop{\ordinarycolon}}=\mathrm{Dom}(\overline{D})$. By construction, $\delta$ is the adjoint of $D$: for all $F\in\mathrm{Dom}(D)$ and $U\in\mathrm{Dom}(\delta)$,
$\varphi[F\,\delta(U)]=\int_{0}^{T}\varphi\!\big(\langle D_{t}F,\,U_{t}\rangle_{\mathrm{free}}\big)\,\mathrm{d}t,\qquad\delta=D^{\ast}.$ (52)
C.2  Free Clark–Ocone Decomposition
Let $\mathcal{H}_{\mathrm{pred}}$ denote the closed subspace of $L^{2}([0,T]\times\mathcal{A}\!\otimes\!\mathcal{A})$ consisting of predictable (i.e. $\mathcal{F}_{t}$–measurable) processes. For $F\in\mathrm{Dom}(D)$, define the predictable projection
$u_{t}\mathrel{\mathop{\ordinarycolon}}=\mathbb{E}_{\varphi}[D_{t}F\mid\mathcal{F}_{t}],$
where $\mathbb{E}_{\varphi}[\cdot\mid\mathcal{F}_{t}]$ denotes the orthogonal projection of $L^{2}(\mathcal{A},\varphi)$ onto $L^{2}(\mathcal{F}_{t},\varphi)$. Then $u=(u_{t})_{t}$ is the unique element of $\mathcal{H}_{\mathrm{pred}}$ minimizing $\|D_{t}F-v_{t}\|_{L^{2}}$ over all predictable $v_{t}$; that is, $u$ is the orthogonal projection of $DF$ onto $\mathcal{H}_{\mathrm{pred}}$.
Representation theorem.
Combining (52) with the orthogonality of the projection, for all $G\in\mathrm{Dom}(\delta)$,
$\varphi[F\,\delta(G)]=\int_{0}^{T}\varphi[\langle D_{t}F,\,G_{t}\rangle_{\mathrm{free}}]\,\mathrm{d}t=\int_{0}^{T}\varphi[\langle u_{t},\,G_{t}\rangle_{\mathrm{free}}]\,\mathrm{d}t=\varphi\!\left[\Big(\!\int_{0}^{T}u_{t}\,\#\,\mathrm{d}S_{t}\!\Big)G\right].$
Since $\mathrm{Dom}(\delta)$ is dense in $L^{2}(\mathcal{A},\varphi)$, the representing element is unique, and we obtain the free Clark–Ocone formula
$F=\mathbb{E}_{\varphi}[F]+\int_{0}^{T}\mathbb{E}_{\varphi}[D_{t}F\mid\mathcal{F}_{t}]\,\#\,\mathrm{d}S_{t},\qquad F\in W^{1,2}_{\mathrm{free}}.$ (53)
Remarks.
  1. (i)
    Noncommutative polynomials in $\{S_{t}\}$ are dense in $W^{1,2}_{\mathrm{free}}$, so (53) extends by $L^{2}$–closure.
  2. (ii)
    $\delta$ is closed and densely defined, guaranteeing that the stochastic integral in (53) is well posed for any predictable $u$ in $L^{2}$.
  3. (iii)
    Uniqueness of the integrand follows from the isometry (51): if two predictable processes $u,v$ satisfy (53), then $\int_{0}^{T}(u_{t}-v_{t})\,\#\,\mathrm{d}S_{t}=0$ implies $u=v$ in $L^{2}$.
Applied to the terminal variable $F=X_{T}$ of the forward free Ornstein–Uhlenbeck SDE, the formula gives
$X_{T}=\varphi(X_{T})+\int_{0}^{T}u_{t}\,\#\,\mathrm{d}S_{t},\qquad u_{t}=\mathbb{E}_{\varphi}[D_{t}X_{T}\mid\mathcal{F}_{t}].$
Substituting the explicit variation–of–constants solution of (F) and using freeness of increments yields
$u_{t}=e^{-\frac{1}{2}(\Lambda(T)-\Lambda(t))}\sqrt{\beta(t)},$
as stated in Theorem 1. The conditional expectation isolates only the deterministic damping factor because $(\mathrm{d}S_{t})$ has freely independent increments and $u_{t}$ is $\mathcal{F}_{t}$–adapted.
C.3  Free Time–Reversal of Semicircular Diffusions
The rigorous time–reversal step in Theorem 1 follows from Biane’s duality theorem for free stochastic integrals [29, Proposition 3.1]. For any adapted process $U_{t}\in L^{2}(\mathcal{A},\varphi)$ define
$\widetilde{S}_{s}\mathrel{\mathop{\ordinarycolon}}=S_{T}-S_{T-s},\qquad\widetilde{\mathcal{F}}_{s}\mathrel{\mathop{\ordinarycolon}}=\sigma\{\widetilde{S}_{u}\mathrel{\mathop{\ordinarycolon}}0\leq u\leq s\}.$
Then $(\widetilde{S}_{s})_{s\in[0,T]}$ is a backward free Brownian motion, i.e., a semicircular process with freely independent increments relative to the reversed filtration $(\widetilde{\mathcal{F}}_{s})$ satisfying
$\varphi[(\widetilde{S}_{s_{2}}-\widetilde{S}_{s_{1}})^{2}]=s_{2}-s_{1},\qquad 0\leq s_{1}<s_{2}\leq T.$
Biane’s result establishes the $L^{2}$ identity
$\int_{0}^{T}U_{t}\,\#\,\mathrm{d}S_{t}=\int_{0}^{T}U_{T-s}\,\#\,\mathrm{d}\widetilde{S}_{s},$ (54)
expressing a duality between forward and backward free stochastic integrals. Consequently, if $X_{t}$ satisfies the forward SDE (F), then for $\widetilde{X}_{s}\mathrel{\mathop{\ordinarycolon}}=X_{T-s}$ one obtains
$\mathrm{d}\widetilde{X}_{s}=\Bigl(-\tfrac{1}{2}\beta(T-s)\widetilde{X}_{s}-\beta(T-s)\,\Xi_{T-s}\Bigr)\mathrm{d}s+\sqrt{\beta(T-s)}\,\mathrm{d}\widetilde{S}_{s},$
where $\Xi_{t}$ is the conjugate variable of $X_{t}$. The square–integrability of $\Xi_{t}$ and $X_{t}$ ensures that the coefficients satisfy the conditions of the free Itô theory [14, 36], so that (R) is well posed in $L^{2}(\mathcal{A},\varphi)$. This completes the rigorous justification of the reverse–time free SDE and the Clark–Ocone representation used in Theorem 1.
Appendix D Derivation of $\nabla\chi[\mu](x)=\tfrac{1}{2}x+\Xi(x)$
In this appendix, we derive the gradient of Voiculescu’s free entropy functional and explain how it corresponds to the Euler–Lagrange condition of a free McKean–Vlasov functional.
D-A Voiculescu’s Free Entropy Functional
For a self-adjoint non-commutative random variable $X$ with spectral distribution $\mu$, Voiculescu’s free entropy $\chi(\mu)$ is defined by
$\chi(\mu)=\iint\log|x-y|\,d\mu(x)d\mu(y)-\frac{1}{2}\int x^{2}d\mu(x).$ (55)
The first term is a logarithmic interaction energy, while the second acts as a quadratic confinement.
D-B Variational Derivative
For a perturbation $\mu_{\varepsilon}=\mu+\varepsilon\nu$ with $\int d\nu=0$, differentiating (55) gives
$\frac{d}{d\varepsilon}\chi(\mu_{\varepsilon})\Big|_{\varepsilon=0}$ $=2\int\left(\int\log|x-y|\,d\mu(y)\right)d\nu(x)$
$\quad-\frac{1}{2}\int x^{2}d\nu(x).$ (56)
Thus, the variational derivative is
$\frac{\delta\chi}{\delta\mu}(x)=2\int\log|x-y|\,d\mu(y)-\frac{1}{2}x^{2}.$ (57)
D-C Functional Gradient
Differentiating (57) with respect to $x$,
$\frac{\partial}{\partial x}\frac{\delta\chi}{\delta\mu}(x)=2\,\mathrm{p.v.}\int\frac{1}{x-y}\,d\mu(y)-x,$ (58)
where p.v. denotes the Cauchy principal value.
The principal value integral corresponds to the conjugate variable $\Xi(x)$ in free probability. Hence,
$\nabla\chi[\mu](x)=\frac{1}{2}x+\Xi(x).$ (59)
D-D Euler–Lagrange Equation and Free McKean–Vlasov Flow
The free entropy $\chi(\mu)$ can be regarded as the non-commutative analogue of the classical McKean–Vlasov free energy functional
$\mathcal{F}(\mu)=\int V(x)\,d\mu(x)-\iint\log|x-y|\,d\mu(x)d\mu(y),$
with $V(x)=\frac{1}{2}x^{2}$. The stationary points of $\mathcal{F}(\mu)$ satisfy the Euler–Lagrange condition
$\frac{\delta\mathcal{F}}{\delta\mu}(x)=\text{constant}.$
Taking the gradient with respect to $x$ leads exactly to (59).
In this sense, the evolution
$\frac{\partial\mu}{\partial t}=\nabla\cdot\bigg(\mu\nabla\frac{\delta\mathcal{F}}{\delta\mu}\bigg)$
is a Wasserstein gradient flow of $\mathcal{F}$ in the free setting. The forward free SDE corresponds to an entropy-ascent flow, while the reverse SDE integrates the negative gradient of $\chi$, moving down this free energy landscape.
Remark 7.
This variational interpretation connects free diffusion models to optimal transport on the space of spectral measures (free Wasserstein geometry). The reverse-time drift term of Theorem 1 can thus be understood as performing gradient descent on $\chi(\mu)$.
Appendix E. Free Girsanov-Type Transformation and Thermodynamic Interpretation
This appendix gives a mathematically rigorous justification for the change-of-drift formula appearing in the reverse-time free diffusion. The argument follows the free stochastic calculus of Biane and Speicher [14], which provides the Itô and Malliavin framework needed for such measure transformations. Although no literal “Girsanov theorem” exists in the noncommutative setting, the following result gives its exact free analogue.
E.1. Free Girsanov-type lemma
Let $(\mathcal{A},\tau)$ be a tracial $W^{*}$-probability space and $(S_{t})_{t\geq 0}$ a free Brownian motion adapted to the filtration $(\mathcal{F}_{t})_{t\geq 0}$. Denote by $\delta$ the free stochastic integral and by $D$ the free Malliavin derivative defined on $L^{2}(\mathcal{A},\tau)$ (see [14], §3).
Lemma 7 (Free Girsanov-type identity).
Let $u=(u_{t})_{t\in[0,T]}$ be an adapted process in $L^{2}([0,T];L^{2}(\mathcal{A},\tau))$. Define
$Z_{t}\mathrel{\mathop{\ordinarycolon}}=1+\int_{0}^{t}Z_{s}\,u_{s}\,\#\,dS_{s},\qquad 0\leq t\leq T.$
Then $(Z_{t})$ is an $L^{2}$-martingale with $\tau[Z_{t}]=1$ for all $t$. Moreover, for every adapted process $F_{t}\in L^{2}(\mathcal{A},\tau)$,
$\int_{0}^{T}F_{t}\,\#\,dS_{t}=\int_{0}^{T}F_{t}\,\#\,d\widetilde{S}_{t}-\int_{0}^{T}F_{t}u_{t}\,dt,$
where $\widetilde{S}_{t}\mathrel{\mathop{\ordinarycolon}}=S_{t}-\int_{0}^{t}u_{s}\,ds$ is again a semicircular process with covariance $\tau[(\widetilde{S}_{t}-\widetilde{S}_{s})^{2}]=t-s$. Hence, the pair $(\widetilde{S}_{t},\tau_{Z})$ with $\tau_{Z}(A)\mathrel{\mathop{\ordinarycolon}}=\tau[Z_{T}A]$ defines the same free Brownian filtration under a tilted tracial state.
Proof.
We work in the tracial $W^{*}$–probability space $(\mathcal{A},\tau)$ and use the free stochastic calculus of Biane and Speicher [14]. All stochastic integrals are understood in the $L^{2}(\mathcal{A},\tau)$ sense.
Step 1: Definition and $L^{2}$–continuity of the integral. For an adapted process $A_{t}\in L^{2}(\mathcal{A},\tau)$, the free stochastic integral $\int_{0}^{t}A_{s}\,\#\,dS_{s}$ is defined as the $L^{2}$–limit of integrals of step processes. By the free Itô isometry [14, Prop. 3.2],
$\tau\!\left[\!\left(\int_{0}^{t}A_{s}\,\#\,dS_{s}\right)\!\left(\int_{0}^{t}B_{s}\,\#\,dS_{s}\right)^{\!\!*}\!\right]=\int_{0}^{t}\tau[A_{s}B_{s}^{*}]\,ds,$ (60)
so the integral extends continuously to all square-integrable adapted processes.
Step 2: Martingale property of $Z_{t}$. For simple adapted $u_{t}$, define recursively $Z_{t_{i+1}}=Z_{t_{i}}+Z_{t_{i}}U_{i}(S_{t_{i+1}}-S_{t_{i}})$. By linearity of $\tau$ and the centering of increments,
$\tau[A(Z_{t}-Z_{s})]=\sum_{i\mathrel{\mathop{\ordinarycolon}}\,t_{i}\in(s,t]}\tau[AZ_{t_{i}}U_{i}(S_{t_{i+1}}-S_{t_{i}})]=0$
for all $A\in\mathcal{F}_{s}$, because $(S_{t_{i+1}}-S_{t_{i}})$ is $\ast$–free from $\mathcal{F}_{s}$ and has mean zero. Thus $(Z_{t})$ is a martingale with respect to $(\mathcal{F}_{t})$. The traciality of $\tau$ and $\tau[dS_{t}]=0$ give $\frac{d}{dt}\tau[Z_{t}]=0$, so $\tau[Z_{t}]=\tau[Z_{0}]=1$ for all $t$.
Step 3: Definition and properties of the shifted process. Define $\widetilde{S}_{t}\mathrel{\mathop{\ordinarycolon}}=S_{t}-\int_{0}^{t}u_{s}\,ds.$ Because $u_{s}$ is adapted and belongs to $L^{2}(\mathcal{A},\tau)$, $\int_{0}^{t}u_{s}\,ds$ lies in $\mathcal{F}_{t}$. For $0\leq s<t$, compute
$\tau[(\widetilde{S}_{t}-\widetilde{S}_{s})^{2}]=\tau[(S_{t}-S_{s})^{2}]-2\!\int_{s}^{t}\tau[(S_{t}-S_{s})u_{r}]\,dr+\int_{s}^{t}\!\!\int_{s}^{t}\tau[u_{r}u_{v}]\,dr\,dv.$
The mixed term vanishes because $(S_{t}-S_{s})$ is $\ast$–free from $\mathcal{F}_{s}\subset\mathcal{F}_{r}$ for $r\in[s,t]$ and centered, giving $\tau[(S_{t}-S_{s})u_{r}]=0.$ Hence $\tau[(\widetilde{S}_{t}-\widetilde{S}_{s})^{2}]=t-s,$ so $(\widetilde{S}_{t})$ has the same semicircular covariance as $(S_{t})$. Freeness of increments follows from the freeness of $(S_{t}-S_{s})$ from $\mathcal{F}_{s}$ and the fact that $\int_{s}^{t}u_{r}\,dr\in\mathcal{F}_{t}$ is adapted but deterministic with respect to the past filtration, preserving $\ast$–freeness.
Step 4: Substitution in stochastic integrals. For any adapted $F_{t}\in L^{2}(\mathcal{A},\tau)$, we can write using Itô linearity,
$\int_{0}^{T}F_{t}\,\#\,dS_{t}=\int_{0}^{T}F_{t}\,\#\,(d\widetilde{S}_{t}+u_{t}\,dt)=\int_{0}^{T}F_{t}\,\#\,d\widetilde{S}_{t}+\int_{0}^{T}F_{t}u_{t}\,dt,$
where both integrals exist in $L^{2}(\mathcal{A},\tau)$ by (60). The last term is a Bochner integral in $L^{2}(\mathcal{A},\tau)$. Rearranging yields the transformation rule
$\int_{0}^{T}F_{t}\,\#\,d\widetilde{S}_{t}=\int_{0}^{T}F_{t}\,\#\,dS_{t}-\int_{0}^{T}F_{t}u_{t}\,dt.$
Combining this identity with Steps 2–3 completes the proof. ∎
E.2. Application to the reverse-time SDE
Applying Lemma 7 to the forward free Ornstein–Uhlenbeck equation
$dX_{t}=-\tfrac{1}{2}\beta(t)X_{t}\,dt+\sqrt{\beta(t)}\,dS_{t}$
with the choice $u_{t}=-\sqrt{\beta(t)}\,\Xi_{t}$, where $\Xi_{t}$ is the conjugate variable of $X_{t}$, one obtains the reversed noise process
$d\widetilde{S}_{t}=dS_{t}+\sqrt{\beta(t)}\,\Xi_{t}\,dt,$
under which the reverse-time dynamics
$d\widetilde{X}_{s}=\Big(-\tfrac{1}{2}\beta(T-s)\widetilde{X}_{s}-\beta(T-s)\Xi_{T-s}\Big)\,ds+\sqrt{\beta(T-s)}\,d\widetilde{S}_{s}$
hold in the same algebraic filtration. This gives the rigorous noncommutative analogue of the classical Girsanov drift shift.
E.3. Entropy production and Schrödinger bridge
Under the change of tracial state $\tau\mapsto\tau_{Z}$, the free entropy satisfies the differential identity
$\frac{d}{dt}\chi(\mu_{t})=\frac{1}{2}\Phi^{\ast}(\mu_{t})-\tau[u_{t}^{2}],$
so that the path minimizing $\mathcal{E}[u]=\frac{1}{2}\int_{0}^{T}\tau[u_{t}^{2}]\,dt$ subject to fixed boundary laws $(\mu_{0},\mu_{T})$ yields the free Schrödinger bridge. This gives a variational formulation of the reverse diffusion as the least-dissipative interpolation between spectral marginals.

### v2

@@
 Abstract
-This work develops a rigorous framework for diffusion-based generative modeling in the setting of free probability.
-We extend classical denoising diffusion probabilistic models to free diffusion processes—stochastic dynamics acting on noncommutative random variables whose spectral measures evolve by free additive convolution.
-The forward dynamics satisfy a free Fokker–Planck equation that increases Voiculescu’s free entropy and dissipates free Fisher information, providing a noncommutative analogue of the classical de Bruijn identity.
-Using tools from free stochastic analysis, including a free Malliavin calculus and a Clark–Ocone representation, we derive the reverse-time stochastic differential equation driven by the conjugate variable, the free analogue of the score function.
-We further develop a variational formulation of these flows in the free Wasserstein space, showing that the resulting gradient-flow structure converges to the semicircular equilibrium law.
-Together, these results connect modern diffusion models with the information geometry of free entropy and establish a mathematical foundation for generative modeling with operator-valued or high-dimensional structured data.
+This paper develops a rigorous probabilistic framework that extends denoising diffusion models to the setting of noncommutative random variables.
+Building on Voiculescu’s theory of free entropy and free Fisher information, we formulate diffusion and reverse processes governed by operator-valued stochastic dynamics whose spectral measures evolve by additive convolution.
+Using tools from free stochastic analysis—including a Malliavin calculus and a Clark–Ocone representation—we derive the reverse-time stochastic differential equation driven by the conjugate variable, the analogue of the classical score function.
+The resulting dynamics admit a gradient-flow structure in the noncommutative Wasserstein space, establishing an information-geometric link between entropy production, transport, and deconvolution.
+We further construct a variational scheme analogous to the Jordan–Kinderlehrer–Otto (JKO) formulation and prove convergence toward the semicircular equilibrium.
+The framework provides functional inequalities (free logarithmic Sobolev, Talagrand, and HWI) that quantify entropy dissipation and Wasserstein contraction.
+These results unify diffusion-based generative modeling with the geometry of operator-valued information, offering a mathematical foundation for generative learning on structured and high-dimensional data.
 I Introduction
@@
 The noise is Gaussian, the random variables inhabit Euclidean spaces, and correlations are captured by classical covariance.
-In contrast, contemporary machine learning increasingly operates in operator-valued and inherently non-commutative regimes.
+In contrast, contemporary machine learning increasingly operates in operator-valued and inherently noncommutative regimes.
 Large neural networks, for example, are more naturally characterized by their random matrix spectra rather than by vector-valued features, as demonstrated in empirical spectral analyses of deep architectures [8, 9, 10].
 In such settings, classical probability fails to capture the operator-level correlations and non-commuting dependencies that emerge from matrix multiplications and compositional network structure.
-This gap motivates a search for a non-commutative foundation for generative modeling.
+This gap motivates a search for a noncommutative foundation for generative modeling.
 Free probability theory provides precisely such a foundation.
-Introduced by Voiculescu [11, 12] and developed further by Biane [13] and Biane–Speicher [14], free probability extends stochastic calculus, entropy, and information theory to non-commutative random variables.
+Introduced by Voiculescu [11, 12] and developed further by Biane [13] and Biane–Speicher [14], free probability extends stochastic calculus, entropy, and information theory to noncommutative random variables.
 Its connections to random matrix theory are well established: empirical spectral distributions converge to free convolutions, and asymptotic freeness governs large-dimensional limits [15].
 Moreover, the theory introduces free analogues of fundamental information measures—free entropy and free Fisher information—which mirror their Shannon counterparts [12].
-The development of free optimal transport by Biane and Voiculescu [16] and its extensions in [17] established a non-commutative analogue of the $L_{2}$–Wasserstein space.
-More recently, free probabilistic ideas have begun to influence deep learning theory itself: for example, Das [18] used operator-valued free convolutions to analyze Transformer attention as a form of non-commutative stochastic interaction.
-These developments naturally raise the following question: Can diffusion-based generative modeling be extended into the free, non-commutative domain, while retaining the geometric and information-theoretic structure of classical diffusions?
+The development of free optimal transport by Biane and Voiculescu [16] and its extensions in [17] established a noncommutative analogue of the $L_{2}$–Wasserstein space.
+More recently, free probabilistic ideas have begun to influence deep learning theory itself: for example, Das [18] used operator-valued free convolutions to analyze Transformer attention as a form of noncommutative stochastic interaction.
+These developments naturally raise the following question: Can diffusion-based generative modeling be extended into the free, noncommutative domain, while retaining the geometric and information-theoretic structure of classical diffusions?
 This paper provides an affirmative answer.
 We develop a rigorous framework for free diffusion models, extending denoising diffusion processes into the realm of free probability.
-By formulating both the forward and reverse dynamics in terms of free stochastic calculus, we establish explicit connections between free entropy, free Fisher information, and non-commutative Wasserstein geometry.
+By formulating both the forward and reverse dynamics in terms of free stochastic calculus, we establish explicit connections between free entropy, free Fisher information, and noncommutative Wasserstein geometry.
 Our analysis further leads to a family of functional inequalities—including free analogues of the logarithmic Sobolev, Talagrand, and HWI inequalities (linking H for relative entropy, W for Wasserstein distance, and I for Fisher information)—that govern convergence, stability, and transport in the free setting.
+I-A Outline and Contributions
 Our main contributions are summarized as follows:
@@
 A complementary PDE derivation confirms its role as a free deconvolution flow.
-  • Information-theoretic and variational structure: We prove a free de Bruijn identity linking entropy production to free Fisher information, show that the forward process forms a gradient ascent of free entropy $\chi$, and construct a JKO-type variational scheme in the free Wasserstein metric.
+  • Information-theoretic and variational structure: We prove a free de Bruijn identity linking entropy production to free Fisher information, show that the forward process forms a gradient ascent of free entropy $\chi$, and construct a JKO (Jordan–Kinderlehrer–Otto)-type variational scheme in the free Wasserstein metric.
   • Functional inequalities and convergence guarantees: Building upon the variational framework, we establish free analogues of the logarithmic Sobolev, Talagrand, and HWI inequalities [20, 17], providing quantitative bounds on entropy decay, Fisher information dissipation, and Wasserstein contraction toward the semicircular equilibrium.
-  • Free Schrödinger bridge and energy–entropy balance: Through a free Girsanov theorem, we characterize the reverse diffusion as a free Schrödinger bridge, revealing a precise energy–entropy trade-off in the non-commutative setting.
+  • Free Schrödinger bridge and energy–entropy balance: Through a free Girsanov theorem, we characterize the reverse diffusion as a free Schrödinger bridge, revealing a precise energy–entropy trade-off in the noncommutative setting.
   • Universality and spectral convergence: Under mild noise-schedule conditions, the forward process converges to the semicircular law, while the reverse dynamics can approximate any compactly supported spectral distribution.
-Taken together, these results position free diffusion models as Wasserstein gradient flows of Voiculescu’s free entropy, unifying stochastic diffusion, functional inequalities, and non-commutative information geometry under a single rigorous framework.
+I-B Information-Theoretic Implications
+The correspondence between free and classical entropy flows reveals a deep information-theoretic parallel.
+In classical diffusion-based generative models, the Fokker–Planck evolution drives Shannon entropy increase, and the reverse-time process recovers structure by following the gradient of the log-likelihood—a manifestation of entropy dissipation balanced by Fisher information flow.
+In the free setting, Voiculescu’s free entropy $\chi(\mu_{t})$ obeys an analogous dissipation law $\frac{\mathrm{d}}{\mathrm{d}t}\chi(\mu_{t})=\tfrac{1}{2}\Phi^{\ast}(\mu_{t})$, where $\Phi^{\ast}$ is the free Fisher information.
+This duality mirrors the de Bruijn identity and encodes how spectral measures lose and regain information under noncommutative diffusion.
+The reverse free SDE thus performs an information-preserving transport analogous to generative decoding, while the forward flow acts as a compression map toward a semicircular prior.
+Consequently, free diffusion models naturally admit a rate–distortion interpretation: entropy decay quantifies compression of operator-valued data, and the reverse flow reconstructs high-entropy signals along optimal information-geometric trajectories.
+Taken together, these results position free diffusion models as Wasserstein gradient flows of Voiculescu’s free entropy, unifying stochastic diffusion, functional inequalities, and noncommutative information geometry under a single rigorous framework.
 Beyond serving as a conceptual extension of classical diffusion models, this theory introduces new analytical tools for operator-valued and high-dimensional structured data, opening a research program at the intersection of free probability, optimal transport, and information-theoretic generative modeling.
@@
 This duality between entropy production (forward) and entropy dissipation (reverse) provides a geometric and information-theoretic foundation for diffusion-based generative modeling.
-It also sets the stage for the non-commutative generalization developed later in this paper, where Shannon entropy and Fisher information are replaced by their free-probabilistic counterparts.
+It also sets the stage for the noncommutative generalization developed later in this paper, where Shannon entropy and Fisher information are replaced by their free-probabilistic counterparts.
 II-D Applications and Impact
-Diffusion models have transformed several areas of artificial intelligence:
-  • Image, audio, and video generation: Samples from the reverse SDE (4) achieve state-of-the-art realism in large-scale models such as Imagen, DALL-E 2, and Stable Diffusion [24, 25, 26].
-  • Conditional generation: Conditioning mechanisms, such as classifier guidance, enable controllable text-to-image or text-to-audio synthesis.
-  • Uncertainty quantification: Score-based methods provide flexible tools for approximate Bayesian inference and posterior sampling [4].
-  • Scientific and physical modeling: Diffusion-based SDEs have been used to generate molecular conformations, simulate turbulent flows, and model protein structures.
-From a theoretical standpoint, the universality of diffusion models lies in their two complementary limits: the forward diffusion maps any data distribution to a Gaussian prior, while the reverse process—given an accurate score estimate—can recover arbitrarily complex distributions.
-This classical framework will serve as the benchmark for the free diffusion models developed in subsequent sections, where we extend these principles to non-commutative, operator-valued random variables.
+Diffusion models have rapidly emerged as one of the most influential frameworks in modern artificial intelligence, reshaping both the theoretical and practical landscape of generative modeling.
+Their success stems from a simple but powerful principle: by learning to reverse a gradual corruption process that transforms structured data into noise, these models can synthesize entirely new samples that are statistically indistinguishable from real data.
+Recent surveys such as [7] have highlighted their broad theoretical foundations and the unifying stochastic framework they provide across disciplines.
+  • High-fidelity image, audio, and video generation: Diffusion-based architectures now dominate large-scale generative modeling.
+Systems such as Imagen, Stable Diffusion, and DALL–E 2 produce photorealistic images, audio signals, and video clips through the reverse stochastic differential equation (SDE) (4).
+The ability to traverse from Gaussian noise to rich data manifolds via controlled stochastic dynamics has become a cornerstone of creative AI, surpassing the fidelity of GAN and flow-based models [24, 25, 26].
+  • Conditional and controllable synthesis: By integrating classifier guidance, cross-modal conditioning, or attention-based mechanisms, diffusion models enable text-to-image, image-to-speech, and even multimodal generation.
+The probabilistic structure of these models naturally accommodates conditioning and interpolation, providing a flexible platform for controllable synthesis and style transfer.
+  • Bayesian inference and uncertainty quantification: The connection between diffusion models and stochastic control [27, 4] allows them to perform approximate Bayesian inference.
+Score-based sampling schemes provide uncertainty estimates and enable posterior sampling in inverse problems, leading to applications in reinforcement learning, climate modeling, and medical imaging.
+  • Scientific and physical simulation: Diffusion-based stochastic dynamics have been successfully applied to model complex physical phenomena, ranging from molecular conformations and quantum energy landscapes to turbulent and fluid flows.
+Recent works have demonstrated diffusion generative models for molecular structure prediction [28, 29] and high-energy physics through calorimeter simulation frameworks [30].
+In these cases, diffusion models act as statistical surrogates for high-dimensional dynamical systems, blending physics-informed priors with generative flexibility.
+Theoretical significance.
+From a mathematical standpoint, diffusion models exhibit a remarkable universality: the forward diffusion maps any data distribution to a simple Gaussian reference, while the reverse-time SDE—parameterized by the score function—can, in principle, recover any target distribution.
+This bidirectional structure links the theory of diffusion models to the geometry of the Wasserstein space and to entropy-regularized optimal transport [31, 32].
+It also provides a continuous-time view of deep generative learning that bridges stochastic control, variational inference, and information geometry.
+Motivation for the free probabilistic framework.
+However, many emerging machine learning settings involve data that are inherently noncommutative or operator-valued.
+Examples include covariance matrices in signal processing, weight matrices in deep networks, density operators in quantum information, and graph Laplacians in network analysis.
+In such regimes, classical diffusion models—built on Euclidean and commutative probability—cannot capture the higher-order spectral interactions or non-commuting dependencies that drive these systems.
+The free probabilistic framework developed in this paper extends diffusion modeling into this operator-valued domain.
+By replacing Gaussian noise with semicircular noise and classical entropy with free entropy, we obtain a noncommutative analogue of the diffusion process that evolves along the geometry of the free Wasserstein space [16, 17].
+Practitioners can view this as a theoretical foundation for generative modeling of structured, matrix- or tensor-valued data—offering new analytical tools for understanding generalization, stability, and the spectral behavior of large models.
+Ultimately, free diffusion models bridge stochastic generative modeling with the deep structure of random matrix theory and noncommutative information geometry, suggesting a unified perspective on data generation across classical and operator-valued domains.
 III Free Probability Formulation
@@
 Their entries are highly correlated, yet their large-dimensional behavior often exhibits a form of asymptotic freeness rather than classical independence.
-This observation motivates the use of free probability theory as a natural non-commutative extension of diffusion modeling.
+This observation motivates the use of free probability theory as a natural noncommutative extension of diffusion modeling.
 Free probability, introduced by Voiculescu [11, 12], provides an operator-algebraic framework for describing random variables that do not commute but still exhibit a form of “statistical independence” called freeness.
@@
 $\varphi(f(X))=\int_{\mathbb{R}}f(\lambda)\,\mathrm{d}\mu_{X}(\lambda)\quad\text{for every bounded Borel function }f.$
-This spectral law generalizes the notion of a probability density to the non-commutative world.
+This spectral law generalizes the notion of a probability density to the noncommutative world.
 III-B Freeness
@@
 whenever $a_{j}\in\mathcal{A}_{i_{j}}$, $\varphi(a_{j})=0$, and consecutive indices differ ($i_{1}\neq i_{2}\neq\cdots\neq i_{k}$).
-This property means that alternating, centered elements from distinct subalgebras have zero expectation—the non-commutative analogue of the classical product rule $\mathbb{E}[XY]=\mathbb{E}[X]\mathbb{E}[Y]$ for independent variables.
+This property means that alternating, centered elements from distinct subalgebras have zero expectation—the noncommutative analogue of the classical product rule $\mathbb{E}[XY]=\mathbb{E}[X]\mathbb{E}[Y]$ for independent variables.
 III-C Free Brownian Motion and Semicircular Law
@@
 III-D Free Diffusion as a Model of Operator-Valued Noise
-Consider now a non-commutative diffusion process
+Consider now a noncommutative diffusion process
 $\mathrm{d}X_{t}=b_{t}\,\mathrm{d}t+\sigma_{t}\,\mathrm{d}S_{t},$
 where $X_{t},b_{t},\sigma_{t}\in\mathcal{A}_{\mathrm{sa}}$ and $S_{t}$ is free Brownian motion.
-The spectral measure $\mu_{X_{t}}$ evolves by free additive convolution—a non-commutative analogue of Gaussian convolution.
+The spectral measure $\mu_{X_{t}}$ evolves by free additive convolution—a noncommutative analogue of Gaussian convolution.
 When $b_{t}=-\tfrac{1}{2}\beta(t)X_{t}$ and $\sigma_{t}=\sqrt{\beta(t)}$, this becomes a free Ornstein–Uhlenbeck process [19], whose stationary law is the semicircular distribution.
@@
 III-F Summary
-Free probability thus provides a rigorous language for describing diffusion in non-commutative or matrix-valued settings.
+Free probability thus provides a rigorous language for describing diffusion in noncommutative or matrix-valued settings.
 The forward process corresponds to the flow of spectral measures under free additive convolution, while the reverse process represents an entropy-gradient descent in the geometry of free Wasserstein space.
@@
 V Free Fokker–Planck Dynamics and Free Clark–Ocone Formula
-Once the forward diffusion is lifted from classical probability to a non-commutative setting, the law of a free diffusion process $X_{s}$ evolves according to a free Fokker–Planck equation.
-Let $\mu_{X_{s}}$ denote the spectral distribution of $X_{s}$ in a tracial $W^{*}$-probability space $(\mathcal{A},\varphi)$, with dynamics
-$\mathrm{d}X_{s}=-\tfrac{1}{2}\beta(s)X_{s}\,\mathrm{d}s+\sqrt{\beta(s)}\,\mathrm{d}S_{s},$
-where $(S_{s})_{s\geq 0}$ is a free Brownian motion [13]. Then $\mu_{X_{s}}$ satisfies, in the distributional sense,
-$\frac{\partial}{\partial s}\mu_{X_{s}}(x)=\frac{1}{2}\beta(s)\frac{\partial}{\partial x}\big(x\mu_{X_{s}}(x)\big)+\frac{1}{2}\beta(s)\frac{\partial^{2}}{\partial x^{2}}\mu_{X_{s}}(x),$ (8)
-which is a free counterpart of the classical Fokker–Planck equation.
-Equation (8) governs the evolution of the spectral law, and can be shown to correspond to a gradient flow of Voiculescu’s free entropy $\chi(\mu_{X_{s}})$ in the metric space of spectral distributions endowed with the free 2-Wasserstein distance [14, 27].
+Once the forward diffusion is lifted from classical to noncommutative probability, the law of a free diffusion process $X_{s}$ evolves according to a free Fokker–Planck equation.
+Let $\mu_{X_{s}}$ denote the spectral measure of $X_{s}$ in a tracial $W^{*}$-probability space $(\mathcal{A},\phi)$ satisfying
+$dX_{s}=-\tfrac{1}{2}\beta(s)X_{s}\,ds+\sqrt{\beta(s)}\,dS_{s},$
+where $(S_{s})_{s\geq 0}$ is a free Brownian motion. Then the family of spectral measures $\{\mu_{X_{s}}\}_{s\geq 0}$ obeys, in the distributional sense,
+$\frac{\partial\mu_{X_{s}}}{\partial s}(x)=\frac{1}{2}\beta(s)\frac{\partial}{\partial x}\!\left[x\,\mu_{X_{s}}(x)\right]+\frac{1}{2}\beta(s)\frac{\partial^{2}\mu_{X_{s}}}{\partial x^{2}}(x),$ (8)
+which represents the commutative projection of the noncommutative evolution.
+Equation (8) is the gradient flow of Voiculescu’s free entropy $\chi(\mu_{X_{s}})$ in the metric space $(\mathcal{P}_{2}^{\mathrm{free}},W_{2,\mathrm{free}})$ introduced by Biane and Voiculescu [33].
+Equation (8) governs the evolution of the spectral law, and can be shown to correspond to a gradient flow of Voiculescu’s free entropy $\chi(\mu_{X_{s}})$ in the metric space of spectral distributions endowed with the free $L_{2}$-Wasserstein distance [14, 34].
 The derivation of the reverse-time SDE in the free setting requires conditioning with respect to the free filtration and exploiting the free analogue of the martingale representation theorem.
@@
 Lemma 1 (Free Clark–Ocone formula).
-Let $(S_{t})_{t\geq 0}$ be a free Brownian motion on $(\mathcal{A},\varphi)$, and let $F\in L^{2}(\mathcal{A},\varphi)$ be square-integrable with respect to the non-commutative filtration $(\mathcal{F}_{t})$ generated by $S_{t}$.
-Then there exists a unique adapted process $\mathbb{E}[\mathcal{D}_{u}F\mid\mathcal{F}_{u}]$ such that
-$F=\mathbb{E}[F]+\int_{0}^{T}\mathbb{E}[\mathcal{D}_{u}F\mid\mathcal{F}_{u}]\,\mathrm{d}S_{u},$ (9)
-where $\mathcal{D}_{u}$ denotes the free Malliavin derivative. This decomposition is the non-commutative analogue of the classical Clark–Ocone representation [14, 28].
+Let $(S_{t})_{t\geq 0}$ be a free Brownian motion on a tracial $W^{*}$-probability space $(\mathcal{A},\phi)$ with canonical filtration $(\mathcal{F}_{t})_{t\geq 0}$.
+For any $F\!\in\!L^{2}(\mathcal{A},\phi)$ adapted to $(\mathcal{F}_{t})$ whose free Malliavin derivative $D_{t}F$ belongs to $L^{2}(\mathcal{A}\,\bar{\otimes}\,\mathcal{A},\phi\!\otimes\!\phi)$, there exists a unique adapted process $u_{t}\in L^{2}_{\mathrm{ad}}([0,T];\mathcal{A})$ such that
+$F=\phi[F]+\int_{0}^{T}\mathbb{E}_{\phi}[D_{t}F\mid\mathcal{F}_{t}]\,\#\,dS_{t},$ (9)
+where #\# denotes the free stochastic integral and $\mathbb{E}_{\phi}[\cdot\mid\mathcal{F}_{t}]$ the conditional expectation in $(\mathcal{A},\phi)$.
+The proof follows standard arguments in free Malliavin calculus and is given in Appendix C (see also [14, 35]).
 Lemma 1 is the cornerstone of free stochastic calculus: it allows one to express functionals of the free Brownian motion as stochastic integrals, which in turn is essential for deriving the drift correction term in the reverse-time SDE.
@@
 $\mathrm{Cov}_{\tau}(F,G)=\mathbb{E}_{\tau}\!\left[\langle DF,DG\rangle_{\mathrm{free}}\right]=\mathbb{E}_{\tau}\!\left[F\big(\Xi_{t}+\tfrac{1}{2}X_{t}\big)G\right]-\mathbb{E}_{\tau}[F]\,\mathbb{E}_{\tau}[G]$ (11)
-extends Voiculescu’s classical free integration-by-parts identity from non-commutative polynomials to the full Sobolev space $W^{1,2}_{\mathrm{free}}(\mu_{t})$.
+extends Voiculescu’s classical free integration-by-parts identity from noncommutative polynomials to the full Sobolev space $W^{1,2}_{\mathrm{free}}(\mu_{t})$.
 Proof.
-Let $F\in\mathrm{Dom}(D)$ and approximate it by a sequence of self-adjoint non-commutative polynomials $\{P_{n}(X_{t})\}_{n\geq 1}$ such that $P_{n}(X_{t})\to F$ in $L^{2}(\mathcal{A},\tau)$ and $DP_{n}(X_{t})\to DF$ in $L^{2}(\mathcal{A}\otimes\mathcal{A})$.
+Let $F\in\mathrm{Dom}(D)$ and approximate it by a sequence of self-adjoint noncommutative polynomials $\{P_{n}(X_{t})\}_{n\geq 1}$ such that $P_{n}(X_{t})\to F$ in $L^{2}(\mathcal{A},\tau)$ and $DP_{n}(X_{t})\to DF$ in $L^{2}(\mathcal{A}\otimes\mathcal{A})$.
 For each polynomial $P_{n}$, Voiculescu’s original integration-by-parts formula [12, Prop.
@@
 where $\mu_{X_{0}}$ is compactly supported with $\chi(\mu_{X_{0}})>$-$\infty$ and $\Phi^{\ast}(\mu_{X_{0}})<\infty$.
-Assume that $(X_{t})_{0\leq t\leq T}$ exists uniquely in the free Itô sense, and that each $\mu_{X_{t}}$ admits a conjugate variable $\Xi_{t}\in L^{2}(W^{\ast}(X_{t}),\varphi)$.
-Then the time-reversed process $\widetilde{X}_{s}\mathrel{\mathop{\ordinarycolon}}=X_{T-s}$ satisfies the stochastic differential equation
+  1. (i)
+    The forward equation (F) admits a unique strong solution $(X_{t})_{0\leq t\leq T}\subset L^{2}(\mathcal{A},\varphi)$ in the sense of free Itô calculus, and for each $t$ the spectral law $\mu_{X_{t}}$ possesses a conjugate variable $\Xi_{t}\in L^{2}(W^{\ast}(X_{t}),\varphi)$.
+Existence and uniqueness in law follow from Kargin [36] and Biane–Speicher [14].
+  2. (ii)
+    The time-reversed process $\widetilde{X}_{s}\mathrel{\mathop{\ordinarycolon}}=X_{T-s}$ satisfies the reverse-time free stochastic differential equation
     $\mathrm{d}\widetilde{X}_{s}=\Bigl(-\tfrac{1}{2}\beta(T-s)\widetilde{X}_{s}-\beta(T-s)\,\Xi_{T-s}\Bigr)\,\mathrm{d}s+\sqrt{\beta(T-s)}\,\mathrm{d}\widetilde{S}_{s},$ (R)
     where $(\widetilde{S}_{s})_{s\in[0,T]}$ is a backward free Brownian motion adapted to the reversed filtration $\widetilde{\mathcal{F}}_{s}=\sigma\{\widetilde{X}_{u}\mathrel{\mathop{\ordinarycolon}}0\leq u\leq s\}$.
+Equation (R) is well posed in $L^{2}(\mathcal{A},\varphi)$ and its solution is unique in law.
 Proof.
-Step 1: Existence and forward representation. Let $\Lambda(t)=\int_{0}^{t}\beta(r)\,\mathrm{d}r$. By free Itô integration theory [14], the SDE (F) admits the explicit solution
-$X_{t}=e^{-\frac{1}{2}\Lambda(t)}X_{0}+\!\int_{0}^{t}e^{-\frac{1}{2}(\Lambda(t)-\Lambda(u))}\sqrt{\beta(u)}\,\mathrm{d}S_{u}.$
-The free Itô isometry ensures that $\|X_{t}\|_{2}^{2}<\infty$ and thus $X_{t}\in L^{2}(\mathcal{A},\varphi)$ for all $t\in[0,T]$.
-Step 2: Clark–Ocone representation. For each fixed $T$, the free Clark–Ocone formula (see Appendix C for a rigorous derivation) gives an adapted representation of $X_{T}$:
+Step 1: Existence and forward representation.
+Let $\Lambda(t)=\int_{0}^{t}\beta(r)\,\mathrm{d}r$.
+By the free Itô integration theory of Biane and Speicher [14] and the well-posedness results of Kargin [36], the SDE (F) admits a unique strong $L^{2}$ solution
+$X_{t}=e^{-\frac{1}{2}\Lambda(t)}X_{0}+\!\int_{0}^{t}e^{-\frac{1}{2}(\Lambda(t)-\Lambda(u))}\sqrt{\beta(u)}\,\mathrm{d}S_{u},$
+which satisfies $\|X_{t}\|_{2}<\infty$ for all $t\in[0,T]$ by the free Itô isometry.
+Step 2: Clark–Ocone representation. For fixed $T$, the free Clark–Ocone formula (Lemma 1, proved in Appendix C) provides the decomposition
 $X_{T}=\varphi(X_{T})+\int_{0}^{T}u_{t}\,\#\,\mathrm{d}S_{t},\qquad u_{t}=\mathbb{E}_{\varphi}[D_{t}X_{T}\mid\mathcal{F}_{t}].$
-Substituting the explicit solution of $X_{T}$ into the stochastic integral and applying the free Itô isometry yields
+Inserting the explicit forward solution and applying the free Itô isometry yields
 $u_{t}=e^{-\frac{1}{2}(\Lambda(T)-\Lambda(t))}\sqrt{\beta(t)},\qquad t\in[0,T].$
-This follows because $(\mathrm{d}S_{t})_{t\geq 0}$ has freely independent increments and $u_{t}$ is adapted; hence the conditional expectation $\mathbb{E}_{\varphi}[\cdot\mid\mathcal{F}_{t}]$ eliminates all stochastic terms beyond time $t$, leaving only the deterministic damping factor $e^{-\frac{1}{2}(\Lambda(T)-\Lambda(t))}$.
+Since the increments $(\mathrm{d}S_{t})$ are freely independent and $u_{t}$ is $\mathcal{F}_{t}$-adapted, conditional expectation removes future stochastic terms, leaving the deterministic factor $e^{-\frac{1}{2}(\Lambda(T)-\Lambda(t))}$.
 Step 3: Identification of the drift correction.
-By Lemma 3, for the log-density functional $G=\log\pi_{t}(X_{t})$ one has $\mathbb{E}_{\varphi}[D_{t}G\mid\mathcal{F}_{t}]=-\,\Xi_{t}$.
-Combining this with the derivative term $\tfrac{1}{2}\beta(t)X_{t}$ in the forward SDE (F) gives the backward drift $\tfrac{1}{2}\beta(t)X_{t}-\beta(t)\Xi_{t}$.
-Step 4: Time reversal. Define $\widetilde{X}_{s}=X_{T-s}$. By free stochastic calculus with respect to backward filtrations (see Biane [29], Sec. 3), the differential of $\widetilde{X}_{s}$ satisfies
-$\mathrm{d}\widetilde{X}_{s}=\Bigl(-\tfrac{1}{2}\beta(T-s)\widetilde{X}_{s}-\beta(T-s)\,\Xi_{T-s}\Bigr)\mathrm{d}s+\sqrt{\beta(T-s)}\,\mathrm{d}\widetilde{S}_{s},$
-where $\widetilde{S}_{s}\mathrel{\mathop{\ordinarycolon}}=S_{T}-S_{T-s}$ is a free Brownian motion under the reversed filtration.
-The integrability of the drift and diffusion coefficients follows from the square-integrability of $X_{t}$ and $\Xi_{t}$, ensuring well-posedness of (R) in $L^{2}(\mathcal{A},\varphi)$.
-Hence, the time-reversed process $\widetilde{X}_{s}$ is a weak solution of (R), completing the proof. ∎
+By Lemma 3, for the log-density functional $G=\log\pi_{t}(X_{t})$ we have $\mathbb{E}_{\varphi}[D_{t}G\mid\mathcal{F}_{t}]=-\,\Xi_{t}$.
+Combining this with the forward drift term $\tfrac{1}{2}\beta(t)X_{t}$ in (F) produces the backward drift $\tfrac{1}{2}\beta(t)X_{t}-\beta(t)\Xi_{t}$.
+Step 4: Time reversal and uniqueness in law.
+Define $\widetilde{X}_{s}=X_{T-s}$.
+By Biane’s time-reversal duality for free stochastic integrals [37, Prop.
+3.1], the process $\widetilde{S}_{s}\mathrel{\mathop{\ordinarycolon}}=S_{T}-S_{T-s}$ is a backward free Brownian motion adapted to $\widetilde{\mathcal{F}}_{s}$.
+Applying free Itô calculus for reversed filtrations yields
+$\mathrm{d}\widetilde{X}_{s}=\Bigl(-\tfrac{1}{2}\beta(T-s)\widetilde{X}_{s}-\beta(T-s)\,\Xi_{T-s}\Bigr)\mathrm{d}s+\sqrt{\beta(T-s)}\,\mathrm{d}\widetilde{S}_{s}.$
+The square-integrability of $X_{t}$ and $\Xi_{t}$ implies integrability of all coefficients, ensuring existence and uniqueness in law of the solution to (R) in $L^{2}(\mathcal{A},\varphi)$ by Kargin’s general theory of free SDEs [36].
+Hence $\widetilde{X}_{s}$ is the unique weak solution of (R), completing the proof.
+∎
 Completion of the proof. To make the time-reversal argument rigorous, recall that under the free Itô isometry [14, Prop. 3.2],
@@
 Remark 3.
-Both derivations show that the term $\Xi_{s}$ plays the role of the non-commutative score function in the reverse-time dynamics, and that the reversed diffusion acts as a free deconvolution flow.
+Both derivations show that the term $\Xi_{s}$ plays the role of the noncommutative score function in the reverse-time dynamics, and that the reversed diffusion acts as a free deconvolution flow.
 Corollary: Reverse Drift as a Gradient of Free Entropy
@@
 which establishes the lower semicontinuity in item 1.
-The argument is the free-probability analogue of the classical lower semicontinuity of $W_{2}$ on $\mathcal{P}_{2}(\mathbb{R})$, see also [31, Ch.
+The argument is the free-probability analogue of the classical lower semicontinuity of $W_{2}$ on $\mathcal{P}_{2}(\mathbb{R})$, see also [38, Ch.
 2].
@@
 For each fixed $R>0$, the sets of $m\times m$ microstates $\Gamma_{m}(\nu_{n};R,\varepsilon)$ corresponding to $\nu_{n}$ converge in the Hausdorff sense (with respect to the normalized trace topology) to $\Gamma_{m}(\nu;R,\varepsilon)$ as $n\to\infty$, because narrow convergence with uniformly bounded second moments implies convergence of all *-moments of bounded degree.
-By the monotonicity of $\mathrm{Vol}[\Gamma_{m}(\nu;R,\varepsilon)]$ in $R$ and $\varepsilon$, and the subadditivity of $\frac{1}{m^{2}}\log\mathrm{Vol}$ in $m$, one can apply a diagonal argument in $(m,R,\varepsilon)$ to pass to the limit (see [34, Thm.
+By the monotonicity of $\mathrm{Vol}[\Gamma_{m}(\nu;R,\varepsilon)]$ in $R$ and $\varepsilon$, and the subadditivity of $\frac{1}{m^{2}}\log\mathrm{Vol}$ in $m$, one can apply a diagonal argument in $(m,R,\varepsilon)$ to pass to the limit (see [41, Thm.
 6.5.1] and [40]), obtaining
@@
 Hence, the free entropy is upper semicontinuous along narrowly convergent sequences with uniformly bounded second moments.
-(iii) Consequences..
+(iii) Consequences.
 Since the free energy functional $\mathcal{F}[\nu]\mathrel{\mathop{\ordinarycolon}}=\tfrac{1}{2}W_{2,\mathrm{free}}^{2}(\nu,\mu_{\mathrm{sc}})-\chi[\nu]$ is the sum of a lower semicontinuous and an upper semicontinuous term, it is itself lower semicontinuous on $\mathcal{P}_{2}^{\mathrm{free}}$.
@@
 Proof.
-We will prove the three items in order. The proof strategy is standard (cf. Ambrosio–Gigli–Savaré [31, Ch. 4]), but we carefully check each step in the free setting using Lemma 4.
+We will prove the three items in order. The proof strategy is standard (cf. Ambrosio–Gigli–Savaré [38, Ch. 4]), but we carefully check each step in the free setting using Lemma 4.
 (I) Pointwise convergence $\mathcal{F}^{\tau}(\mu)\to\mathcal{F}(\mu)$.
@@
 $\mu_{k+1}^{\tau}\in\arg\min_{\nu\in\mathcal{P}_{2}^{\mathrm{free}}}\Big\{\frac{1}{2\tau}W_{2,\mathrm{free}}^{2}(\nu,\mu_{k}^{\tau})+\mathcal{F}[\nu]\Big\}.$ (20)
-Theorems 2 ($\Gamma$-convergence of $\mathcal{F}^{\tau}$) and standard minimizing-movement compactness arguments (see Ambrosio–Gigli–Savaré [31, Ch. 4]) imply the following.
+Theorems 2 ($\Gamma$-convergence of $\mathcal{F}^{\tau}$) and standard minimizing-movement compactness arguments (see Ambrosio–Gigli–Savaré [38, Ch. 4]) imply the following.
 Having established the $\Gamma$-convergence of the discrete functionals $\mathcal{F}_{\tau}$ to $\mathcal{F}$, we now turn to the dynamic aspect of the minimizing-movement scheme.
@@
 Theorem 3 (Existence and convergence of the JKO scheme in the free setting).
-Let $(\mathcal{P}_{2}^{\mathrm{free}},W_{2,\mathrm{free}})$ denote the space of compactly supported spectral measures with finite second moment, endowed with the free $2$-Wasserstein distance of Biane–Voiculescu.
+Let $(\mathcal{P}_{2}^{\mathrm{free}},W_{2,\mathrm{free}})$ denote the space of compactly supported spectral measures with finite second moment, endowed with the free $L_{2}$-Wasserstein distance of Biane–Voiculescu.
 Assume that the free energy functional
@@
 Step 6: $\lambda$-convexity and uniqueness (EVI).
-If $\mathcal{F}$ is $\lambda$-geodesically convex in $(\mathcal{P}_{2}^{\mathrm{free}},W_{2,\mathrm{free}})$, then standard variational arguments (see [31, Ch.
+If $\mathcal{F}$ is $\lambda$-geodesically convex in $(\mathcal{P}_{2}^{\mathrm{free}},W_{2,\mathrm{free}})$, then standard variational arguments (see [38, Ch.
 4]) yield the differential contractivity estimate
@@
 $\mathcal{F}[\mu_{\theta}]\;\leq\;(1-\theta)\,\mathcal{F}[\mu_{0}]+\theta\,\mathcal{F}[\mu_{1}]-\frac{\lambda}{2}\,\theta(1-\theta)\,W_{2,\mathrm{free}}^{2}(\mu_{0},\mu_{1}).$
-This holds, for instance, for the quadratic potential plus free entropy under compact support (see the discussion in [34, Sec. 6] and [16].
+This holds, for instance, for the quadratic potential plus free entropy under compact support (see the discussion in [41, Sec. 6] and [16].
 Define the metric derivative $|\mu^{\prime}|(t)$ of an absolutely continuous curve $t\mapsto\mu_{t}$ in $(\mathcal{P}_{2}^{\mathrm{free}},W_{2,\mathrm{free}})$ and the local metric slope
 $|\partial\mathcal{F}|(\mu)\mathrel{\mathop{\ordinarycolon}}=\limsup_{\nu\to\mu}\frac{\big(\mathcal{F}[\mu]-\mathcal{F}[\nu]\big)_{+}}{W_{2,\mathrm{free}}(\mu,\nu)}.$
-In the present free setting, using the continuity equation representation and Lemma 4, one identifies (cf. [31, Ch. 10])
+In the present free setting, using the continuity equation representation and Lemma 4, one identifies (cf. [38, Ch. 10])
 $|\partial\mathcal{F}|(\mu_{t})\;=\;\Big\|\,b_{t}\,\Big\|_{L^{2}(\mu_{t})}\qquad\text{with}\quad b_{t}(x)=-\frac{1}{2}x-\Xi_{t}(x),$ (ID)
@@
 Proof.
-The standard minimizing-movement compactness and a priori bounds from [31, Chs.
+The standard minimizing-movement compactness and a priori bounds from [38, Chs.
 2–3 & 4] apply verbatim by Lemma 4, yielding (up to subsequences) $\mu^{\tau}(t)\rightharpoonup\mu_{t}$ and the discrete energy dissipation estimate
@@
 Summing over $k$ and passing to the limit gives the EDI (25).
-Under (VII), the AGS theory yields the EVI (26) and uniqueness of the gradient flow [31, Thm.
+Under (VII), the AGS theory yields the EVI (26) and uniqueness of the gradient flow [38, Thm.
 4.0.4 & Ch.
@@
 Remark 5 (Energy dissipation equality).
-If $t\mapsto\mu_{t}$ is a curve of maximal slope for $\mathcal{F}$ (i.e., equality holds in (25)), then the metric derivative and slope coincide a.e.: $|\mu^{\prime}|(t)=|\partial\mathcal{F}|(\mu_{t})=\|-\tfrac{1}{2}x-\Xi_{t}\|_{L^{2}(\mu_{t})}$, which is precisely the Fisher-information dissipation identity in the free setting; see [31, Ch.
-10] and [34, Sec.
+If $t\mapsto\mu_{t}$ is a curve of maximal slope for $\mathcal{F}$ (i.e., equality holds in (25)), then the metric derivative and slope coincide a.e.: $|\mu^{\prime}|(t)=|\partial\mathcal{F}|(\mu_{t})=\|-\tfrac{1}{2}x-\Xi_{t}\|_{L^{2}(\mu_{t})}$, which is precisely the Fisher-information dissipation identity in the free setting; see [38, Ch.
+10] and [41, Sec.
 6].
@@
 They also form the analytical basis for the summary given in Appendix D.
-The forward free diffusion process defined in (7) admits a natural interpretation in terms of non-commutative information measures.
+The forward free diffusion process defined in (7) admits a natural interpretation in terms of noncommutative information measures.
 In classical information theory, Gaussian heat flow increases Shannon entropy and satisfies the de Bruijn identity, which relates the derivative of the entropy to the Fisher information.
@@
 and $\Xi_{s}$ is the conjugate variable of $X_{s}$.
-Equation (29) is the non-commutative analogue of the classical de Bruijn identity and shows that the forward free diffusion is an entropy-increasing flow.
+Equation (29) is the noncommutative analogue of the classical de Bruijn identity and shows that the forward free diffusion is an entropy-increasing flow.
 In particular, $\chi(X_{s})$ is non-decreasing in $s$, with strict monotonicity unless $X_{s}$ is already semicircular.
@@
 The conjugate variable $\Xi_{s}$ plays the same role in free probability as the score function $\nabla\log p(x)$ in classical statistics.
-It represents the $L^{2}$-gradient of the log non-commutative density and encodes how the spectral law of $X_{s}$ deviates from the equilibrium semicircular law.
+It represents the $L^{2}$-gradient of the log noncommutative density and encodes how the spectral law of $X_{s}$ deviates from the equilibrium semicircular law.
 The reverse-time dynamics of Theorem 1 can thus be viewed as a free analogue of the score-based generative modeling framework.
@@
 Proof.
-For a single self-adjoint non-commutative random variable, Voiculescu’s free entropy has the variational derivative
+For a single self-adjoint noncommutative random variable, Voiculescu’s free entropy has the variational derivative
 $\nabla\chi[\mu](x)=\frac{1}{2}x+\Xi(x),$
@@
 $\frac{\mathrm{d}}{\mathrm{d}t}\chi(X_{t})=\frac{1}{2}\Phi^{*}(X_{t}),$
-and the free Fisher information is connected to the derivative of $\chi$ through the non-commutative analogue of the classical de Bruijn identity.
+and the free Fisher information is connected to the derivative of $\chi$ through the noncommutative analogue of the classical de Bruijn identity.
 The free analog of the Fisher information inequality, due to Biane and Voiculescu, states that for free convolution with a semicircular element $S_{t}$,
@@
 $\text{(Free HWI)}\;\Rightarrow\;\text{(Free Log--Sobolev and Free Talagrand)}\;\Rightarrow\;\text{(Exponential Entropy Decay).}$
-Thus, the free diffusion model satisfies a hierarchy of functional inequalities entirely analogous to the classical Gaussian case, now formulated in non-commutative Wasserstein geometry.
+Thus, the free diffusion model satisfies a hierarchy of functional inequalities entirely analogous to the classical Gaussian case, now formulated in noncommutative Wasserstein geometry.
 This establishes the complete correspondence between free stochastic dynamics, functional inequalities, and information geometry.
@@
 defines a generative process.
-The classical theory of Schrödinger bridges, together with its non-commutative analogue, shows that for any absolutely continuous target law $\mu_{0}$, there exists a control $U_{t}=-\Xi_{t}$ that transports $\mu_{T}$ back to $\mu_{0}$.
+The classical theory of Schrödinger bridges, together with its noncommutative analogue, shows that for any absolutely continuous target law $\mu_{0}$, there exists a control $U_{t}=-\Xi_{t}$ that transports $\mu_{T}$ back to $\mu_{0}$.
 Theorem 7 (Universality of Free Diffusion Models).
@@
 $\chi(\mu)$ Voiculescu’s free entropy of a measure $\mu$
-$W_{2,\mathrm{free}}$ Free 2-Wasserstein distance on $\mathcal{P}_{2}^{\mathrm{free}}$
+$W_{2,\mathrm{free}}$ Free $L_{2}$-Wasserstein distance on $\mathcal{P}_{2}^{\mathrm{free}}$
 $\mathcal{F}[\mu]$ Free energy functional $\mathcal{F}[\mu]=-\chi(\mu)$
@@
 The reverse-time SDE can also be obtained from the PDE governing the evolution of the spectral law of $X_{t}$ under the forward SDE.
-This derivation requires no stochastic calculus beyond identifying the conjugate variable with the non-commutative gradient of the log-density.
+This derivation requires no stochastic calculus beyond identifying the conjugate variable with the noncommutative gradient of the log-density.
 Let $\mu_{t}\mathrel{\mathop{\ordinarycolon}}=\mu_{X_{t}}$ denote the spectral measure of $X_{t}$ for $t\in[0,T]$.
-Under the forward SDE (LABEL:(F)), $\mu_{t}$ evolves according to the free Fokker–Planck equation
+Under the forward SDE (7), $\mu_{t}$ evolves according to the free Fokker–Planck equation
 $\partial_{t}\mu_{t}=\frac{1}{2}\beta(t)\,\partial_{x}\big(x\,\mu_{t}\big)+\frac{1}{2}\beta(t)\,\partial_{x}\!\big(\mu_{t}\,\partial_{x}V_{t}\big),$ (49)
@@
 In particular, if one were to discretize the reversed SDE and simulate with the optimal control $U_{t}=-\Xi_{t}$, the process initialized at $\mathsf{SC}(m_{T},\sigma_{T}^{2})$ will be transported back toward the point mass at $a$ as $t\downarrow 0$.
-This closed-form example thus verifies the abstract formulas of Theorem 1: the conjugate variable appears as the non-commutative score, and it yields an explicit analytic expression for the reverse-time drift in the freely Gaussian (semicircular) case.
+This closed-form example thus verifies the abstract formulas of Theorem 1: the conjugate variable appears as the noncommutative score, and it yields an explicit analytic expression for the reverse-time drift in the freely Gaussian (semicircular) case.
 ∎
@@
 This appendix presents the analytic foundation underlying Theorem 1.
-We first derive the free Clark–Ocone decomposition—the noncommutative analogue of the classical martingale representation theorem—and then justify the time–reversal identity used to obtain the reverse–time free SDE.
+We first derive and prove Lemma 1—the free Clark–Ocone decomposition, a noncommutative analogue of the classical martingale representation theorem—and then justify the time–reversal identity used to obtain the reverse–time free SDE.
 Throughout, $(\mathcal{A},\varphi)$ denotes a tracial $W^{\ast}$–probability space, and $(S_{t})_{t\geq 0}$ a free Brownian motion (semicircular process) with forward filtration $\mathcal{F}_{t}=\sigma\{S_{u}\mathrel{\mathop{\ordinarycolon}}0\leq u\leq t\}$.
@@
 $\varphi[F\,\delta(U)]=\int_{0}^{T}\varphi\!\big(\langle D_{t}F,\,U_{t}\rangle_{\mathrm{free}}\big)\,\mathrm{d}t,\qquad\delta=D^{\ast}.$ (52)
-C.2  Free Clark–Ocone Decomposition
-Let $\mathcal{H}_{\mathrm{pred}}$ denote the closed subspace of $L^{2}([0,T]\times\mathcal{A}\!\otimes\!\mathcal{A})$ consisting of predictable (i.e.
-$\mathcal{F}_{t}$–measurable) processes.
-For $F\in\mathrm{Dom}(D)$, define the predictable projection
-$u_{t}\mathrel{\mathop{\ordinarycolon}}=\mathbb{E}_{\varphi}[D_{t}F\mid\mathcal{F}_{t}],$
-where $\mathbb{E}_{\varphi}[\cdot\mid\mathcal{F}_{t}]$ denotes the orthogonal projection of $L^{2}(\mathcal{A},\varphi)$ onto $L^{2}(\mathcal{F}_{t},\varphi)$.
-Then $u=(u_{t})_{t}$ is the unique element of $\mathcal{H}_{\mathrm{pred}}$ minimizing $\|D_{t}F-v_{t}\|_{L^{2}}$ over all predictable $v_{t}$; that is, $u$ is the orthogonal projection of $DF$ onto $\mathcal{H}_{\mathrm{pred}}$.
-Representation theorem.
-Combining (52) with the orthogonality of the projection, for all $G\in\mathrm{Dom}(\delta)$,
-$\varphi[F\,\delta(G)]=\int_{0}^{T}\varphi[\langle D_{t}F,\,G_{t}\rangle_{\mathrm{free}}]\,\mathrm{d}t=\int_{0}^{T}\varphi[\langle u_{t},\,G_{t}\rangle_{\mathrm{free}}]\,\mathrm{d}t=\varphi\!\left[\Big(\!\int_{0}^{T}u_{t}\,\#\,\mathrm{d}S_{t}\!\Big)G\right].$
-Since $\mathrm{Dom}(\delta)$ is dense in $L^{2}(\mathcal{A},\varphi)$, the representing element is unique, and we obtain the free Clark–Ocone formula
-$F=\mathbb{E}_{\varphi}[F]+\int_{0}^{T}\mathbb{E}_{\varphi}[D_{t}F\mid\mathcal{F}_{t}]\,\#\,\mathrm{d}S_{t},\qquad F\in W^{1,2}_{\mathrm{free}}.$ (53)
+C.2  Proof of Lemma 1 (Free Clark–Ocone Formula)
+The operator $D$ is densely defined and closable on $L^{2}(\mathcal{A},\varphi)$, and $\delta=D^{\ast}$ by (52).
+For $F\in\mathrm{Dom}(D)$, define the predictable projection $u_{t}\mathrel{\mathop{\ordinarycolon}}=\mathbb{E}_{\varphi}[D_{t}F\mid\mathcal{F}_{t}]$.
+Then $u\in L^{2}_{\mathrm{ad}}([0,T];\mathcal{A})$ and $\delta(u)=\int_{0}^{T}u_{t}\,\#\,\mathrm{d}S_{t}$ is well defined.
+Using orthogonality of the conditional expectation and the adjointness relation, one obtains
+$\varphi[F\,\delta(G)]=\int_{0}^{T}\varphi[\langle D_{t}F,G_{t}\rangle_{\mathrm{free}}]\,\mathrm{d}t=\int_{0}^{T}\varphi[\langle u_{t},G_{t}\rangle_{\mathrm{free}}]\,\mathrm{d}t=\varphi\!\left[\!\Big(\!\int_{0}^{T}u_{t}\,\#\,\mathrm{d}S_{t}\!\Big)G\!\right],$
+for all $G\in\mathrm{Dom}(\delta)$.
+By density of $\mathrm{Dom}(\delta)$ in $L^{2}(\mathcal{A},\varphi)$, we identify $F-\varphi[F]=\delta(u)$, yielding (9).
+Uniqueness of $u$ follows from the free Itô isometry (51): if $\int_{0}^{T}(u_{t}-v_{t})\#\mathrm{d}S_{t}=0$ then $u=v$ in $L^{2}$.
+$\square$
 Remarks.
@@
     $\delta$ is closed and densely defined, guaranteeing that the stochastic integral in (9) is well posed for any predictable $u$ in $L^{2}$.
-  3. (iii)
-    Uniqueness of the integrand follows from the isometry (51): if two predictable processes $u,v$ satisfy (53), then $\int_{0}^{T}(u_{t}-v_{t})\,\#\,\mathrm{d}S_{t}=0$ implies $u=v$ in $L^{2}$.
 Applied to the terminal variable $F=X_{T}$ of the forward free Ornstein–Uhlenbeck SDE, the formula gives
@@
 C.3  Free Time–Reversal of Semicircular Diffusions
-The rigorous time–reversal step in Theorem 1 follows from Biane’s duality theorem for free stochastic integrals [29, Proposition 3.1].
+The rigorous time–reversal step in Theorem 1 follows from Biane’s duality theorem for free stochastic integrals [37, Proposition 3.1].
 For any adapted process $U_{t}\in L^{2}(\mathcal{A},\varphi)$ define
@@
 D-A Voiculescu’s Free Entropy Functional
-For a self-adjoint non-commutative random variable $X$ with spectral distribution $\mu$, Voiculescu’s free entropy $\chi(\mu)$ is defined by
+For a self-adjoint noncommutative random variable $X$ with spectral distribution $\mu$, Voiculescu’s free entropy $\chi(\mu)$ is defined by
 $\chi(\mu)=\iint\log|x-y|\,d\mu(x)d\mu(y)-\frac{1}{2}\int x^{2}d\mu(x).$ (54)
@@
 D-D Euler–Lagrange Equation and Free McKean–Vlasov Flow
-The free entropy $\chi(\mu)$ can be regarded as the non-commutative analogue of the classical McKean–Vlasov free energy functional
+The free entropy $\chi(\mu)$ can be regarded as the noncommutative analogue of the classical McKean–Vlasov free energy functional
 $\mathcal{F}(\mu)=\int V(x)\,d\mu(x)-\iint\log|x-y|\,d\mu(x)d\mu(y),$